### v1

# Organizing Unstructured Image Collections using Natural Language

Abstract
Organizing unstructured visual data into semantic clusters is a key challenge in computer vision. Traditional deep clustering (DC) approaches focus on a single partition of data, while multiple clustering (MC) methods address this limitation by uncovering distinct clustering solutions. The rise of large language models (LLMs) and multimodal LLMs (MLLMs) has enhanced MC by allowing users to define clustering criteria in natural language. However, manually specifying criteria for large datasets is impractical. In this work, we introduce the task Semantic Multiple Clustering (SMC) that aims to automatically discover clustering criteria from large image collections, uncovering interpretable substructures without requiring human input. Our framework, Text Driven Semantic Multiple Clustering (TeDeSC), uses text as a proxy to concurrently reason over large image collections, discover partitioning criteria, expressed in natural language, and reveal semantic substructures. To evaluate TeDeSC, we introduce the COCO-4c and Food-4c benchmarks, each containing four grouping criteria and ground-truth annotations. We apply TeDeSC to various applications, such as discovering biases and analyzing social media image popularity, demonstrating its utility as a tool for automatically organizing image collections and revealing novel insights.
Disclaimer: Potentially sensitive content.
Figure 1: Top: Semantic Multiple Clustering (SMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable groups (or semantic clusters), under multiple shared themes (or clustering criteria), without requiring any prior information. Bottom: Our proposed SMC system enables various applications like discovering novel biases in text-to-image generative (T2I) models.
1 Introduction
Organizing large volumes of unstructured visual data into semantic clusters is a crucial problem in computer vision and has traditionally been addressed through the lens of unsupervised deep clustering (DC) (Caron et al., 2018; Van Gansbeke et al., 2020). Despite the advancements and improved performance, DC remains limited due to its inherently ill-posed nature (Estivill-Castro, 2002), as a dataset can be clustered in multiple ways (e.g., a dataset of a standard deck of cards can be grouped by suit or rank). Furthermore, the clusters discovered by DC rely on factors such as the inductive biases of the network, augmentations, and pretext tasks; and do not necessarily adhere to the a particular partitioning criteria that a user may have in mind (e.g., suit).
The limitation of DC in exploring visual data only along a single partition is addressed by multiple clustering (MC) methods (Yu et al., 2024), which uncover distinct clustering solutions. Multiple partitions reveal hidden patterns in the data and allow for analysis from different perspectives. The recent advent of large language models (LLMs) and multimodal LLMs (MLLMs) has further enhanced MC by enabling users to specify the clustering criterion (e.g., group by suit) in natural language (Kwon et al., 2024; Yao et al., 2024). While incorporating a human in the loop is appealing, as it grants the user control over the clustering results, specifying a meaningful text criterion still requires prior understanding of the image collections. This task becomes cumbersome, especially due to the proliferation and high scene complexity inherent in large volumes of visual data.
[…]
Prior User-provided Criteria
Knowledge # Clusters
Out. Multiple Clustering
Interpretability
In this work, we propose to ease the burden of specifying text criteria off the user by automatically discovering the clustering criteria, and introduce the task of Semantic Multiple Clustering (SMC), as shown in Fig. 1(top). To tackle SMC, we propose a general two-stage framework Text Driven Semantic Multiple Clustering (TeDeSC), powered by cutting-edge MLLMs and LLMs, that first discovers latent criteria (e.g., Activity and Location) from a given collection of images, and then groups the images into semantic clusters (e.g., “Surfing”,“Skateboarding” under Activity) for every discovered clustering criterion. To reason at a dataset level, TeDeSC first translates the visual data into textual data (e.g., captions), which in turn are then collectively used as a proxy for the LLM to discover the hidden patterns in the image collection. The uniqueness of TeDeSC lies in its: (i) comprehensiveness, as it can exhaustively discover multiple clustering criteria that may not be evident to a user; (ii) interpretability, as it outputs cluster names in natural language as opposed to traditional MC methods; (iii) flexibility, as it can discover clusters at multiple granularities (coarse to fine-grained), without needing to specify the number of clusters a priori; and (iv) generality, as, unlike the existing work (Yao et al., 2024), it is not limited to object-centric images but can handle complex scenes having fine-grained details. We summarize the key differences among various clustering solutions in Tab. 1 and elaborate in App. A.
Current benchmarks for evaluating multiple clustering either lack realism (Clevr-4c, Fruit-2c, Card-2c) or offer a limited number of criteria (Action-3c), as shown in Fig. 2. To advance research in SMC, we introduce two challenging new benchmarks, COCO-4c and Food-4c, which depict images in daily life contexts and allow the dataset to be clustered as per four distinct criteria. Extensive experimental analyses on both existing and novel benchmarks demonstrate that the proposed TeDeSC can effectively discover meaningful clustering criteria and successfully group semantically similar images (some examples of clustering results by TeDeSC are shown in Fig. 1 (top)).
Lastly, we apply TeDeSC to a variety of applications, including discovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images. For example, as shown in Fig. 1(bottom), TeDeSC can uncover less commonly studied biases (e.g., Hair color) in DALL·E3-generated (Betker et al., 2023) images, beyond the well-known biases (e.g., Gender) that may have already been corrected for. It achieves this by discovering various human-interpretable semantic clusters (e.g., “Dark Hair”, “Grey Hair”) across different discovered criteria and identifying overpopulated clusters. Similarly, it can identify the semantic visual elements that contribute to the popularity of social images, offering valuable insights to related practitioners. These results suggest that TeDeSC is an automatic, versatile, and highly practical tool, opening up numerous new application opportunities for future research and providing the potential to generate insights from unstructured visual data at scale across various domains.
2 Semantic Multiple Clustering
In this section, we first formally define the task of Semantic Multiple Clustering (SMC), and then introduce the existing and newly proposed benchmarks for SMC.
Task Definition: Given a collection of $N$ unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of SMC is to categorize $\mathcal{D}$ into semantic clusters based on multiple criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$, where $L$ is the total number of criteria. In detail, a criterion $R_{l}$ refers to the grouping theme according to which a set of images can be organized (or partitioned), such as Activity, Location, Mood, or Time of Day. All semantic clusters $\mathcal{C}_{l}$ under a given criterion $R_{l}$ should align with the theme of that criterion. For instance, the semantic clusters under the criterion Activity should reflect activities depicted in the images (e.g., “Kayaking”, “Cooking”). The same dataset $\mathcal{D}$, when grouped by a different criterion such as Location, could be clustered into categories like “Outdoor”, “Indoor”, and so on.
[…]
Figure 2: Overview of the SMC benchmarks. We show all clustering criteria and the corresponding ground-truth labels for the example images. We introduce two new challenging benchmarks: COCO-4c and Food-4c.
Benchmarks: Evaluating SMC methods requires benchmarks that can be partitioned under multiple valid criteria. Currently, only a few benchmarks (Yu et al., 2024) support the evaluation of SMC methods: Fruit-2c (Muresan & Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), and Clevr-4c (Vaze et al., 2024). As shown in Fig. 2, these benchmarks are limited by their object-centric nature with simple backgrounds (e.g., Fruit-2c), an insufficient number of criteria (e.g., up to three in Action-3c), and a lack of photorealism due to synthetic generation (e.g., Clevr-4c).
Given that the data encountered in real-world applications is more complex, we propose two new benchmarks for SMC: Food-4c and COCO-4c. Food-4c is sourced from Food-101 (Bossard et al., 2014), which includes 101 Food type (original annotations), along with new annotations for 15 Cuisine types, 5 Courses types, and 4 Diet preferences, totaling four clustering criteria. Additionally, we constructed COCO-4c using images from COCO-val (Lin et al., 2014), where we annotated four criteria with varying number of clusters: 64 Activity, 19 Location, 20 Mood, and 6 Time of day. Examples of these newly constructed benchmarks are shown in Fig. 2. Further details, including the full list of cluster names and the annotation pipeline, are provided in App. B.
3 Methodology: Text Driven Semantic Multiple Clustering
Partitioning an unstructured image collection into semantic clusters based on different (unknown) criteria is challenging, as it requires reasoning over the visual contents of all the images concurrently. Specifically, the system needs to first find commonalities among the images for discovering the partitioning criterion (or theme) and then group the images into semantic clusters according to the discovered criterion. To tackle the challenging SMC task, we propose a two-stage framework, named Text Driven Semantic Multiple Clustering (TeDeSC), that significantly deviates from the commonly used technique of deep feature-based clustering, and instead uses text descriptions as a proxy to reason over the images and uncover hidden patterns in $\mathcal{D}$. We elaborate the key design differences between TeDeSC and related work in App. A.
System overview: As illustrated in Fig. 3, the proposed TeDeSC consists of two modules: Criteria Proposer and Semantic Grouper. The Criteria Proposer (or Proposer in short) processes the entire image set $\mathcal{D}$ to discover diverse common themes among the images and proposes grouping criteria $\mathcal{R}$ in natural language (e.g., Location). Once the criteria are proposed, the Semantic Grouper (or Grouper in short) discovers distinct semantic clusters that adhere to each criterion $R_{l}$ at varying levels of semantic granularity and assigns images to their respective clusters (e.g., “Climbing gym”). In this work, we explore various design choices for both the Proposer and Grouper, detailed in the following subsections. Full implementation details, including exact prompts, are provided in App. C.
Figure 3: TeDeSC is composed of a Criteria Proposer and a Semantic Grouper. (Left) The Proposer processes the entire image set to discover and propose diverse grouping criteria expressed in natural language, which are accumulated in the Criteria Pool. (Middle) The Grouper takes these proposed criteria from the pool, discovers semantic clusters linked to the criteria at three different levels of granularity, and assigns images to their respective clusters. (Right) Multiple substructures of the image set, indicated by the discovered criteria, are revealed across different granularities by combining the cluster assignment results. We explore various design choices, and set the best-performing methods from our experiments as the main configuration (marked with Refer to caption).
3.1 Criteria Proposer
As shown in Fig. 3(left), the Proposer takes the target image collection as input and outputs a set of discovered grouping criteria, expressed in natural language. The core of the Proposer design lies in its ability to concurrently analyze and reason over the visual content of a large set of images. With this in mind, we explore three systematic approaches, which we describe below.
Image-based Proposer: We start with a baseline that leverages a state-of-the-art MLLM (Li et al., 2024) designed for multi-image reasoning out-of-the-box to directly infer the criteria given a set of images. To reason over a set of images, which is crucial for discovering the criterion, we stitch a batch of images into a $8\times 8$ image grid and input it to the MLLM as a single image. We then prompt the MLLM to propose grouping criteria for the images in the grid. All the resulting criteria from all such image grids are accumulated in a criteria pool, denoted as $\widetilde{\mathcal{R}}$.
Tag-based Proposer: Next, we explore a tag-based approach that uses an open-vocabulary tagger (Radford et al., 2021) to assign 10 tags to each image in $\mathcal{D}$, using the vocabulary from WordNet (Miller, 1995). These tags serve as semantic descriptors, translating the visual content of each image into textual form. We then gather all the assigned tags from the images, input them into a LLM, and prompt it to discover grouping criteria based on these tags.
Caption-based Proposer (Main): While image tags effectively capture certain visual semantics, we found that they predominantly reflect object-related content. To encompass a broader spectrum of visual information—such as environmental settings or interactions—we instead use a MLLM to generate captions for each image in $\mathcal{D}$. Descriptive captions provide a richer and more holistic semantic context. Staying within the 128k token limit of modern LLMs (Meta, 2024b; OpenAI, 2024), we feed a subset of captions into a LLM, which we prompt to elicit partitioning criteria. The criteria generated for each subset are accumulated in $\widetilde{\mathcal{R}}$. Experiments in Sec. 4.1 show that the Caption-based Proposer is the most effective; so we use it as our main method and consider the others as baselines.
Criteria refinement: Since the Proposers operate on subsets, the criteria accumulated in $\widetilde{\mathcal{R}}$ may include redundant or noisy entries. To address this, we input all initially proposed criteria from $\widetilde{\mathcal{R}}$ into a LLM (Meta, 2024b), prompting it to consolidate similar criteria and discard noisy ones. This step refines and updates the criteria pool into $\mathcal{R}$, which is ready to be used in the subsequent stage.
3.2 Semantic Grouper
The automatically discovered criteria $\mathcal{R}$ serve as thematic indicators for revealing the semantic substructures (or clusters) within the image set $\mathcal{D}$. To uncover these substructures, as shown in Fig. 3(right), the Grouper takes $\mathcal{D}$ and each criterion $R_{l}$ as inputs. It then discovers cluster names $\mathcal{C}_{l}=\{s_{k}^{l}\}_{k=1}^{K_{l}}$ conditioned on $R_{l}$ and assigns each image to its corresponding cluster. The core design of the Grouper focuses on aligning semantic substructure discovery with a specified partitioning criterion. Similar to the Proposer, we also explore three distinct approaches for the Grouper.
Furthermore, as clusters under a given criterion can be formed at varying semantic granularities, depending on user preferences, we have designed our Grouper to cluster $\mathcal{D}$ at three levels of granularity: coarse, middle, and fine-grained. This design enables TeDeSC to provide new insights into the data at different granularities. For example, for the criterion food Cuisine, TeDeSC can organize images at a coarse-grained continental level (e.g., “European” or “Asian”), a middle-grained regional level (e.g., “Mediterranean” or “Southeast Asian”), or a fine-grained national level (e.g., “Italian” or “Thai”).
Image-based Grouper: Given a target criterion $R_{l}\in\mathcal{R}$, we first prompt a LLM to generate a question $q_{l}$ specific to $R_{l}$-i.e., for criterion Mood the generated question is: “What mood is conveyed by this image? Answer with an abstract, common, and specific category name, respectively.” We then use $q_{l}$ to guide a visual question answering (VQA) model to infer semantic cluster names for each image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$. Consequently, by aggregating the cluster assignment results at each granularity level across $\mathcal{D}$, we can derive multi-granularity semantic substructures.
Tag-based Grouper: Given a criterion, we prompt a LLM to generate a list of common, middle-grained tags specific to that criterion (e.g., “Recreational facility”). Following Liu et al. (2024d), we then obtain coarse- and fine-grained tags by querying the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag. Unlike lexical databases such as WordNet or ConceptNet (Speer et al., 2017), which do not support free-form input and may be limited in accommodating certain criteria, tag synthesis using a LLM provides a flexible and reliable alternative. Subsequently, we use an image tagger to assign the most relevant tag from the candidate tags at each granularity to each image, resulting in multi-granularity substructures after aggregation.
Caption-based Grouper (Main): We prompt a MLLM to generate captions that specifically focus on the target criterion for each image, as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$. Next, we use a LLM in a three-step process to assign images to clusters at multiple granularity levels: i) Initial Naming: First, we prompt the LLM to assign a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, resulting in an initial set of names $\mathcal{S}_{\text{init}}^{l}$ of size $N$; ii) Multi-granularity Cluster Refinement: Using these initial names as basis, we prompt the LLM to refine them into three structured granularity levels: $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$. These structured names serve as candidates for cluster assignment; iii) Final Assignment: Each image is then assigned a class name from each granularity level based on its caption, as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$. Experiments in Sec. 4.2 show that the Caption-based Grouper performs the best; thus we use it as our main method.
4 Experiments
[…]
Evaluation metrics for Criteria Discovery: Given the discovered criteria set $\mathcal{R}$ of the proposer, we assess its quality from two dimensions: i) Comprehensiveness: We compute True Positive Rate (TPR) (Csurka et al., 2024) for the predicted criteria relative to the annotated ones as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, to assess to what extent the predicted set covers the ground-truth set $\mathcal{Y}$; ii) Diversity: We compute the pairwise semantic similarity between the predicted criteria $\mathcal{R}$, using Sentence-BERT (Reimers & Gurevych, 2019), and convert it into a diversity measure by subtracting the similarity from 1. The final score is the average of these pairwise values, reflecting how well the proposed criteria avoid redundancy and capture distinct aspects of the data—i.e., criteria such as Location and Place are nearly identical and provide little additional insight. A higher score indicates better diversity. It is important to note that the number of valid grouping criteria is subjective and potentially unlimited, making False Positives difficult to define. Thus, we use TPR as the primary evaluation metric.
Evaluation metrics for substructure uncovering: Following Conti et al. (2023) and Liu et al. (2024e), for each criterion and its substructure uncovered by the grouper, we evaluate its alignment with the substructure defined by the ground-truth labels along two dimensions: i) Semantic Consistency: For each image, we measure the semantic similarity between its assigned cluster name and the ground-truth label under the current criterion using Sentence-BERT. The average similarity across the dataset is reported as Semantic Accuracy (SAcc), reflecting how well the predicted substructure semantically aligns with the ground-truth. ii) Structural Consistency: We compute the clustering accuracy (CAcc) between the predicted and ground-truth substructures (clusterings) to evaluate how accurately the predicted structure groups images. By combining CAcc and SAcc, we evaluate the overall alignment of the predicted substructure with the ground-truth across various criteria.
Evaluation strategy: Since the granularity of the ground-truth annotations is unknown during evaluation, we choose the substructure from the three predicted granularities that achieves the highest CAcc relative to the ground-truth for assessing our TeDeSC. Unlike methods (Kwon et al., 2024; Yao et al., 2024), which rely on the ground-truth number of clusters to create flat clusterings that perfectly match the ground-truth granularity, we believe our approach offers a fair and practical evaluation strategy for systems capable of providing multi-granularity output.
Figure 4: Comparison of Criteria Proposers. (a) Comprehensiveness: TPR performance of each proposer evaluated on Basic and Hard ground-truth criteria. (b) Diversity: Diversity scores of the proposed criteria for each proposer. (c) Influence of Image Quantity: TPR performance of the caption-based proposer on Hard ground-truth criteria across varying image scales used for criteria discovery. See expanded results in App. E.1.
4.1 Study of the Criteria Proposer
Expanding ground-truth criteria for comprehensive evaluation: The newly introduced benchmarks provide up to four ground-truth criteria to evaluate the predicted criteria, but for image collections with high scene complexity (e.g.,, COCO-4c), these are still insufficient to cover all valid grouping options. To address this, we expanded the ground-truth criteria for each of the six benchmarks in Sec. 2 by using human annotators to identify additional criteria. This resulted in expanded sets that have: {10, 4, 11, 7, 17, 11} distinct criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}. We refer to the per-image annotated criteria set (see Fig. 2) as Basic ground truth and the expanded set as Hard. See App. B.2 for the exact criteria and annotation details.
Which Criteria Proposer is best? Our comparative results, presented in Fig. 4(a-b), show that the Caption-based Criteria Proposer consistently outperforms its counterparts by a significant margin across all six benchmarks. As shown in Fig. 4(a), the caption-based proposer achieves an average of 100% TPR on the Basic criteria set and 75.1% TPR on the Hard set across the benchmarks. Its superior performance is particularly evident under the Hard criteria, where it surpasses the second-best tag-based proposer by +32.2% TPR. Meanwhile, the criteria predicted by the caption-based proposer also show greater diversity in Fig. 4(b), coming closest to the diversity of human-annotated criteria.
Although the tag-based proposer performs well on object-centric benchmarks like Clevr-4c or Food-4c, it significantly lags behind the caption-based proposer on more challenging benchmarks such as COCO-4c and Action-3c, which feature more complex, realistic images. This discrepancy arises because image tagging is a more object-centric way of representing image semantics, potentially missing the rich context depicted in images. Conversely, while the image-based proposer outperforms the tag-based proposer on COCO-4c and Action-3c, it is less effective overall because it is limited to reasoning over a small subset of images and loses visual details when combining images into a grid.
Studying the influence of image quantity on Criteria Discovery: Fig. 4(c) shows the TPR performance of the caption-based proposer across varying image scales (100% to only 1 image) used for criteria proposing, tested on the Hard criteria sets of six benchmarks. Interestingly, satisfactory performance is achieved with just a few images in object-centric benchmarks like Card-2c and Clevr-4c. In fact, even a single image often provides sufficient information for reasonable criteria discovery in these object-centric datasets, where a brief glimpse often represents the entirety adequately. For example, seeing one playing card allows the proposer to easily suggest criteria like “Rank” and “Suit” in the Card-2c dataset. However, this does not hold for more complex datasets like COCO-4c, Food-4c, and Action-3c, which contain diverse and realistic scenarios. Here, a reduction in image scale leads to a clear drop in TPR performance, as these datasets require a larger set of images to capture their intricate and varied thematic criteria. Since TeDeSC operates without prior knowledge of the dataset, we default to using the entire dataset to ensure comprehensive criteria discovery.
Figure 5: Comparison of the Semantic Groupers. We show CAcc, SAcc, their Harmonic Mean (HM) evaluated on the basic criteria. The Caption-based Proposer is used by all groupers. See expanded results in App. E.2.
4.2 Study of the Semantic Grouper
Which Semantic Grouper is best? To determine which grouper best uncovers the substructures indicated by the discovered criteria and aligns most closely with the ground truth, we use the Harmonic Mean (HM) of the CAcc and SAcc for each criterion. This allows us to identify the best overall performer for each criterion. In the comparison, to further illustrate the relative performance position of our framework, we first establish a pseudo upper-bound reference by leveraging the zero-shot classification capabilities of CLIP ViT-L14 (Radford et al., 2021) with ground-truth labels for each criterion, where the grouping criteria, cluster names, and number of clusters are all known. Additionally, we apply KMeans (Ahmed et al., 2020) with the ground-truth number of clusters, using robust visual features from CLIP ViT-L14, DINOv1-B/16 (Li et al., 2022), and DINOv2-G/14 (Oquab et al., 2023) to establish baseline positions for clustering performance.
Our comparative analysis, presented in Fig. 5, shows that the Caption-based Semantic Grouper performs best, ranking first in 12 out of 19 tested criteria based on the HM, across six benchmarks. Specifically, in terms of CAcc, the caption-based grouper achieves an average of 59.9% across all tested criteria, closely aligning with the pseudo-upper-bound average of 58.1% and significantly outperforming feature representation-driven baselines, with the highest being DINOv2 at 40.6% CAcc. This demonstrates the effectiveness of our text-driven approach, as the cluster names inferred by the caption-based provide robust guidance for grouping. For SAcc, while the caption-based grouper falls short of the upper-bound reference average of 74.2%, where exact ground-truth class names are used, it still achieves a notable average of 60.5% across all tested criteria, surpassing its counterparts. To some extent, this gap in semantic consistency is expected due to the open nature of the semantic space—i.e., names like “Joyful”, “Happy”, and “Cheerful” are often used to describe the same Mood, yet they do not exhibit full semantic equivalence. The BLIP-2 image-based grouper also shows commendable performance, ranking first in 5 of the 19 tested criteria. This suggests that the criterion-customized prompts generated by LLM help BLIP-2 accurately label relevant visual content. However, its per-image labeling approach may lead to noisy clusters due to a lack of overall alignment between them. The tag-based grouper lags behind in all benchmarks likely due to a mismatch between the generated tags and the concepts in the dataset.
Table 2: Comparison with criterion-conditioned clustering methods. For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria. We provide CLIP ViT-L/14 zero-shot performance as the pseudo upper-bound reference (UB). See expanded per-criterion results in App. E.3.
COCO-4c Food-4c Clevr-4c Action-3c Card-2c Fruit-2c Avg
CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
UB 40.1 60.6 64.1 80.2 56.7 72.5 79.8 82.3 41.4 66.9 69.4 88.3 50.2 64.4
[…]
MMaP 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 41.3 -
TeDeSC (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 53.0 53.4
[…]
Figure 6: Ablation study of multi-granularity refinement. For each benchmark, we show average CAcc and SAcc across all criteria. See expanded per-criterion numerical results in App. E.4.
Necessity of multi-granularity cluster refinement: To validate the effectiveness of the multi-granularity cluster refinement design, we design control experiments with our Caption-based Grouper, using three different methods for constructing cluster names to organize images: i) Initial Names: using the initially assigned names as the final output, ii) Flat Refinement: prompting the LLM to refine the initial names into a flat list with a unified granularity, and our iii) Multi-granularity Refinement. We then compare the performance of these approaches in Fig. 6. We observe that both refinement methods significantly improve clustering accuracy compared to using noisy initial names, highlighting the importance of having granularity-consistent cluster names for accurately revealing substructures. Additionally, our proposed multi-granularity method surpasses flat refinement, as it enables the grouper to generate clustering results at different granularities, providing greater flexibility in aligning with the label granularity-the user-preferred level of detail.
Additional studies: In App. F, we provide further studies on: i) Fine-grained Scenario: We examine how to integrate our framework with more advanced cross-modal Chain-of-Thought prompting strategies to better handle fine-grained criteria. ii) Sensitivity Analysis: We analyze the system sensitivity to various MLLMs and LLMs. Additionally, Sec. G offers a Qualitative Analysis by visualizing clustering results for different criteria, and Sec. H includes a Failure Case Analysis.
5 Applications
Discovering and mitigating dataset bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024).
Figure 7: Results of dataset bias discovery and mitigation. Worst group and average accuracies(%) are reported.
As a case study, we applied the proposed TeDeSC framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split. Additional details are provided in App. I.1.
Findings: As expected, our method identified the grouping criteria Hair color and Gender. Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters. As shown in Fig. 7(a), we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%). Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”. To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels. As shown in Fig. 7(b), our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
Figure 8: Bias discovery results in T2I models. Bias evaluation results are shown for the associated occupations.
Discovering novel bias in Text-to-Image Diffusion Models: Stereotypical biases related to gender or race (Naik & Nushi, 2023) in images generated by widely accessible text-to-image (T2I) models like Stable Diffusion XL (SDXL) (Podell et al., 2024) and DALL·E3 (Betker et al., 2023) have sparked intense discussions (Nicoletti & Bass, 2023) and significant research (Bianchi et al., 2023). However, beyond these well-known biases extensively studied in social science, we wonder: Are there other biases present in T2I-generated images? To explore this question, we selected nine occupations (e.g.,, Nurse or CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) as subjects for our case study on bias. For each occupation, we generated 100 images using the prompt “A portrait photo of a $<$OCCUPATION$>$” with both DALL-E3 and SDXL, resulting in a total of 1.8k images (see Fig. 20 and Fig. 21 for examples). We applied TeDeSC to these images and automatically discovered 10 grouping criteria (possible bias dimensions) along with their predicted distributions for each occupation image set. To inspect the biases, we calculated the normalized entropy of each distribution to quantify the bias intensity, as proposed by D’Incà et al. (2024), and identified the dominant cluster—the one containing the most images—as the possible bias direction. Furthermore, we conducted a user study with 54 anonymous participants to assess how well our bias discovery results align with human judgment. We found that our system’s bias intensity scores closely matched human ratings—with an Absolute Mean Error (AME) of 0.1396 on a scale from 0 to 1—and that its predicted bias directions agreed with human evaluations 72.3% of the time. We highlight some key findings below and provide complete results and evaluations in App. I.2.
Findings: In Fig. 8, we present some of our notable findings. Without any predefined biases, our method successfully identifies well-known social biases associated with certain occupations. As shown in Fig. 8(a–c), the SDXL-generated images exhibit near-total social bias for occupations like Nurse, Firefighter, and Basketball Player, displaying gender and racial imbalances that exceed official statistics such as those from U.S. Bureau of Labor Statistics (2021). In contrast, DALL·E3 shows markedly improved performance in mitigating social biases—likely due to its bias mitigation “guardrails” (OpenAI, 2022). More importantly, as shown in Fig. 8(d–f), our method revealed previously unnoticed novel bias dimensions where both models exhibit moderate biases. For example, we found that SDXL strongly associates CEOs with “Grey” hair color, while DALL·E3 favors “Dark” hair. Interestingly, and contrary to trends observed in well-known social biases, DALL·E3 shows stronger biases than SDXL in Hair style and Grooming for Nurse and Teacher. For instance, 91% of the Teacher images generated by DALL·E3 depict individuals who are clean-shaven (Fig. 8(f)). The gap in DALL·E3’s performance between addressing well-known and novel biases suggests that even industrial T2I models with guardrail systems may overlook some novel biases, underscoring the importance of thoroughly examining biases beyond the commonly recognized ones.
Figure 9: Analysis of social media photo popularity on SPID dataset. We show the viral and major popular clusters along with the popularity distribution of data points within these clusters across five criteria (in Grey).
Analyzing Social media image popularity: Given a set of social media photos with significantly varying popularity for each photo, it is natural to ask: What visual element makes a photo popular? As a case study, we applied TeDeSC to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), each accompanied by a popularity score based on the number of views. After our method grouped the photos into semantic clusters based on 10 discovered criteria, we calculated two scores for each cluster: i) $\bar{P}$: the average popularity of photos within the cluster, and ii) $\bar{P}_{w}$: the average popularity $\bar{P}$ weighted by the cluster’s proportion relative to the entire dataset. Here, we consider the cluster with the highest $\bar{P}$ as viral, meaning its photos are highly popular but may be few in number, and the cluster with the highest $\bar{P}_{w}$ as major popular, where the majority of popular photos are concentrated. We discuss the key findings below and provide complete results in App. I.3.
Findings: As shown in our analysis in Fig. 9, by combining the interpretable groupings from our method with popularity scores, the visual elements that drive virality (clusters with the highest $\bar{P}$) and the common traits shared among widely popular images (clusters with the highest $\bar{P}_{w}$) across different dimensions clearly emerged. Interestingly, we further observe that viral visual elements often exhibit sharp semantic contrast with those of the majority of popular images, such as Musical activities vs. Rest and relaxation, and High-intensity expressions vs. Neutral emotion. This suggests that viral content is more attention-grabbing, likely due to novel, unexpected, or intense visuals that stand out, especially in an era where the average attention span is less than 9 seconds (McSpadden, 2015; Farid, 2024). During our analysis, as shown in Fig. 9, we unexpectedly found that some of the most popular images in certain clusters contained not safe for work (NSFW) content, which had not been previously identified in the SPID dataset. This suggests that some images gain popularity due to negative or provocative visuals. It also highlights the need for more thorough inspection of visual datasets to avoid harm, and our proposed framework can be a useful tool for this purpose.
6 Conclusion
In this work, we introduce the task of semantic multiple clustering and propose TeDeSC, a system that automatically discovers grouping criteria in natural language from large image collections and uncovers interpretable data substructures based on these criteria. We rigorously evaluate various design choices of TeDeSC on four existing and two newly proposed benchmarks, and demonstrate its ability to reveal valuable insights that might otherwise go unnoticed in various real-world applications.
Acknowledgments
This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGenerationEU. E.R. acknowledges partial support from the PRECRISIS project, funded by the EU Internal Security Fund (ISFP-2022-TFI-AG-PROTECT-02-101100539), the EU project SPRING (No. 871245), and the PRIN project LEGO-AI (Prot. 2020TA3K9N). This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN. We extend our gratitude to Feng Xue for his valuable suggestions on plot creation. M.L. would like to thank Margherita Potrich for her unwavering support.
Ethics Statement
We do not anticipate any immediate negative societal impacts from our work. However, we encourage future researchers building on this work to remain vigilant, as we have, about the potential for TeDeSC, which integrates LLMs and MLLMs–particularly their human-like reasoning abilities– to be used both for good and for harm.
The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs, as well as our exploration of social media image popularity, is to reveal and address these biases and the presence of sensitive or not-safe-for-work (NSFW) content that objectively exist in the datasets and models. We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content. Specifically, in Sec. 5, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016). Furthermore, our framework reveals previously undisclosed sensitive and sexual content in the SPID dataset (Ortis et al., 2019). We responsibly present these findings in Sec. 5, applying significant blurring to disturbing content, with the intention of raising community awareness about the need to further scrutinize NSFW content in existing benchmarks. However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups. We strongly oppose any such misuse or misrepresentation of our work. Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
For the creation of our two new benchmarks, COCO-4c and Food-4c, we sourced images exclusively from the COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) datasets, strictly adhering to their licensing agreements. Additionally, we utilized voluntary human annotators for proposing valid grouping criteria and creating annotations along these criteria, rather than employing annotators from crowdsourcing platforms. This decision was made to ensure sustainability, fair compensation, and high-quality work, as well as to safeguard the psychological well-being of participants. Similarly, for our user study on T2I model bias evaluation, we recruited voluntary participants via questionnaires to collect human evaluation results. The user study was conducted entirely anonymously, with participants providing informed consent. Our project, including data annotation and the user study involving human subjects, was approved by the Ethical Review Board of our university.
Lastly, we emphasize that our proposed framework, TeDeSC, relies on open-source LLMs and MLLMs, allowing full deployment on local machines. We refrain from using APIs from industrial LLMs or MLLMs, both to ensure reproducibility and to protect data privacy.
Reproducibility Statement
We will release all essential resources required to reproduce the experimental results presented in this project, including source code, exact prompts, benchmarks with their data splits, and generated images, upon publication. Our proposed framework, TeDeSC, is built on open-source, publicly accessible models to ensure reproducibility. In Sec. 3, we provide a detailed description of how our framework is constructed. Additionally, App. C contains further implementation details, including exact prompts, to help practitioners easily reproduce our method. Details regarding the implementation of the compared methods are also provided in App. D. Moreover, we present extended numerical experimental results in App. E, alongside comprehensive findings for the application study in App. I. We believe that the thorough descriptions of our methodology, the extensive presentation of experimental results, and the open-source nature of our framework ensure that this work is highly reproducible, enabling future researchers and practitioners to readily apply our method to various domains.
Appendix
[…]
This appendix provides detailed supplementary information supporting the implementations, experiments, and findings presented in the paper. First, in App. A, we offer an extensive discussion of Related Work, covering tasks and methods pertinent to our study. In App. B, we describe the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, along with the process for creating hard ground-truth criteria for evaluating different proposers. App. C outlines the prompts and implementation details for our proposed framework, including both Criteria Proposers and Semantic Groupers. App. D provides a deep dive into the implementation specifics of the compared methods used in our experiments. In App. E, we present additional quantitative results that supplement the findings reported in the main paper, covering evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods. App. F extends our analysis of the proposed framework, including system sensitivity and studies on fine-grained image collections. App. G presents further qualitative results for the predicted clusters, while App. H offers a more detailed analysis of the failure cases encountered by our method. App. I provides additional findings, along with implementation details and user study results for the three applications explored in this work. Finally, App. J presents a deeper discussion of related settings pertinent to our research.
We will release all essential resources for reproducing this work, including code, prompts, benchmarks, and annotations, upon publication.
Appendix A Related Work
Image Clustering. Image clustering (Xu & Wunsch, 2005) discovers hidden grouping structures within large, unstructured, and unlabeled image collections, serving as a tool for various data-driven applications (Wazarkar & Keshavamurthy, 2018). To achieve this, deep clustering (DC) methods such as DEC (Xie et al., 2016) and SCAN (Van Gansbeke et al., 2020) focus on simultaneously learning feature representations and cluster assignments using deep neural networks via self-supervised techniques (Caron et al., 2020; Zhong et al., 2021; Ren et al., 2024). Furthermore, large-scale pre-trained feature representations like DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and CLIP (Radford et al., 2021) have also been shown to be effective at clustering image collections in a zero-shot fashion with the help of KMeans (Vaze et al., 2022; Liu et al., 2024f; Han et al., 2023).
However, it is well-known that “clusters are in the eye of the beholder” (Estivill-Castro, 2002); there often exist multiple ways to partition the same image collection into clusters, and what constitutes a cluster depends on the user’s needs. This insight has led to the study of Multiple Clustering (MC) (Ren et al., 2022; Yao et al., 2023; Yu et al., 2024), which aims to simultaneously learn feature representations and cluster assignments from different perspectives to find various ways of grouping the same data, enabling alternative interpretations from different viewpoints. Although existing MC methods have achieved impressive results (Yao et al., 2023) on some benchmarks, they share similar limitations with deep clustering approaches (Xie et al., 2016; Van Gansbeke et al., 2020). Their results require intensive manual post-analysis, and they also hold strong assumptions: users must specify (i) the number of potential clusterings and (ii) the number of clusters within each clustering. However, when dealing with millions of unstructured images, it is infeasible for users—who are trying to understand the data—to know this information a priori.
[…]
In stark contrast to prior work, we introduce and study the task of semantic multiple clustering (SMC). Instead of requiring users to specify the grouping criteria, SMC seeks to actively and automatically discover criteria expressed in natural language from large visual data and uncover the corresponding semantic substructures, without access to any of the aforementioned human priors. As demonstrated in Sec. 5, the flexibility provided by our proposed SMC framework adds significant value to various data-driven applications, unveiling novel insights about the data that might not have been noticed before.
Topic Discovery. The setting of semantic multiple clustering (SMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund & Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g.,, documents). Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data. In contrast, our work focuses on discovering thematic criteria from large visual content. However, indeed, the core challenges of SMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data. Nevertheless, SMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them. Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024b).
Multimodal Large Language Model. Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data. The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024a; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024b) alongside a pre-trained vision encoder (Radford et al., 2021). A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs. Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2023), and the LLaVA series (Liu et al., 2024c; b; Li et al., 2024). In our proposed TeDeSC framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria. Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2024) to establish a baseline image-based proposer for the SMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2023) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
[…]
Text-Driven Image Retrieval. Given a query text (e.g.,, “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query. In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques. In this context, the query can be considered as a sort of “cluster name”. However, this differs significantly from the proposed task of semantic multiple clustering (SMC), because SMC requires both discovering the textual criteria and the corresponding textual clusters. Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish SMC.
Appendix B Benchmark Details
B.1 Benchmark Construction of COCO-4c and Food-4c
Table 3: Summary of number of classes for the basic criteria annotation across the six benchmarks.
Dataset Basic Criterion Number of Classes
COCO-4c Activity 64
Location 19
Mood 20
Time of Day 6
Food-4c Food Type 101
Cuisine 15
Course 5
Diet 4
Action-3c Action 40
Location 10
Mood 4
Clevr-4c Color 10
Texture 10
Shape 10
Count 10
Card-2c Rank 14
Suit 5
Fruit-2c Species 34
Color 15
To create high-quality benchmarks for COCO-4c and Food-4c, we designed a four-step annotation pipeline:
(1) Criteria Identification: We first split COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images into batches of 100. Each batch was stitched into a 10$\times$10 grid to form a single image. These grid images were then distributed to 5 human annotators, who were tasked with identifying grouping criteria. For each dataset, we selected the 4 most frequently occurring criteria, as shown in Tab. 3, to proceed with per-image annotation.
(2) Label Candidate Generation: To facilitate the annotation process, we used GPT-4V (OpenAI, 2023) to generate an initial list of candidate labels for each criterion. Specifically, for each criterion of COCO-4c and Food-4c, GPT-4V was prompted to assign a label that reflected the criterion for each image. This resulted in a list of criterion-specific label candidates for each dataset.
(3) Image Annotation: Next, 10 human annotators were tasked with assigning a label from the criterion-specific candidates to each image in COCO-4c and Food-4c for each criterion. The entire annotation process took 25 days to complete.
[…]
Following these steps, we constructed COCO-4c and Food-4c. Note that we used the official COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images for our benchmarks and did not collect any new images. We adhered strictly to the licenses of the datasets during their creation. The exact number of classes is presented in Tab. 3. Additionally, the annotated class names for each criterion of COCO-4c are provided in Tab. 4, and for Food-4c in Tab. 5.
Table 4: Full class names for COCO-4c across the four basic criteria.
Criterion COCO-4c
Activity “repairing a toilet”, “playing volleyball”, “playing guitar”, “haircutting”, “cutting a cigar”, “kayaking”, “applauding”, “tying a tie”, “playing basketball”, “washing dishes”, “gardening”, “texting messages”, “repairing a car”, “peeing”, “cleaning the floor”, “writing on a book”, “feeding a horse”, “singing”, “baking”, “hiking”, “smoking”, “riding an elephant”, “pouring liquid”, “waving hands”, “swimming”, “meditating”, “fixing a bike”, “cutting vegetables”, “walking a dog”, “reading a book”, “celebrating”, “queuing”, “cutting a cake”, “brushing teeth”, “playing soccer”, “jumping”, “snowboarding”, “playing”, “touching animals”, “pushing a cart”, “watching tv”, “rowing a boat”, “taking photos”, “running”, “flying a kite”, “riding a horse”, “playing video games”, “holding up an umbrella”, “throwing a frisbee”, “lying down”, “riding a bike”, “drinking”, “cooking”, “phoning”, “chatting”, “skiing”, “driving”, “surfing”, “skateboarding”, “playing baseball”, “playing tennis”, “using a computer”, “posing”, “eating”
Location “amusement or theme park”, “healthcare facility”, “virtual or digital space”, “educational institution”, “industrial area”, “historical landmark”, “public event or gathering”, “store or market”, “underground or enclosed space”, “transportation hub”, “zoo”, “water body”, “office or workplace”, “park or recreational area”, “restaurant or dining area”, “sports facility”, “natural environment”, “urban area or city street”, “residential area”
Mood “anxious”, “sombre”, “contemplative”, “suspenseful”, “serene”, “nostalgic”, “inspired”, “whimsical”, “romantic”, “mysterious”, “melancholic”, “chaotic”, “humorous”, “vibrant”, “peaceful”, “energetic”, “focused”, “joyful”, “relaxed”, “adventurous”
Time of Day “evening”, “afternoon”, “night”, “morning”, “indoor lighting”, “midday”
Table 5: Full class names for Food-4c across the four basic criteria.
Criterion Food-4c
Food Type “apple pie”, “baby back ribs”, “baklava”, “beef carpaccio”, “beef tartare”, “beet salad”, “beignets”, “bibimbap”, “bread pudding”, “breakfast burrito”, “bruschetta”, “caesar salad”, “cannoli”, “caprese salad”, “carrot cake”, “ceviche”, “cheesecake”, “cheese plate”, “chicken curry”, “chicken quesadilla”, “chicken wings”, “chocolate cake”, “chocolate mousse”, “churros”, “clam chowder”, “club sandwich”, “crab cakes”, “creme brulee”, “croque madame”, “cup cakes”, “deviled eggs”, “donuts”, “dumplings”, “edamame”, “eggs benedict”, “escargots”, “falafel”, “filet mignon”, “fish and chips”, “foie gras”, “french fries”, “french onion soup”, “french toast”, “fried calamari”, “fried rice”, “frozen yogurt”, “garlic bread”, “gnocchi”, “greek salad”, “grilled cheese sandwich”, “grilled salmon”, “guacamole”, “gyoza”, “hamburger”, “hot and sour soup”, “hot dog”, “huevos rancheros”, “hummus”, “ice cream”, “lasagna”, “lobster bisque”, “lobster roll sandwich”, “macaroni and cheese”, “macarons”, “miso soup”, “mussels”, “nachos”, “omelette”, “onion rings”, “oysters”, “pad thai”, “paella”, “pancakes”, “panna cotta”, “peking duck”, “pho”, “pizza”, “pork chop”, “poutine”, “prime rib”, “pulled pork sandwich”, “ramen”, “ravioli”, “red velvet cake”, “risotto”, “samosa”, “sashimi”, “scallops”, “seaweed salad”, “shrimp and grits”, “spaghetti bolognese”, “spaghetti carbonara”, “spring rolls”, “steak”, “strawberry shortcake”, “sushi”, “tacos”, “takoyaki”, “tiramisu”, “tuna tartare”, “waffles”
Cuisine “japanese”, “indian”, “american”, “greek”, “spanish”, “mexican”, “italian”, “vietnamese”, “canadian”, “korean”, “chinese”, “middle eastern”, “french”, “thai”, “general”
Course “appetizer”, “main course”, “side dish”, “dessert”, “breakfast”
Diet “omnivore”, “vegan”, “vegetarian”, “gluten free”
B.2 Details on Hard Grouping Criteria Annotation
Table 6: Annotated criteria for the six benchmarks. The basic criteria are annotated on per-image level for each benchmark, while the hard criteria (those not in the basic criteria) are further exhaustively annotated by human annotators for further evaluating the performance of the rule proposer in SMC task.
COCO-4c Food-4c Action-3c
Basic criteria Hard criteria Basic criteria Hard criteria Basic criteria Hard criteria
Total: 4 Total: 17 Total: 4 Total: 11 Total: 3 Total: 11
Activity Activity Food Type Food Type Action Action
Location Location Cuisine Cuisine Mood Mood
Mood Mood Course Course Location Location
Time of Day Time of Day Diet Diet Clothing Style
Interaction Tableware Type Number of People Present
Number of People Present Presentation Style Age or Age Composition
Group Dynamics Color Palette Race or Race Composition
Clothing Style Setting/Theme Occasion or Event Type
Occasion or Event Type Primary Taste Group Dynamics
Photo Style Primary Ingredient Lighting Condition
Type of Animal Present Cooking Method Gender or Gender Composition
Weather
Type of Primary Object
Continent
Age or Age Composition
Race or Race Composition
Gender or Gender Composition
       Clevr-4c Card-2c Fruit-2c
Basic criteria Hard criteria Basic criteria Hard criteria Basic criteria Hard criteria
Total: 4 Total: 7 Total: 2 Total: 4 Total: 2 Total: 8
  Color Color Rank Rank Species Species
Texture Texture Suit Suit Color Color
Shape Shape Color Size
Count Count Illustration Style Seasonality
Spatial Positioning Primary Taste
Count of Surface Texture
Complexity of Geometry Ripeness
Fruit Quantity and Arrangement
In Tab. 6, we present the additional annotated Hard grouping criteria ground truth alongside the Basic criteria for each benchmark.
While we have established more rigorous and challenging benchmarks such as COCO-4c and Food-4c, which feature up to four distinct grouping criteria, these annotated criteria sets do not encompass all potential grouping criteria within the image collections. This is particularly true for more complex and realistic datasets like COCO-4c, Food-4c, and Action-3c. As a result, the performance differences between different criteria proposers on these basic criteria, as shown in Fig. 4, tend to be close to each other, limiting our understanding of each proposer’s ability to generate comprehensive grouping criteria.
To address this limitation, we employed human annotators to further identify and propose grouping criteria across the six benchmarks, resulting in a more extensive ground-truth set for each benchmark. This provides a better basis for evaluating the comprehensiveness of the different proposers. We refer to this set of larger annotation criteria as the Hard criteria, in contrast to the Basic criteria, which involve per-image annotations. Note that for the Hard criteria, per-image label annotation is not provided due to the high cost of annotation. The procedure for obtaining the Hard grouping criteria is as follows:
(1) Criteria Discovery: We divided each dataset into batches of 100 images, displaying each batch in a 10$\times$10 grid. Five human annotators were assigned to each batch and instructed to identify as many valid grouping criteria as possible. The proposed criteria from each annotator were then combined to form a comprehensive set of grouping criteria for the dataset.
(2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g.,, Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in Tab. 6. This resulted in sets containing 8 criteria for Fruit-2c, 4 criteria for card, 11 criteria for Action-3c, 7 criteria for Clevr-4c, 17 criteria for COCO-4c, and 11 criteria for Food-4c. These expanded ground-truth sets enable us to more effectively evaluate the capabilities of various criteria discovery methods, providing a clearer understanding of different criteria proposers.
Appendix C Further Implementation Details
In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App. C.1 and the Semantic Grouper in App. C.2.
C.1 Prompts and Implementation Details of Criteria Proposer
Image-based Proposer:
Table 7: Prompts for the MLLM in the image-based proposer for criteria proposing.
  Prompt purpose Prompt
  System Prompt You are a helpful AI assistant
Input Explanation This image contains 64 individual images arranged in 8 columns and 8 rows.
Goal Explanation I am a machine learning researcher trying to identify all the possible clustering criteria or rules that could be used to group these images so I can better understand my data.
Task Instruction Your job is to carefully analyze the entire set of the 64 images, and identify five distinct clustering criteria or rules that could be used to cluster or group these images. Please consider different characteristics.
Output Instruction Please write a list of the 5 identified clustering criteria or rules (separated by bullet points “*”).
Task Reinforcement Again, I want to identify all the possible clustering criteria or rules that could be used to group these images. List the 5 distinct clustering criteria or rules that you identified from the 64 images. Answer with a list (separated by bullet points “*”). Your response:
In Tab. 7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2024). Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images. Each subset is then stitched into an $8\times 8$ image grid, treated as a single image, and fed into the MLLM. For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab. 7. After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set. Finally, we deduplicate the discovered criteria and accumulate them into the criteria pool.
Tag-based Proposer:
Table 8: Prompts for the LLM used in the tag-based proposer for criteria proposing. We embed the exact image captions by replacing the placeholders "{TAGS}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation The following are the tagging results of a set of images in the format of “Image ID: tag 1, tag 2, …, tag 10”. These assigned tags reflect the visible semantic content of each image:
Tag Embedding Image 1: "{TAGS}" Image 2: "{TAGS}" Image N: "{TAGS}"
Goal Explanation I am a machine learning researcher trying to figure out the potential clustering or grouping criteria that exist in these images. So I can better understand my data and group them into different clusters based on different criteria.
Task Instruction Please analyze these images by using their assigned tags. Come up with an array of distinct clustering criteria that exist in this set of images.
Output Instruction Please write a list of clustering criteria (separated by bullet points “*”).
Task Reinforcement Again, I want to figure out what are the potential clustering or grouping criteria that I can use to group these images into different clusters. List an array of clustering or grouping criteria that often exist in this set of images based on the tagging results. Answer with a list (separated by bullet points “*”). Your response:
In Tab. 8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024b). For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image. These tags are selected from the WordNet (Miller, 1995) vocabulary, which contains 118k English synsets, and represent the semantic content of the images. We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging. Next, we embed the assigned tags into the prompt shown in Tab. 8 to carry the semantics of the entire image set and query the LLM to propose grouping criteria. The criteria proposed by the LLM are then added to the criteria pool. Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g.,, 128k for Llama-3.1-8B) for the datasets used in our experiments. However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
Caption-based Proposer:
Table 9: Prompts for the MLLM in the caption-based proposer for generating detailed descriptions of the images.
  Prompt purpose Prompt
  System Prompt You are a helpful AI assistant
Task Instruction Describe the following image in detail.
Table 10: Prompts for the LLM used in the caption-based proposer for criteria proposing. We embed the exact image captions by replacing the placeholders "{CAPTION}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation The following are the result of captioning a set of images:
Caption Embedding Image 1: "{CAPTION}" Image 2: "{CAPTION}" Image N: "{CAPTION}"
Goal Explanation I am a machine learning researcher trying to figure out the potential clustering or grouping criteria that exist in these images. So I can better understand my data and group them into different clusters based on different criteria.
Task Instruction Come up with ten distinct clustering criteria that exist in this set of images.
Output Instruction Please write a list of clustering criteria (separated by bullet points “*”).
Task Reinforcement Again I want to figure out what are the potential clustering/grouping criteria that I can use to group these images into different clusters. List ten clustering or grouping criteria that often exist in this set of images based on the captioning results. Answer with a list (separated by bullet points “*”). Your response:
We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab. 9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024b) in Tab. 10. Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language. The generated captions are then randomly shuffled and split into disjoint subsets, each containing 400 captions. Next, we embed the captions from each subset into the prompt shown in Tab. 10 and use it to query the LLM to propose grouping criteria for the images represented by the captions. After iterating through all subsets, we take the union of the proposed criteria across subsets as the discovered criteria for the target image set. Finally, we deduplicate these criteria and add them to the criteria pool. Due to the context window limitations of LLMs, embedding all captions into a single prompt is infeasible. To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset. This strategy allows us to remain within the context length limits of modern LLMs (e.g.,, 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
Criteria Pool Refinement:
Table 11: Prompts for the LLM used in Proposed Criteria Refinement step We embed the exact initially discovered criteria by replacing the placeholders "{CRITERION}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation I am a machine learning researcher working with a set of images. I aim to cluster this set of images based on the various clustering criteria present within them. Below is a preliminary list of clustering criteria that I’ve discovered to group these images:
Criteria Embedding: * Criterion 1: "{CRITERION}" * Criterion 2: "{CRITERION}" * Criterion L: "{CRITERION}"
Goal Explanation My goal is to refine this list by merging similar criteria and rephrasing them using more precise and informative terms. This will help create a set of distinct, optimized clustering criteria.
Task Instruction Your task is to first review and understand the initial list of clustering criteria provided. Then, assist me in refining this list by: * Merging similar criteria. * Expressing each criterion more clearly and informatively.
Output Instruction Please respond with the cleaned and optimized list of clustering criteria, formatted as bullet points (using “*”). Your response:
In Tab. 11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024b). Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab. 11 and ask the LLM to merge similar criteria and rephrase their names to enhance clarity. This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
C.2 Prompts and Implementation Details of Semantic Grouper
Image-based Grouper:
Table 12: Prompts for the LLM used in the image-based grouper for automatic criterion-specific VQA question generation. We embed the exact discovered criterion by replacing the placeholder "{CRITERION}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Goal Explanation Hello! I am a machine learning researcher focusing on image categorization based on the aspect of "{CRITERION}" depicted in images.
Task Instruction Therefore, I need your assistance in designing a prompt for the Visual Question Answering (VQA) model to help it identify the "{CRITERION}" category in a given image at three different granularity. Please help me design and generate this prompt using the following template: "Question: [Generated VQA Prompt Question] Answer (reply with an abstract, a common, and a specific category name, respectively):". The generated prompt should be simple and straightforward.
Output Instruction Please respond with only the generated prompt using the following format “* Answer *”. Your response:
[…]
Rather than manually creating these questions, we embed the target criterion into the prompt shown in Tab. 12 and query the LLM to automatically generate high-quality, criterion-specific questions. These questions are then used to direct the VQA model, enabling it to accurately label each image according to the visual content relevant to the target criterion.
Tag-based Grouper:
Table 13: Prompts for the LLM used in the tag-based grouper for generating middle-grained criterion-specific tags. We embed the exact discovered criterion by replacing the placeholder "{CRITERION}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Goal Explanation Hello! I am a machine learning researcher focusing on image categorization of a certain aspect. I’m interested in generating a list of tags specifically for categorizing the types of "{CRITERION}" depicted in images.
Task Instruction Please provide a list of potential "{CRITERION}" category names. Please generate diverse category names. Do not include too general or specific category names such as “Sports”.
Output Instruction Please respond with the list of category names. Each category should be formatted as follows: “* Category Name”. Your response:
Table 14: Prompts for the LLM used in the tag-based grouper for generating coarse-grained criterion-specific tags. We embed the exact discovered criterion and middle-grained category by replacing the placeholder "{CRITERION}" and "{MIDDLE-GRAINED CATEGORY NAME}" in the prompt, respectively.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Task Instruction Generate a list of three more abstract or general "{CRITERION}" super-categories that the following "{CRITERION}" category belongs to and output the list separated by “&” (without numbers): "{MIDDLE-GRAINED CATEGORY NAME}"
Output Instruction Your response:
Table 15: Prompts for the LLM used in the tag-based grouper for generating fine-grained criterion-specific tags. We embed the exact discovered criterion and middle-grained category by replacing the placeholder "{CRITERION}" and "{MIDDLE-GRAINED CATEGORY NAME}" in the prompt, respectively.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Task Instruction Generate a list of ten more detailed or specific "{CRITERION}" sub-categories of the following "{CRITERION}" category and output the list separated by “&” (without numbers): "{MIDDLE-GRAINED CATEGORY NAME}"
Output Instruction Your response:
We present the prompts used in the tag-based grouper for querying the LLM Llama-3.1-8B. The prompt for generating criterion-specific tags is shown in Tab. 13, while the prompts for generating coarse-grained and fine-grained tags are shown in Tab. 14 and Tab. 15, respectively.
In the tag-based grouper, we begin by embedding the target criterion into the prompt from Tab. 13 to generate criterion-specific tags at a middle granularity. To enhance the diversity and coverage of the tags, we query the LLM 10 times and take the union of the generated tags after deduplication as candidates. Following the SHiNe framework (Liu et al., 2024d), for each middle-grained tag, we further embed it into the prompts from Tab. 14 and Tab. 15 to generate 3 super-categories (coarse-grained) and 10 sub-categories (fine-grained) for each tag.
After generating coarse and fine-grained categories for all middle-grained tags, we take the union of the super-categories as the coarse-grained tag candidates and the union of the sub-categories as the fine-grained tag candidates. Lastly, we use the open-vocabulary tagger CLIP ViT-L/14 to assign the most relevant tags to each image based on cosine similarity, using candidates from each granularity level. After tagging all the images, we group those sharing the same tag into clusters, yielding the clustering result. Note that we do not utilize lexical databases such as WordNet (Miller, 1995) or ConceptNet (Speer et al., 2017) for tag generation, as they do not support free-form input and may not capture certain discovered criteria.
Caption-based Grouper:
Table 16: Prompts for the MLLM used in the caption-based grouper for generating criterion-specific captions. We embed the exact discovered criterion by replacing the placeholder "{CRITERION}" in the prompt.
  Prompt purpose Prompt
  System Prompt You are a helpful AI assistant.
Task Instruction Analyze the image focusing specifically on the "{CRITERION}". Provide a detailed description of the "{CRITERION}" depicted in the image. Highlight key elements and interactions relevant to the "{CRITERION}" that enhance the understanding of the scene.
Output Instruction Your response:
Table 17: Prompts for the LLM used in the caption-based grouper at the Initial Naming step for initially assigning a criterion-based category name to the image based on its criterion-specific caption. We embed the exact discovered criterion and the corresponding criterion-specific caption by replacing the placeholder "{CRITERION}" and "{CRITERION-SPECIFIC CAPTION}" in the prompt, respectively.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation The following is the description about the "{CRITERION}" of an image:
Caption Embedding "{CRITERION-SPECIFIC CAPTION}"
Goal Explanation I am a machine learning researcher trying to assign a label to this image based on what is the "{CRITERION}" depicted in this image.
Task Instruction Understand the provided description carefully and assign a label to this image based on what is the "{CRITERION}" depicted in this image.
Output Instruction Please respond in the following format within five words: "*Answer*". Do not talk about the description and do not respond long sentences. The answer should be within five words.
Task Reinforcement Again, your job is to understand the description and assign a label to this image based on what is the "{CRITERION}" shown in this image. Your response:
Table 18: Prompts for the LLM used in the caption-based grouper at the Multi-granularity Cluster Generation step for refining the initially assigned names to a structured three granularity levels. We embed the exact discovered criterion and the initially assigned name categories by replacing the placeholder "{CRITERION}" and "{MIDDLE-GRAINED CATEGORY NAME}" in the prompt, respectively.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation The following is an initial list of "{CRITERION}" categories. These categories might not be at the same semantic granularity level. For example, category 1 could be “cutting vegetables”, while category 2 is simply “cutting”. In this case, category 1 is more specific than category 2.
Category Embedding * "{MIDDLE-GRAINED CATEGORY NAME}" * "{MIDDLE-GRAINED CATEGORY NAME}" * "{MIDDLE-GRAINED CATEGORY NAME}"
Task Instruction These categories might not be at the same semantic granularity level. For example, category 1 could be “cutting vegetables”, while category 2 is simply “cutting”. In this case, category 1 is more specific than category 2. Your job is to generate a three-level class hierarchy (class taxonomy, where the first level contains more abstract or general coarse-grained classes, the third level contains more specific fine-grained classes, and the second level contains intermediate mid-grained classes) of "{CRITERION}" based on the provided list of "{CRITERION}" categories. Follow these steps to generate the hierarchy.
Sub-task Instruction Follow these steps to generate the hierarchy: Step 1 - Understand the provided initial list of "{CRITERION}" categories. The following three-level class hierarchy generation steps are all based on the provided initial list. Step 2 - Generate a list of abstract or general "{CRITERION}" categories as the first level of the class hierarchy, covering all the concepts present in the initial list. Step 3 - Generate a list of middle-grained "{CRITERION}" categories as the second level of the class hierarchy, in which the middle-grained categories are the subcategories of the categories in the first level. The categories in the second-level are more specific than the first level but should still cover and reflect all the concepts present in the initial list. Step 4 - Generate a list of more specific fine-grained "{CRITERION}" categories as the third level of the class hierarchy, in which the categories should reflect more specific "{CRITERION}" concepts that you can infer from the initial list. The categories in the third-level are subcategories of the second-level. Step 5 - Output the generated three-level class hierarchy as a JSON object where the keys are the level numbers and the values are a flat list of generated categories at each level, structured like: { “level 1": [“categories”], “level 2": [“categories”], “level 3": [“categories”] }
Output Instruction Please only output the JSON object in your response and simply use a flat list to store the generated categories at each level. Your response:
Table 19: Prompts for the LLM used in the caption-based grouper at the Final Assignment step. We embed the exact discovered criterion and the refined category names from each granularity level, by replacing the placeholder "{CRITERION}" and "{CANDIDATE CATEGORY NAME}" in the prompt, respectively.
  Prompt purpose Prompt
  System Prompt You are a helpful assistant.
Input Explanation The following is a detailed description about the "{CRITERION}" of an image.
Caption Embedding "{CRITERION-SPECIFIC CAPTION}"
Task Instruction Based on the content and details provided in the description, classify the image into one of the specified "{CRITERION}" categories listed below:
Candidate Category Embedding "{CRITERION}" categories: * "{CANDIDATE CATEGORY NAME}" * "{CANDIDATE CATEGORY NAME}" * "{CANDIDATE CATEGORY NAME}"
Output Instruction Ensure that your classification adheres to the details mentioned in the image description. Respond with the classification result in the following format: “*category name*”. Your response:
We first present the MLLM prompt used for LLaVA-NeXT-7B (Liu et al., 2024b) to generate criterion-specific captions in Tab. 16. Following this, we present the LLM Llama-3.1-8B prompts used in the caption-based grouper for the Initial Naming step in Tab. 17, the Multi-granularity Cluster Refinement step in Tab. 18, and the Final Assignment step in Tab. 19.
Specifically, we begin by generating criterion-specific captions for each image using LLaVA-NeXT-7B with the prompt shown in Tab. 16. For each image, we then embed its criterion-specific caption and the relevant criterion into the LLM prompt shown in Tab. 17, querying the LLM to assign an initial name based on the target criterion. Once the initial names for all images in the dataset are obtained, we embed these names along with the target criterion into the prompt in Tab. 18 to query the LLM for cluster name refinement across three semantic granularity levels: coarse, middle, and fine.
Finally, for each image, we embed the target criterion, its criterion-specific caption, and cluster candidates from each granularity level into the prompt shown in Tab. 19, and use this to query the LLM for final cluster assignment at each granularity level.
Appendix D Further Implementation Details of the Compared Methods
[…]
Implementation details of MMaP (Yao et al., 2024): We closely followed the training configuration outlined in the original MMaP paper. Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset. We then prompt-tuned CLIP-ViT/B32 using Adam with a momentum of 0.9, training the model for 1,000 epochs for each criterion across all datasets. Hyperparameters were optimized according to the loss score of MMaP, with the learning rate searched in {0.1, 0.05, 0.01, 0.005, 0.001, 0.0005}, weight decay in {0.0005, 0.0001, 0.00005, 0.00001, 0}, $\alpha$ and $\beta$ in {0.0, 0.1, 0.2, …, 1.0}, and $\lambda$ fixed at 1 for all experiments. After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
Appendix E Supplementary Results of the Quantitative Experiments
In this section, we present additional numerical experiment results to supplement the figures in the main paper. In Sec. E.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework. In Sec. E.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks. Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods. Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in Sec. E.4.
E.1 Supplementary Results for Criteria Proposer Evaluation
We provide detailed numerical results corresponding to Fig. 4(a) in Tab. 20 and Fig. 4(c) in Tab. 21 for the six tested benchmarks.
Although captions generated by the MLLM may exhibit some information loss (e.g.,, ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g.,, introducing objects not present in the images) Liu et al. (2024a), these issues generally occur at the object or fine-grained attribute level. However, when reasoning about grouping criteria for SMC task, the focus is on identifying general thematic elements shared across the image set. As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
Table 20: Comparison of True Positive Rate (TPR) (%) for criteria proposers across the six SMC benchmarks. TPR performance is reported for both Basic and Hard ground-truth criteria. The best performance is highlighted in bold.
COCO-4c Food-4c Action-3c Clevr-4c Card-2c Fruit-2c Average
Basic Hard Basic Hard Basic Hard Basic Hard Basic Hard Basic Hard Basic Hard
Image-based 100.0 52.9 25.0 36.4 66.7 54.6 50.0 28.6 50.0 25.0 50.0 20.0 56.9 36.2
Tag-based 50.0 35.3 100.0 72.7 66.7 36.4 75.0 42.9 50.0 50.0 50.0 20.0 65.3 42.9
Caption-based 100.0 64.7 100.0 81.8 100.0 72.7 100.0 71.4 100.0 100.0 100.0 60.0 100.0 75.1
Table 21: Study of the impact of data scale on criteria discovery. The Caption-based Proposer is used for criteria discovery, and TPR performance (%) is reported on the Hard ground-truth criteria sets across the six SMC benchmarks for different data scales. The best performance is highlighted in bold.
Data scales COCO-4c Food-4c Action-3c Clevr-4c Card-2c Fruit-2c Average
100% 64.7 81.8 72.7 71.4 100.0 60.0 75.1
80% 47.1 72.7 54.6 71.4 75.0 30.0 58.5
60% 52.9 63.6 54.6 71.4 100.0 50.0 65.4
40% 41.2 45.5 45.5 85.7 100.0 40.0 59.6
20% 35.3 45.5 36.4 42.9 100.0 40.0 50.0
1 img 23.5 36.4 27.3 57.1 75.0 50.0 44.9
E.2 Supplementary Results for Semantic Grouper Evaluation
In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Fig. 5. Specifically, we provide detailed results for the evaluation of the six tested datasets as follows:
  • COCO-4c (Fig. 5(a)) in Tab. 22
  • Card-2c (Fig. 5(b)) in Tab. 23
  • Action-3c (Fig. 5(c)) in Tab. 24
  • Food-4c (Fig. 5(d)) in Tab. 25
  • Fruit-2c (Fig. 5(e)) in Tab. 26
  • Clevr-4c (Fig. 5(f)) in Tab. 27
Table 22: Comparison of Semantic Groupers on COCO-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(a).
Methods Activity Location Mood Time of Day
CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 62.6 73.5 67.6 34.3 51.5 41.1 22.4 43.3 29.5 40.6 74.1 52.4
KMeans CLIP 34.4 - - 32.7 - - 18.9 - - 38.6 - -
KMeans DINOv1 34.8 - - 37.5 - - 17.9 - - 36.5 - -
KMeans DINOv2 38.2 - - 37.9 - - 22.5 - - 43.8 - -
Img-based BLIP-2 48.7 64.1 55.3 39.6 48.0 43.4 30.2 37.5 33.4 40.7 60.3 48.6
Img-based LLaVA 46.5 61.8 53.1 34.0 46.3 39.2 28.0 24.7 26.3 39.4 51.7 44.7
Tag-based 43.2 51.5 47.0 28.6 46.6 35.5 13.0 25.6 17.2 19.3 48.8 27.7
Caption-based 44.1 48.9 46.4 55.2 55.6 55.4 38.1 32.6 35.2 67.6 56.7 61.7
Table 23: Comparison of Semantic Groupers on Card-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(b).
Methods Suit Rank
CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 47.9 69.5 56.7 35.0 64.2 45.3
KMeans CLIP 45.0 - - 28.6 - -
KMeans DINOv1 38.5 - - 20.7 - -
KMeans DINOv2 36.7 - - 22.3 - -
Img-based BLIP-2 66.7 77.7 71.8 47.5 54.4 50.7
Img-based LLaVA 36.8 65.8 47.2 24.6 49.8 32.9
Tag-based 39.2 32.9 35.8 22.3 39.1 28.4
Caption-based 54.5 73.6 62.6 92.1 95.1 93.6
Table 24: Comparison of Semantic Groupers on Action-3c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(c).
Methods Action Location Mood
CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 97.1 99.2 98.1 66.7 67.1 66.9 75.5 80.7 78.0
KMeans CLIP 62.3 - - 58.3 - - - -
KMeans DINOv1 49.3 - - 61.4 - - - -
KMeans DINOv2 75.7 - - 67.6 - - - -
Img-based BLIP-2 79.7 80.9 80.3 43.3 42.4 42.8 43.1 43.8 43.4
Img-based LLaVA 70.1 60.5 65.0 45.8 42.8 44.2 32.0 38.0 34.7
Tag-based 70.2 55.0 61.6 36.8 48.1 41.7 50.7 47.6 49.1
Caption-based 82.8 82.8 82.8 69.8 55.2 61.6 52.3 50.2 51.2
Table 25: Comparison of Semantic Groupers on Food-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(d).
Methods Food Type Cuisine Course Diet
CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 90.6 94.6 92.6 54.9 81.4 65.6 63.5 84.7 72.6 47.6 59.9 53.0
KMeans CLIP 66.1 - - 29.8 - - 49.5 - - 36.9 - -
KMeans DINOv1 33.6 - - 15.3 - - 38.1 - - 41.4 - -
KMeans DINOv2 72.7 - - 22.5 - - 47.6 - - 43.4 - -
Img-based BLIP-2 54.2 71.4 61.6 54.8 73.3 62.7 42.3 71.0 53.0 34.2 53.8 41.9
Img-based LLaVA 42.2 64.0 50.9 33.7 57.6 42.6 46.9 73.1 57.1 27.0 40.5 32.4
Tag-based 45.0 63.3 52.6 48.8 42.1 45.2 42.7 70.1 53.1 25.2 34.1 29.0
Caption-based 34.6 54.2 42.2 47.0 65.9 54.9 69.1 85.7 76.5 41.5 54.0 46.9
Table 26: Comparison of Semantic Groupers on Fruit-2c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(e).
Methods Species Color
CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 84.0 93.1 88.3 54.8 83.5 66.1
KMeans CLIP 67.1 - - 39.6 - -
KMeans DINOv1 53.8 - - 36.0 - -
KMeans DINOv2 71.2 - - 36.7 - -
Img-based BLIP-2 70.7 68.3 69.5 40.9 70.6 51.8
Img-based LLaVA 63.9 67.8 65.8 51.0 83.2 63.2
Tag-based 64.0 67.1 65.5 54.1 44.1 48.6
Caption-based 76.9 70.7 73.7 53.3 51.5 52.4
Table 27: Comparison of Semantic Groupers on Clevr-4c. We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%). These results are plotted in Fig. 5(f).
Methods Color Texture Count Shape
CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM CAcc SAcc HM
CLIP Zero-shot 77.7 94.0 85.1 34.1 41.9 37.6 43.7 81.5 56.9 71.1 72.7 71.9
KMeans CLIP 48.8 - - 61.4 - - 44.2 - - 56.1 - -
KMeans DINOv1 53.0 - - 58.4 - - 47.5 - - 67.0 - -
KMeans DINOv2 44.1 - - 46.9 - - 52.5 - - 87.0 - -
Img-based BLIP-2 69.3 76.5 72.7 57.8 34.4 43.1 25.7 55.9 35.2 69.1 62.6 65.7
Img-based LLaVA 56.5 63.5 59.8 51.9 26.9 35.4 53.7 39.4 45.4 64.3 71.3 67.6
Tag-based 66.6 55.3 60.4 57.2 40.2 47.3 47.4 8.3 14.1 62.7 36.5 46.2
Caption-based 70.3 63.4 66.7 65.3 42.1 51.2 65.7 73.3 69.3 58.4 38.5 46.4
E.3 Supplementary Results for Comparison Criteria-Conditioned Clustering Methods
We provide expanded results in Tab. 28 for each criterion and benchmark, detailing the comparison of criteria-conditioned clustering methods presented in Tab. 2 in the main paper.
[…]
CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
COCO-4c Activity 62.6 73.5 51.3 53.2 33.8 - 44.1 48.9
Location 34.3 51.5 58.5 54.0 35.3 - 55.2 55.6
Mood 22.4 43.3 23.2 40.4 20.9 - 38.1 32.6
Time of Day 40.6 74.1 62.8 65.2 45.7 - 67.6 56.7
$Avg.$ 40.1 60.6 48.9 53.2 33.9 - 51.2 48.4
Food-4c Food Type 90.6 94.6 36.0 52.6 48.9 - 34.6 54.2
Cuisine 54.9 81.4 46.8 42.4 31.7 - 47.0 65.9
Course 63.5 84.7 70.5 89.5 48.6 - 69.1 85.7
Diet 47.6 59.9 48.5 62.1 45.9 - 41.5 54.0
$Avg.$ 64.1 80.2 50.5 61.7 43.8 - 48.1 64.9
Clevr-4c Color 77.7 94.0 51.2 43.2 75.3 - 70.3 63.4
Texture 34.1 41.9 64.9 26.4 56.5 - 65.3 42.1
Count 43.7 81.5 46.9 39.0 53.9 - 65.7 73.3
Shape 71.1 72.7 70.0 38.7 65.5 - 58.4 38.5
$Avg.$ 56.7 72.5 58.3 36.8 62.8 - 64.9 54.3
Action-3c Action 97.1 99.2 86.4 58.7 51.3 - 82.8 76.3
Location 66.7 67.1 82.0 52.9 59.4 - 69.8 55.2
Mood 75.5 80.7 60.8 57.4 71.0 - 52.3 50.2
$Avg.$ 79.8 82.3 76.4 56.3 60.6 - 68.3 60.6
Card-2c Suit 47.9 69.5 54.9 65.6 41.3 - 54.5 73.6
Rank 35.0 64.2 94.6 96.8 32.6 - 92.1 95.1
$Avg.$ 41.4 66.9 74.8 81.2 36.9 - 73.3 84.3
Fruit-2c Species 84.0 93.1 69.3 66.9 58.8 - 76.9 70.7
Color 54.8 83.5 57.2 43.3 43.3 - 53.3 51.5
$Avg.$ 69.4 88.3 63.3 55.1 51.0 - 65.1 61.1
E.4 Supplementary Results for Studying the Necessity of Multi-granularity Cluster Generation.
We present expanded results in Tab. 29 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Fig. 6 in the main paper.
Table 29: Ablation study of multi-granularity refinement on the six SMC benchmarks. We compare three ways of constructing cluster names: Initial Names (IN), Flat Refinement (FR), Multi-granularity Refinement (MR). We report Clustering Accuracy (CAcc) and Semantic Accuracy (SAcc)as percentages (%). Average (Avg.) CAcc and SAcc across different criteria on each dataset is also provided. These expanded results correspond to the plotting shown in Fig. 6.
Benchmark Criterion IN FR MR
CAcc SAcc CAcc SAcc CAcc SAcc
COCO-4c Activity 14.1 48.5 34.5 40.5 44.1 48.9
Location 30.0 51.9 41.4 56.0 55.2 55.6
Mood 6.6 34.7 21.9 32.1 38.1 32.6
Time of Day 24.4 50.5 28.2 54.4 67.6 56.7
$Avg.$ 18.8 46.4 31.5 45.8 51.2 48.4
Food-4c Food Type 33.9 52.4 35.5 54.3 34.6 54.2
Cuisine 30.6 39.7 27.6 36.5 47.0 65.9
Course 52.9 81.1 62.8 83.0 69.1 85.7
Diet 14.0 46.6 36.8 58.2 41.5 54.0
$Avg.$ 32.9 55.0 40.7 58.0 48.1 64.9
Clevr-4c Color 56.5 49.7 60.9 53.0 70.3 63.4
Texture 56.5 26.0 60.9 33.0 65.3 42.1
Count 56.5 39.6 56.5 40.8 65.7 73.3
Shape 47.8 33.6 47.8 41.8 58.4 38.5
$Avg.$ 54.3 37.2 56.5 42.2 64.9 54.3
Action-3c Action 72.2 63.6 90.5 63.0 82.8 76.3
Location 46.0 50.4 65.9 59.3 69.8 55.2
Mood 20.6 41.9 46.0 51.0 52.3 50.2
$Avg.$ 46.3 52.0 67.5 57.8 68.3 60.6
Card-2c Suit 40.9 50.1 45.7 45.7 54.5 73.6
Rank 43.0 55.1 47.7 54.6 92.1 95.1
$Avg.$ 42.0 52.6 46.7 50.2 73.3 84.3
Fruit-2c Species 59.2 68.6 64.1 67.0 76.9 70.7
Color 41.8 56.7 44.7 42.3 53.3 51.5
$Avg.$ 50.5 62.7 54.4 54.7 65.1 61.1
Appendix F Further Studies of the Proposed Framework
In this section, we provide additional studies on our proposed framework, using the main configuration (Caption-based Proposer and Caption-based Grouper) for the analysis. In Sec. F.1, we conduct control experiments to examine the system sensitivity of our framework to different multimodal large language models (MLLMs) and large language models (LLMs). In Sec. F.2, we demonstrate how incorporating advanced prompting strategies can further enhance the framework’s performance on fine-grained criteria.
F.1 Further System Sensitivity Analysis of Various MLLMs and LLMs
Figure 10: Sensitivity analysis of different MLLMs and LLMs on the six SMC benchmarks. Top (a): We fix the LLM to Llama-3.1-8B and study the impact of different MLLMs. Bottom (b): We fix the MLLM to LLaVA-NeXT-7B and study the impact of different LLMs. The average clustering accuracy(%) across different criteria is reported on the left, while the average semantic accuracy(%) is reported on the right.
In Fig. 10, we perform a system-level sensitivity analysis using our default system configuration (caption-based proposer and caption-based grouper) to examine the impact of different MLLMs and LLMs on the system performance. Since all variants successfully propose the basic criteria in each benchmark, we report the average clustering accuracy (CAcc) and semantic accuracy (SAcc) across various criteria for comparative analysis.
Specifically, in Fig. 10(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024b) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b). Next, in Fig. 10(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024a), and Llama-3.1-8B.
Findings in Fig. 10(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language. On the other hand, this scalability demonstrates that our system can enhance performance with more robust MLLMs, thanks to its training-free design, which ensures compatibility with any MLLM. Despite this, we use LLaVA-NeXT-7B as our default MLLM due to its reproducibility, being open-source and unaffected by API changes, and its capacity for local deployment, which upholds privacy by not exposing sensitive image data to external entities.
As for the LLMs, as depicted in Fig. 10(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM. Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM. This consistency suggests that the reasoning task for SMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
F.2 Further Study on Fine-grained Image Collections
Image collections may include fine-grained grouping criteria, such as Bird species in bird photography. Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He & Peng, 2017). This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process. The straightforward captioning process in our current framework may not fully capture these subtle visual nuances. However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
We demonstrate this by enhancing our Caption-based Grouper with FineR (Liu et al., 2024e), a cross-modal CoT prompt method specifically designed for fine-grained visual recognition. When the proposer identifies fine-grained criteria, such as Bird species, the framework switches to a FineR-enhanced captioning strategy that provides more detailed attribute descriptions, such as “Wing color: Blue-grey,” to enrich the captions and capture per-attribute visual characteristics to better support the subsequent substructure uncovering process.
Table 30: Study of substructure discovery for fine-grained criteria. We report clustering accuracy (CAcc) and semantic accuracy (SAcc) as percentages (%). The pseudo upper-bound (UB) performance is obtained using CLIP (Radford et al., 2021) ViT-L/14 in a zero-shot transfer setting with the ground-truth class names. †: We compare with FineR (Liu et al., 2024e) without its post-class name refinement step to ensure a fair comparison.
CUB200 Car196
CAcc SAcc CAcc SAcc
UB 57.4 80.5 63.1 66.3
FineR† 44.8 64.5 33.8 52.9
Ours 30.1 56.7 21.3 35.9
Ours + FineR 45.1 68.9 31.1 47.3
We evaluate this on two image collections containing fine-grained criteria: CUB200 (Wah et al., 2011) and Stanford Cars196 (Khosla et al., 2011). Our framework successfully discovers the fine-grained criteria Bird species for CUB200 and Car model for Cars196. As shown in Tab. 30, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself. This demonstrates the flexibility of our system, allowing future adaptations to specific application needs, such as fine-grained image collections.
Appendix G Further Qualitative Results
In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). Specifically, we present example clustering results across different criteria for COCO-4c in Fig. 11, Food-4c in Fig. 12, Action-3c in Fig. 13, Clevr-4c in Fig. 14, and Card-2c in Fig. 15. Additionally, we showcase example clustering results at different predicted granularity levels for COCO-4c in Fig. 16.
Figure 11: Example predicted clusters of COCO-4c.
Figure 12: Example predicted clusters of Food-4c.
Figure 13: Example predicted clusters of Action-3c.
Figure 14: Example predicted clusters of Clevr-4c.
Figure 15: Example predicted clusters of Card-2c.
Figure 16: Example predicted clusters of COCO-4c at different granularities.
Appendix H Further Failure Case Analysis
Figure 17: Failure case analysis. We show wrongly predicted images with their ground-truth label for four clusters.
In Fig. 17, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). As observed, our method frequently mis-assigns “Surfing” to the “Kayaking” cluster under the Activity criterion. Upon examining the intermediate criterion captions generated by the MLLM, we found that this error is largely due to the MLLM incorrectly describing a “Surfboard” as a “Kayak”. This highlights the importance of the MLLM’s ability to accurately describe images, as it is critical for the performance of our system. Potential improvements could include majority voting or model ensembling using different MLLM models.
Another issue arises in crowded scenes. When multiple people are present in an image, the model consistently assigns the Mood label “Communal” to the images. We speculate that this occurs because, in the presence of multiple people, the model struggles to accurately determine the mood of one key individual.
Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels. For example, as shown in Fig. 17, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine. The “Edamame” shown in Fig. 17 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine. This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
Appendix I Further Details on Application Study
In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in Sec. 5 of the main paper. Specifically, Sec. I.1 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020). Sec. I.2 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria. Finally, Sec. I.3 provides additional insights from the analysis of social media image popularity.
I.1 Further Details on Discovering and Mitigating Dataset Bias
In this section, we provide additional evaluation results and implementation details for the application study presented in Sec. 5 of the main paper.
Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset (Liu et al., 2015) to assess the classification accuracy of them. Our method achieved an impressive classification accuracy of 99.1% for gender and 87.4% for hair color on the 162,770 training images, demonstrating its effectiveness for uncovering gender and hair color substructures within the training set.
In addition, we quantified the spurious correlation between hair color and gender using the metric proposed by Yang et al. (2023). Specifically, given the correlated gender attribute distribution $A$ and the target hair color distribution $Y$, we computed the normalized mutual information between $A$ and $Y$ to quantify the spurious correlation as:
$I(A;Y)=\frac{2I(A;Y)}{H(A)+H(Y)}$ (1)
where $H(A)$ and $H(Y)$ represent the normalized entropy of the gender and hair color distributions, respectively. A value of $H(A)$ or $H(Y)$ equal to 1 indicates a uniform distribution (i.e., no class imbalance). We then used the ground-truth distribution from the dataset’s labels and our predicted distribution to estimate the spurious correlation intensity using the score from Eq. 1. For gender and hair color, our method’s predictions yielded a score of $I_{Pred}=0.10$, which is nearly identical to the ground-truth score of $I_{GT}=0.11$. This demonstrates that our method effectively identifies and confirms the bias directly from the training set.
Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO (Sagawa et al., 2020), we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution. We closely followed the training protocol outlined in B2T (Kim et al., 2024) and GroupDRO (Sagawa et al., 2020). Specifically, we fine-tuned a ResNet-50 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), using the training split of the CelebA dataset (Liu et al., 2015). The training was performed using the SGD optimizer (Ruder, 2016) with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$. We applied a weight decay of 0.1 and set the group adjustment parameter to zero. The model was trained over 50 epochs. For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set. The final evaluation and comparison results are provided in Fig. 7.
I.2 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
Figure 18: Bias quantification results and human evaluation for each occupation and criterion across the two studied T2I models, DALL·E3 and SDXL. The bias intensity score is reported.
Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator). We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study. This resulted in a total of 1,800 images. For each occupation, we provide some examples of images generated by DALL·E3 in Fig. 20, while provide some examples of images generated by SDXL in Fig. 21. We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set. For this study, we utilized the mid-granularity output of our system. To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction. We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by D’Incà et al. (2024):
$\mathcal{H}_{bias}^{l}=1+\frac{\sum_{c^{l}\in\mathcal{C}^{l}}\log(p(c^{l}|\mathcal{C}^{l},\mathcal{D}_{\text{Occupation}}))}{\log(|\mathcal{C}^{l}|)}$ (2)
where $\mathcal{D}_{\text{Occupation}}$ represents the generated images for each occupation, $\mathcal{C}^{l}$ denotes the clusters discovered under the $l$-th criterion, and $p(c^{l}|\mathcal{C}^{l})$ is the probability of each cluster under the current distribution. The resulting score is bounded between $\mathcal{H}_{bias}^{l}\in[0,1]$, where 0 indicates no bias towards a specific cluster (concept) under the evaluated criterion, and 1 indicates that the images are completely biased towards a particular cluster (concept) (e.g.,, “Grey” hair color) within the current bias dimension (e.g., Hair color). We used the score defined in Eq. 2 to quantify the biases for each occupation across the 10 discovered grouping criteria. We report the bias intensity score for each occupation and each model across the 10 discovered grouping criteria in Fig. 18.
Human Evaluation Study Details: To assess the alignment between our method’s predictions and human judgments on bias detection, we conducted a user study to gather human evaluation results for the generated images. As shown in the questionnaire example in Fig. 22, participants were presented with images generated by DALL-E3 and SDXL for each occupation and were asked to identify the bias direction (dominant class) for each of the 10 discovered criteria and rate the bias intensity on a scale from 0 to 10. We collected responses from 54 anonymous participants, resulting in 6 human evaluations for each occupation and each criterion.
The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396. Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach. Detailed user study results are provided in Sec. I.2. We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers. In Fig. 18, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
Complete Results and Additional Findings: In Fig. 18, we present the detailed bias detection results for each model, occupation, and criterion, alongside human evaluation scores for reference. A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g.,, Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g.,, Hair color, Mood, Attire, and Accessories).
We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs. However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied. This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
I.3 Further Details on Analyzing Social Media Image Popularity
Figure 19: Complete analysis of social media photo popularity on the SPID dataset. We display the viral and major popular clusters, along with the popularity distribution of data points within these clusters across all ten discovered criteria (in Grey).
With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems. Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts. However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored. In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria). By analyzing these interpretable results alongside image popularity metrics (e.g.,, number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images. These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
To expand on the discussion in Sec. 5 of the main paper, we present the complete findings across all ten discovered criteria in Fig. 19. Notably, we consistently observed a sharp semantic contrast between the visual elements in viral images and those in the majority of popular images across all ten criteria. For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
This recurring observation reinforces the idea that viral content tends to capture more attention, likely because it features novel, surprising, or striking visual elements. Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer & Gore, 2014). On the other hand, widely popular yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality. We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
Figure 20: Samples of DALL·E3 generated images. For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to DALL·E3 to generate 100 images. We present a random sample of 30 generated images.
Figure 21: Samples of SDXL generated images. For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to SDXL to generate 100 images. We present a random sample of 30 generated images.
Figure 22: Example of the questionnaire for human evaluation study.
Appendix J In-depth Discussion of Related Settings
Distinction from Multiple Clustering: Multiple clustering involves finding diverse clusterings from the same dataset in either a semi-supervised (Bae & Bailey, 2006; Ren et al., 2022) or unsupervised (Caruana et al., 2006; Mautz et al., 2020; Yao et al., 2023) manner. However, multiple clustering methods require users to specify the number of clusterings and clusters (or rely on ground-truth annotations for evaluation) (Yu et al., 2024). In other words, $L$ (the number of clusterings) and $K_{l}$ (the number of clusters within each clustering) are assumed as prior knowledge in multiple clustering settings. This strong assumption creates a chicken-or-egg dilemma and is often impractical in real-world applications: Users want to employ multiple clustering methods to understand their data, but how can they predefine the number of valid clusterings and clusters without already having a deep understanding of the entire dataset? Although some strategies exist to determine the number of clusters, most of these methods (Monti et al., 2003; Zhang et al., 2017; Liu et al., 2024e) only work for single clustering. In contrast, SMC requires the model to automatically discover both the number of clusterings and the clusters within them.
Additionally, while multiple clustering methods can reveal diverse sample division patterns and underlying data structures, they do not provide interpretations of the results—specifically, what rules the output clustering follows and the semantic meaning of the clusters. As a result, users often need to manually investigate the clustering outcomes. In stark contrast, SMC methods provide interpretability by describing both the clustering rules and the semantic meanings of the clusters in natural language. This not only offers a more comprehensive understanding of the data but also allows users to combine different clusterings to gain deeper insights into the data distribution.
[…]
Distinction from Multi-label Zero-shot Classification: SMC differs from multi-label zero-shot classification in that the latter (Lee et al., 2018; Huang et al., 2020; Ali & Khan, 2023; Gupta et al., 2023) requires predefined sets of classes under different rules, with the goal being to assign each image to multiple classes from these sets. In contrast, SMC requires discovering the class names (or cluster semantics). In fact, multi-label zero-shot classification can be viewed as a specific instance of SMC when the user explicitly and precisely defines all the clusters and their semantic meanings under different clustering rules.

### v4

@@
 Abstract
-Organizing unstructured visual data into semantic clusters is a key challenge in computer vision.
-Traditional deep clustering (DC) approaches focus on a single partition of data, while multiple clustering (MC) methods address this limitation by uncovering distinct clustering solutions.
-The rise of large language models (LLMs) and multimodal LLMs (MLLMs) has enhanced MC by allowing users to define clustering criteria in natural language.
-However, manually specifying criteria for large datasets is impractical.
-In this work, we introduce the task Semantic Multiple Clustering (SMC) that aims to automatically discover clustering criteria from large image collections, uncovering interpretable substructures without requiring human input.
-Our framework, Text Driven Semantic Multiple Clustering (TeDeSC), uses text as a proxy to concurrently reason over large image collections, discover partitioning criteria, expressed in natural language, and reveal semantic substructures.
-To evaluate TeDeSC, we introduce the COCO-4c and Food-4c benchmarks, each containing four grouping criteria and ground-truth annotations.
-We apply TeDeSC to various applications, such as discovering biases and analyzing social media image popularity, demonstrating its utility as a tool for automatically organizing image collections and revealing novel insights.
-Disclaimer: Potentially sensitive content.
-Figure 1: Top: Semantic Multiple Clustering (SMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable groups (or semantic clusters), under multiple shared themes (or clustering criteria), without requiring any prior information.
-Bottom: Our proposed SMC system enables various applications like discovering novel biases in text-to-image generative (T2I) models.
+Organizing unstructured image collections into semantic clusters is a long-standing challenge.
+Traditional deep clustering techniques address this by producing a single data partition, whereas multiple clustering methods uncover diverse alternative partitions, but only when users predefine the clustering criteria.
+Yet expecting users to specify such criteria a priori for large, unfamiliar datasets is unrealistic.
+In this work, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), which aims to automatically discover clustering criteria from large, unstructured image collections, revealing interpretable substructures without human input.
+Our framework, $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, treats text as a reasoning proxy: it concurrently scans the entire image collection, proposes candidate criteria in natural language, and groups images into meaningful clusters per criterion.
+To evaluate progress, we release COCO-4c and Food-4c benchmarks, each annotated with four grouping criteria.
+Experiments show that $\mathcal{X}$-Cluster effectively reveals meaningful partitions and enables downstream applications such as bias discovery and social media image popularity analysis.
+We will open-source code and data to encourage reproducibility and further research.
+Organizing Unstructured Image Collections using Natural Language
+Mingxuan Liu1    Zhun Zhong4    Jun Li6    Gianni Franchi3    Subhankar Roy5    Elisa Ricci1,2 1 University of Trento   2 Fondazione Bruno Kessler  3 ENSTA Paris, Institut Polytechnique de Paris 4 Hefei University of Technology   5 University of Bergamo   6 Technical University of Munich
+Figure 1: Top: Open-ended Semantic Multiple Clustering (OpenSMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable clusters, under multiple shared themes or criteria, without requiring any prior knowledge.
+Bottom: Our system enables various applications like discovering novel biases in text-to-image (T2I) generative models.
+See our project webpage for code and data: here.
 1 Introduction
-Organizing large volumes of unstructured visual data into semantic clusters is a crucial problem in computer vision and has traditionally been addressed through the lens of unsupervised deep clustering (DC) (Caron et al., 2018; Van Gansbeke et al., 2020).
-Despite the advancements and improved performance, DC remains limited due to its inherently ill-posed nature (Estivill-Castro, 2002), as a dataset can be clustered in multiple ways (e.g., a dataset of a standard deck of cards can be grouped by suit or rank).
-Furthermore, the clusters discovered by DC rely on factors such as the inductive biases of the network, augmentations, and pretext tasks; and do not necessarily adhere to the a particular partitioning criteria that a user may have in mind (e.g., suit).
-The limitation of DC in exploring visual data only along a single partition is addressed by multiple clustering (MC) methods (Yu et al., 2024), which uncover distinct clustering solutions.
-Multiple partitions reveal hidden patterns in the data and allow for analysis from different perspectives.
-The recent advent of large language models (LLMs) and multimodal LLMs (MLLMs) has further enhanced MC by enabling users to specify the clustering criterion (e.g., group by suit) in natural language (Kwon et al., 2024; Yao et al., 2024).
-While incorporating a human in the loop is appealing, as it grants the user control over the clustering results, specifying a meaningful text criterion still requires prior understanding of the image collections.
-This task becomes cumbersome, especially due to the proliferation and high scene complexity inherent in large volumes of visual data.
+When grouping a large, unstructured collection of unlabeled images, several questions arise: i) Which criterion should guide the grouping?
+ii) Do multiple, equally valid criteria exist?
+iii) Given a criterion, how can we uncover the corresponding semantic clusters?
+iv) What should the clusters’ granularity be?
+Because the answers depend on both the collection’s visual content and the objectives of a specific application, each question is inherently open-ended.
+Having the ability to automatically resolve them would unlock hidden insights in the data and enable valuable applications, such as social media recommendation systems (Cheng et al., 2024) and dataset auditing (Bianchi et al., 2023).
+footnotetext: † Corresponding author.
+Previous studies have attempted to answer these questions.
+Deep Clustering (DC) approaches (Caron et al., 2018; Van Gansbeke et al., 2020) iteratively partition an unlabeled dataset but produce only a single partition.
+Multiple Clustering (MC) methods (Yu et al., 2024), instead, uncover several alternative clustering solutions, offering complementary perspectives.
+Yet both DC and MC ultimately reflect the inductive biases of the underlying networks, which can deviate from the partitions a user actually needs.
+Worse, the discovered criteria and clusters, denoted merely by index labels, lack human-interpretable semantics.
+To give users explicit control over clustering, Text Criterion conditioned Multiple Clustering (TCMC) (Yao et al., 2024, 2025) approaches let the users specify the desired criterion in natural language e.g., “group playing cards by Suit.” TCMC methods (Kwon et al., 2024; Luo et al., 2024b; Yao et al., 2024, 2025) first employ large language models (LLMs) to generate candidate cluster semantics conditioned on the user-specified criterion, then use vision–language models to assign images to semantically related clusters.
+Although TCMC methods yield user-guided and interpretable partitions, they presume the user already knows valid grouping criteria for the entire collection.
+As datasets grow and scenes become more complex (e.g., millions of social-media images (Hu et al., 2014)) knowing such prior knowledge becomes impractical.
+This leads to the question: Can clustering criteria and semantic substructures (groups) be discovered directly from visual data, without any human input?
+We formalize this challenge as the task of Open-ended Semantic Multiple Clustering (OpenSMC).
+As depicted in Figure 1(top), the objective in OpenSMC is to discover, in natural language, both the clustering criteria and their associated semantic clusters without relying on any human priors.
+The task is challenging because i) it requires concurrent reasoning over the entire image collection to surface valid criteria, posing a challenge to current vision systems, and ii) it is open-ended as neither user-defined knowledge about the clustering criteria nor clustering granularity is available.
+Tab.
+1 contrasts OpenSMC with existing clustering paradigms.
+In addition, due to the lack of practical benchmarks with diverse criteria for evaluating the OpenSMC, we introduce two realistic and larger benchmarks–COCO-4c and Food-4c–with annotations for up to four criteria.
+Table 1: Comparison of different clustering paradigms. Unlike DC, MC, and TCMC settings, the proposed OpenSMC task does not assume any prior knowledge and offers interpretable results.
+DC MC TCMC OpenSMC
+Prior Knowledge # Criteria
+Text Criteria
+Knowledge # Clusters
+Output Multiple Clustering
+Interpretable
+Open-ended
+To tackle OpenSMC, we leverage text as a compact proxy for reasoning because current vision models cannot yet encode and reliably reason over large image collections.
+Consequently, we propose $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a two-stage framework that combines multimodal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2025, 2023a) with text-only LLMs (Meta, 2024a).
+$\mathcal{X}$-Cluster comprises two modules: a Criteria Proposer, which jointly reasons over all images via their textual representation (embeddings, tags, or captions) to surface valid grouping criteria, and a Semantic Grouper, which uncovers semantic clusters and assigns images to them under each discovered criterion.
+As shown in Figure 1(top), $\mathcal{X}$-Cluster can help users explore unfamiliar image collections at scale, by automatically grouping the images into semantic clusters (e.g., “Surfing” and “Skateboarding”) under diverse, human-interpretable criteria (e.g., Activity).
+In this work, we investigate three designs for each module and demonstrate the flexibility and effectiveness of our framework.
+Finally, we showcase both the practicality and versatility of $\mathcal{X}$-Cluster across several applications, including uncovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images.
+For instance, as shown in Figure 1(bottom), the method uncovers well-known biases (e.g., Gender) alongside less explored ones (e.g., Hair color) in DALL·E3 -generated images (Betker et al., 2023).
+These results confirm that $\mathcal{X}$-Cluster is an automated, practical tool for extracting interpretable insights from large-scale, unstructured visual data, opening new avenues for research and industry.
+2 Related Work
+Image clustering.
+Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self-supervision (Caron et al., 2018, 2020; Ren et al., 2024).
+Multiple clustering extends this idea, seeking multiple non-redundant partitions with data augmentations, diversity losses, or subspace methods (Qi and Davidson, 2009; Hu and Pei, 2018; Yao et al., 2023b; Metaxas et al., 2023; Yu et al., 2024).
+Despite steady progress, both paradigms share key limitations: i) their results reflect inductive biases of the network and training algorithm, hampering generalization beyond object-centric data and often misaligning with user intent; and ii) clusters are returned as numeric indices rather than human-readable names.
+In contrast, $\mathcal{X}$-Cluster derives both meaningful criteria and cluster names directly from unlabeled data.
 […]
+While $\mathcal{X}$-Cluster also uses text as its reasoning medium, it departs from TCMC systems in two key aspects: i) it automatically discovers the grouping criteria rather than relying on a user-supplied one; ii) it infers both the number and the names of clusters, without any user input.
+Together, these features make $\mathcal{X}$-Cluster fully automatic and scalable, and well suited to uncover novel, often overlooked insights in large image collections.
+Topic Discovery.
+OpenSMC is also related to Topic Discovery (Blei et al., 2003) in NLP, which uncovers latent themes in large text corpora.
+Our work likewise seeks to reveal common thematic concepts, but from visual collections, and does so by using LLMs as the reasoning core.
+3 OpenSMC
+Task Definition: Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering (OpenSMC) is to build a system $\mathcal{H}$ that automatically i) discovers a set of $L$ grouping criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them.
+Formally, we define an OpenSMC system:
+$\mathcal{H}:\mathcal{D}\mapsto\left\{\mathcal{O}_{l}=\Bigl{\{}\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\Bigr{\}}_{k=1}^{K_{l}}\,\middle|\,R_{l}\right\}_{l=1}^{L},$
 […]
-Prior User-provided Criteria
-Knowledge # Clusters
-Out. Multiple Clustering
-Interpretability
-In this work, we propose to ease the burden of specifying text criteria off the user by automatically discovering the clustering criteria, and introduce the task of Semantic Multiple Clustering (SMC), as shown in Fig.
-1(top).
-To tackle SMC, we propose a general two-stage framework Text Driven Semantic Multiple Clustering (TeDeSC), powered by cutting-edge MLLMs and LLMs, that first discovers latent criteria (e.g., Activity and Location) from a given collection of images, and then groups the images into semantic clusters (e.g., “Surfing”,“Skateboarding” under Activity) for every discovered clustering criterion.
-To reason at a dataset level, TeDeSC first translates the visual data into textual data (e.g., captions), which in turn are then collectively used as a proxy for the LLM to discover the hidden patterns in the image collection.
-The uniqueness of TeDeSC lies in its: (i) comprehensiveness, as it can exhaustively discover multiple clustering criteria that may not be evident to a user; (ii) interpretability, as it outputs cluster names in natural language as opposed to traditional MC methods; (iii) flexibility, as it can discover clusters at multiple granularities (coarse to fine-grained), without needing to specify the number of clusters a priori; and (iv) generality, as, unlike the existing work (Yao et al., 2024), it is not limited to object-centric images but can handle complex scenes having fine-grained details.
-We summarize the key differences among various clustering solutions in Tab.
-1 and elaborate in App.
-A.
-Current benchmarks for evaluating multiple clustering either lack realism (Clevr-4c, Fruit-2c, Card-2c) or offer a limited number of criteria (Action-3c), as shown in Fig.
-2.
-To advance research in SMC, we introduce two challenging new benchmarks, COCO-4c and Food-4c, which depict images in daily life contexts and allow the dataset to be clustered as per four distinct criteria.
-Extensive experimental analyses on both existing and novel benchmarks demonstrate that the proposed TeDeSC can effectively discover meaningful clustering criteria and successfully group semantically similar images (some examples of clustering results by TeDeSC are shown in Fig.
-1 (top)).
-Lastly, we apply TeDeSC to a variety of applications, including discovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images.
-For example, as shown in Fig.
-1(bottom), TeDeSC can uncover less commonly studied biases (e.g., Hair color) in DALL·E3-generated (Betker et al., 2023) images, beyond the well-known biases (e.g., Gender) that may have already been corrected for.
-It achieves this by discovering various human-interpretable semantic clusters (e.g., “Dark Hair”, “Grey Hair”) across different discovered criteria and identifying overpopulated clusters.
-Similarly, it can identify the semantic visual elements that contribute to the popularity of social images, offering valuable insights to related practitioners.
-These results suggest that TeDeSC is an automatic, versatile, and highly practical tool, opening up numerous new application opportunities for future research and providing the potential to generate insights from unstructured visual data at scale across various domains.
-2 Semantic Multiple Clustering
-In this section, we first formally define the task of Semantic Multiple Clustering (SMC), and then introduce the existing and newly proposed benchmarks for SMC.
-Task Definition: Given a collection of $N$ unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of SMC is to categorize $\mathcal{D}$ into semantic clusters based on multiple criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$, where $L$ is the total number of criteria.
-In detail, a criterion $R_{l}$ refers to the grouping theme according to which a set of images can be organized (or partitioned), such as Activity, Location, Mood, or Time of Day.
-All semantic clusters $\mathcal{C}_{l}$ under a given criterion $R_{l}$ should align with the theme of that criterion.
-For instance, the semantic clusters under the criterion Activity should reflect activities depicted in the images (e.g., “Kayaking”, “Cooking”).
-The same dataset $\mathcal{D}$, when grouped by a different criterion such as Location, could be clustered into categories like “Outdoor”, “Indoor”, and so on.
-[…]
-Figure 2: Overview of the SMC benchmarks.
-We show all clustering criteria and the corresponding ground-truth labels for the example images.
-We introduce two new challenging benchmarks: COCO-4c and Food-4c.
-Benchmarks: Evaluating SMC methods requires benchmarks that can be partitioned under multiple valid criteria.
-Currently, only a few benchmarks (Yu et al., 2024) support the evaluation of SMC methods: Fruit-2c (Muresan & Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), and Clevr-4c (Vaze et al., 2024).
-As shown in Fig.
-2, these benchmarks are limited by their object-centric nature with simple backgrounds (e.g., Fruit-2c), an insufficient number of criteria (e.g., up to three in Action-3c), and a lack of photorealism due to synthetic generation (e.g., Clevr-4c).
-Given that the data encountered in real-world applications is more complex, we propose two new benchmarks for SMC: Food-4c and COCO-4c.
-Food-4c is sourced from Food-101 (Bossard et al., 2014), which includes 101 Food type (original annotations), along with new annotations for 15 Cuisine types, 5 Courses types, and 4 Diet preferences, totaling four clustering criteria.
-Additionally, we constructed COCO-4c using images from COCO-val (Lin et al., 2014), where we annotated four criteria with varying number of clusters: 64 Activity, 19 Location, 20 Mood, and 6 Time of day.
-Examples of these newly constructed benchmarks are shown in Fig.
-2.
-Further details, including the full list of cluster names and the annotation pipeline, are provided in App.
-B.
-3 Methodology: Text Driven Semantic Multiple Clustering
-Partitioning an unstructured image collection into semantic clusters based on different (unknown) criteria is challenging, as it requires reasoning over the visual contents of all the images concurrently.
-Specifically, the system needs to first find commonalities among the images for discovering the partitioning criterion (or theme) and then group the images into semantic clusters according to the discovered criterion.
-To tackle the challenging SMC task, we propose a two-stage framework, named Text Driven Semantic Multiple Clustering (TeDeSC), that significantly deviates from the commonly used technique of deep feature-based clustering, and instead uses text descriptions as a proxy to reason over the images and uncover hidden patterns in $\mathcal{D}$.
-We elaborate the key design differences between TeDeSC and related work in App.
-A.
-System overview: As illustrated in Fig.
-3, the proposed TeDeSC consists of two modules: Criteria Proposer and Semantic Grouper.
-The Criteria Proposer (or Proposer in short) processes the entire image set $\mathcal{D}$ to discover diverse common themes among the images and proposes grouping criteria $\mathcal{R}$ in natural language (e.g., Location).
-Once the criteria are proposed, the Semantic Grouper (or Grouper in short) discovers distinct semantic clusters that adhere to each criterion $R_{l}$ at varying levels of semantic granularity and assigns images to their respective clusters (e.g., “Climbing gym”).
-In this work, we explore various design choices for both the Proposer and Grouper, detailed in the following subsections.
-Full implementation details, including exact prompts, are provided in App.
+An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language.
+By contrast, in the TCMC setting humans predefine the criterion $\mathcal{R}$ and the desired cluster count $K$.
+Figure 2: OpenSMC benchmarks. We introduce two new challenging benchmarks: COCO-4c and Food-4c. We show all annotated criteria and the corresponding labels for the example images.
+Refer to caption(Prompts in App. E.1)(Prompts in App. E.2)
+Figure 3: $\bm{\mathcal{X}}$-Cluster overview.
+(Left) The Criteria Proposer reason over the entire image set via their textual proxies and deposits diverse, natural language grouping criteria in a Criteria Pool.
+(Middle) The Semantic Grouper draws each criterion from the pool, discovers semantic clusters at three levels of granularity, and assigns images to their proper clusters.
+(Right) Aggregating these assignments reveals the collection’s semantic substructure across multiple criteria and granularities.
+We evaluate three design variants for each module and adopt the best-performing configuration (marked Refer to caption) as the default $\mathcal{X}$-Cluster.
+See App.
+E for implementation details.
+Benchmark: Only a few benchmarks exist that are suitable for evaluating OpenSMC.
+As shown in Figure 2, current benchmarks: Fruit-2c (Muresan and Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), Clevr-4c (Vaze et al., 2024) are limited because they are object-centric (Fruit-2c), lacking photorealism (Clevr-4c), and too few criteria (e.g., at most three in Action-3c).
+To alleviate this issue, we introduce two new benchmarks: Food-4c and COCO-4c.
+Examples of images are presented in Figure 2.
+We extended Food-101 dataset (Bossard et al., 2014) to create Food-4c which in total has four grouping criteria.
+Similarly, we extend COCO-val (Lin et al., 2014) to create COCO-4c that has four grouping criteria.
+We believe that the newly introduced more realistic benchmarks will play a critical role in evaluating OpenSMC systems.
+Further benchmark statistics and annotation details are provided in App.
 C.
-Figure 3: TeDeSC is composed of a Criteria Proposer and a Semantic Grouper.
-(Left) The Proposer processes the entire image set to discover and propose diverse grouping criteria expressed in natural language, which are accumulated in the Criteria Pool.
-(Middle) The Grouper takes these proposed criteria from the pool, discovers semantic clusters linked to the criteria at three different levels of granularity, and assigns images to their respective clusters.
-(Right) Multiple substructures of the image set, indicated by the discovered criteria, are revealed across different granularities by combining the cluster assignment results.
-We explore various design choices, and set the best-performing methods from our experiments as the main configuration (marked with Refer to caption).
+4 Method
+To recap, the objective of an OpenSMC system is twofold: i) discover meaningful grouping criteria, or themes, in an unstructured image collection, and ii) group the images into semantic clusters for each discovered criterion.
+This task is particularly challenging because it requires simultaneous reasoning over a large number of images, which the current vision systems are not equipped to handle.
+System overview.
+We propose $\bm{\mathcal{X}}$-Cluster, a two-stage framework that is designed to perform joint reasoning over an entire image collection, as required by the OpenSMC task.
+At a high level, $\mathcal{X}$-Cluster addresses this task by first translating the visual content of the image collection into a lower-dimensional representation, i.e., natural language descriptions.
+We then exploit the large context window of the LLMs/MLLMs to discover the grouping criteria, followed by their semantic clusters, using the derived textual descriptions.
+Concretely, as shown in Figure 3, $\mathcal{X}$-Cluster comprise: (1) a Criteria Proposer that is in charge of discovering the grouping criteria (e.g., Activity, Mood); and (2) a Semantic Grouper that finds semantic clusters by conditioning the predictions on each discovered criterion (e.g., for a given image “Climbing gym” is predicted when conditioned on the criterion Activity, whereas “Adventurous” is predicted when conditioned on Mood).
+Next, we elaborate on each component of $\mathcal{X}$-Cluster.
 4.1 Criteria Proposer
-As shown in Fig.
-3(left), the Proposer takes the target image collection as input and outputs a set of discovered grouping criteria, expressed in natural language.
-The core of the Proposer design lies in its ability to concurrently analyze and reason over the visual content of a large set of images.
-With this in mind, we explore three systematic approaches, which we describe below.
-Image-based Proposer: We start with a baseline that leverages a state-of-the-art MLLM (Li et al., 2024) designed for multi-image reasoning out-of-the-box to directly infer the criteria given a set of images.
-To reason over a set of images, which is crucial for discovering the criterion, we stitch a batch of images into a $8\times 8$ image grid and input it to the MLLM as a single image.
-We then prompt the MLLM to propose grouping criteria for the images in the grid.
-All the resulting criteria from all such image grids are accumulated in a criteria pool, denoted as $\widetilde{\mathcal{R}}$.
-Tag-based Proposer: Next, we explore a tag-based approach that uses an open-vocabulary tagger (Radford et al., 2021) to assign 10 tags to each image in $\mathcal{D}$, using the vocabulary from WordNet (Miller, 1995).
-These tags serve as semantic descriptors, translating the visual content of each image into textual form.
-We then gather all the assigned tags from the images, input them into a LLM, and prompt it to discover grouping criteria based on these tags.
-Caption-based Proposer (Main): While image tags effectively capture certain visual semantics, we found that they predominantly reflect object-related content.
-To encompass a broader spectrum of visual information—such as environmental settings or interactions—we instead use a MLLM to generate captions for each image in $\mathcal{D}$.
-Descriptive captions provide a richer and more holistic semantic context.
-Staying within the 128k token limit of modern LLMs (Meta, 2024b; OpenAI, 2024), we feed a subset of captions into a LLM, which we prompt to elicit partitioning criteria.
-The criteria generated for each subset are accumulated in $\widetilde{\mathcal{R}}$.
-Experiments in Sec.
-4.1 show that the Caption-based Proposer is the most effective; so we use it as our main method and consider the others as baselines.
-Criteria refinement: Since the Proposers operate on subsets, the criteria accumulated in $\widetilde{\mathcal{R}}$ may include redundant or noisy entries.
-To address this, we input all initially proposed criteria from $\widetilde{\mathcal{R}}$ into a LLM (Meta, 2024b), prompting it to consolidate similar criteria and discard noisy ones.
-This step refines and updates the criteria pool into $\mathcal{R}$, which is ready to be used in the subsequent stage.
+We explore three Proposer designs that convert visual information into text, enabling the LLM to infer grouping criteria.
+Each design differs in how we convert the visual content into text and in how we use the text to derive the grouping criteria.
+Image-based Proposer: We prompt a MLLM (Li et al., 2025) to directly infer grouping criteria $\mathcal{R}$ from images, bypassing the translation of visual content into text.
+In detail, we present a mini-batch of images, tiled into a composite of $8{\times}8$ image grid, to the MLLM.
+This allows the MLLM to reason concurrently over a group of images.
+The image-based approach balances broader coverage of criteria with the preservation of fine-grained details, a trade-off introduced by the tiling operation.
+Tag-based Proposer.
+This variant adopts a tag-driven strategy.
+We use an open-vocabulary tagger (Radford et al., 2021) to assign up to ten WordNet (Miller, 1995) tags to each image, yielding concise textual descriptors.
+A LLM (Meta, 2024a) then summarizes all the aggregated tags in $\mathcal{D}$ to propose the grouping criteria $\mathcal{R}$.
+Caption-based Proposer (Main).
+While image tags capture certain visual semantics, they primarily focus on object-related content.
+To incorporate richer visual cues, such as scene context and interactions, we use a MLLM (Liu et al., 2024b) to caption every image in $\mathcal{D}$.
+We then split the captions into the largest chunks that fit the LLM’s context window (e.g.128k or 1M tokens) (Meta, 2024a; OpenAI, 2024) and prompt the LLM on each chunk to propose grouping criteria.
+The deduplicated criteria from all chunks form the pool $\mathcal{R}$.
 4.2 Semantic Grouper
-The automatically discovered criteria $\mathcal{R}$ serve as thematic indicators for revealing the semantic substructures (or clusters) within the image set $\mathcal{D}$.
-To uncover these substructures, as shown in Fig.
-3(right), the Grouper takes $\mathcal{D}$ and each criterion $R_{l}$ as inputs.
-It then discovers cluster names $\mathcal{C}_{l}=\{s_{k}^{l}\}_{k=1}^{K_{l}}$ conditioned on $R_{l}$ and assigns each image to its corresponding cluster.
-The core design of the Grouper focuses on aligning semantic substructure discovery with a specified partitioning criterion.
-Similar to the Proposer, we also explore three distinct approaches for the Grouper.
-Furthermore, as clusters under a given criterion can be formed at varying semantic granularities, depending on user preferences, we have designed our Grouper to cluster $\mathcal{D}$ at three levels of granularity: coarse, middle, and fine-grained.
-This design enables TeDeSC to provide new insights into the data at different granularities.
-For example, for the criterion food Cuisine, TeDeSC can organize images at a coarse-grained continental level (e.g., “European” or “Asian”), a middle-grained regional level (e.g., “Mediterranean” or “Southeast Asian”), or a fine-grained national level (e.g., “Italian” or “Thai”).
-Image-based Grouper: Given a target criterion $R_{l}\in\mathcal{R}$, we first prompt a LLM to generate a question $q_{l}$ specific to $R_{l}$-i.e., for criterion Mood the generated question is: “What mood is conveyed by this image?
-Answer with an abstract, common, and specific category name, respectively.” We then use $q_{l}$ to guide a visual question answering (VQA) model to infer semantic cluster names for each image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$.
-Consequently, by aggregating the cluster assignment results at each granularity level across $\mathcal{D}$, we can derive multi-granularity semantic substructures.
-Tag-based Grouper: Given a criterion, we prompt a LLM to generate a list of common, middle-grained tags specific to that criterion (e.g., “Recreational facility”).
+We explore three Grouper designs that partition the dataset $\mathcal{D}$ into semantic clusters by conditioning the partition on each criterion $R_{l}\in\mathcal{R}$.
+The end result is a partition (or substructure) $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\}_{k=1}^{K_{l}}$, where an image is not only assigned to a cluster, but also has an associated semantic name (e.g., “Climbing gym” instead of a cluster index), making it human interpretable.
+Each design differs in how we condition the predictor to incorporate the grouping criterion.
+Since data can be clustered at varying granularity, we design the Grouper to cluster images at three granular levels: coarse, middle, and fine.
+As an example, under the Cuisine criterion, the Grouper can cluster $\mathcal{D}$ at a coarse continental level (“European”, “Asian”), a middle regional level (“Mediterranean”, “Southeast Asian”), or a fine national level (“Italian”, “Thai”).
+Image-based Grouper.
+We task an LLM to generate a question $q_{l}$ conditioned on $R_{l}$.
+For e.g., for $R_{l}=\leavevmode\nobreak$Mood, the $q_{l}=$ “What mood does this image convey?
+Return an abstract, common, and specific category.” We then query a visual question answering (VQA) model (Li et al., 2023a) model to output a coarse-, mid-, and fine-grained label for every image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$.
+Subsequently, aggregating these labels over $\mathcal{D}$ at each level yields multi-granularity substructures.
+Tag-based Grouper: We first prompt a LLM to generate a list of middle-grained tags (e.g., “Recreational facility”) conditioned on a given criterion $R_{l}$.
 Following Liu et al.
-(2024d), we then obtain coarse- and fine-grained tags by querying the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag.
-Unlike lexical databases such as WordNet or ConceptNet (Speer et al., 2017), which do not support free-form input and may be limited in accommodating certain criteria, tag synthesis using a LLM provides a flexible and reliable alternative.
-Subsequently, we use an image tagger to assign the most relevant tag from the candidate tags at each granularity to each image, resulting in multi-granularity substructures after aggregation.
-Caption-based Grouper (Main): We prompt a MLLM to generate captions that specifically focus on the target criterion for each image, as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$.
-Next, we use a LLM in a three-step process to assign images to clusters at multiple granularity levels: i) Initial Naming: First, we prompt the LLM to assign a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, resulting in an initial set of names $\mathcal{S}_{\text{init}}^{l}$ of size $N$; ii) Multi-granularity Cluster Refinement: Using these initial names as basis, we prompt the LLM to refine them into three structured granularity levels: $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$.
-These structured names serve as candidates for cluster assignment; iii) Final Assignment: Each image is then assigned a class name from each granularity level based on its caption, as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$.
-Experiments in Sec.
-4.2 show that the Caption-based Grouper performs the best; thus we use it as our main method.
+(2024d), we then query the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag, obtaining coarse- and fine-grained tags.
+Given the vocabulary of tags, we use CLIP (Radford et al., 2021) to tag an image with the most likely label, one at each granular level.
+Caption-based Grouper (Main): We condition a MLLM with a given criterion $R_{l}$ to generate a set of captions for each image as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$.
+Next, we use a LLM in a three-step process to group images to clusters at multiple granularity levels: i) Initial Naming: The LLM assigns a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels as $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names each granularity level as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$.
 5 Experiments
 […]
-Evaluation metrics for Criteria Discovery: Given the discovered criteria set $\mathcal{R}$ of the proposer, we assess its quality from two dimensions: i) Comprehensiveness: We compute True Positive Rate (TPR) (Csurka et al., 2024) for the predicted criteria relative to the annotated ones as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, to assess to what extent the predicted set covers the ground-truth set $\mathcal{Y}$; ii) Diversity: We compute the pairwise semantic similarity between the predicted criteria $\mathcal{R}$, using Sentence-BERT (Reimers & Gurevych, 2019), and convert it into a diversity measure by subtracting the similarity from 1.
-The final score is the average of these pairwise values, reflecting how well the proposed criteria avoid redundancy and capture distinct aspects of the data—i.e., criteria such as Location and Place are nearly identical and provide little additional insight.
-A higher score indicates better diversity.
-It is important to note that the number of valid grouping criteria is subjective and potentially unlimited, making False Positives difficult to define.
-Thus, we use TPR as the primary evaluation metric.
-Evaluation metrics for substructure uncovering: Following Conti et al.
-(2023) and Liu et al.
-(2024e), for each criterion and its substructure uncovered by the grouper, we evaluate its alignment with the substructure defined by the ground-truth labels along two dimensions: i) Semantic Consistency: For each image, we measure the semantic similarity between its assigned cluster name and the ground-truth label under the current criterion using Sentence-BERT.
-The average similarity across the dataset is reported as Semantic Accuracy (SAcc), reflecting how well the predicted substructure semantically aligns with the ground-truth.
-ii) Structural Consistency: We compute the clustering accuracy (CAcc) between the predicted and ground-truth substructures (clusterings) to evaluate how accurately the predicted structure groups images.
-By combining CAcc and SAcc, we evaluate the overall alignment of the predicted substructure with the ground-truth across various criteria.
-Evaluation strategy: Since the granularity of the ground-truth annotations is unknown during evaluation, we choose the substructure from the three predicted granularities that achieves the highest CAcc relative to the ground-truth for assessing our TeDeSC.
-Unlike methods (Kwon et al., 2024; Yao et al., 2024), which rely on the ground-truth number of clusters to create flat clusterings that perfectly match the ground-truth granularity, we believe our approach offers a fair and practical evaluation strategy for systems capable of providing multi-granularity output.
-Figure 4: Comparison of Criteria Proposers.
-(a) Comprehensiveness: TPR performance of each proposer evaluated on Basic and Hard ground-truth criteria.
-(b) Diversity: Diversity scores of the proposed criteria for each proposer.
-(c) Influence of Image Quantity: TPR performance of the caption-based proposer on Hard ground-truth criteria across varying image scales used for criteria discovery.
+Evaluation metric for criteria discovery.
+We assess criterion discovery performance of the proposers using the True Positive Rate (TPR) (Csurka et al., 2024), defined as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, i.e., the fraction of ground-truth criteria $\mathcal{Y}$ recovered by the predicted set $\mathcal{R}$.
+Because the space of plausible criteria is inherently open-ended and subjective, False Positives are hard to define, so we use TPR as our primary metric: higher scores indicate more complete coverage of the reference annotated criteria.
+Evaluation metrics for substructure uncovering.
+We measure how well each criterion-specific substructure predicted by the Grouper matches the ground truth along two complementary axes: i) Semantic consistency: We quantify the semantic agreement between predicted cluster names and ground-truth labels by computing their cosine similarity with Sentence-BERT (Reimers and Gurevych, 2019) for each image; the dataset-level mean is reported as Semantic Accuracy (SAcc) (Conti et al., 2023; Liu et al., 2024e).
+ii) Structural consistency: We quantify the structural agreement between predicted and ground-truth clusters with clustering accuracy (CAcc) (Vaze et al., 2022), which computes Hungarian-matched accuracy between the two partitions (Kuhn, 1955).
+Because ground-truth granularity is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for evaluation.
+Unlike TCMC methods, which use the true cluster count during both training and evaluation, our protocol offers a fair and practical benchmark for OpenSMC.
+Figure 4: Comprehensiveness Comparison of Criteria Proposers: TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart.
+Each block represents one ground-truth criterion, with Colored blocks indicating successfully discovered criteria and Gray blocks representing undiscovered criteria.
+5.1 Study of the Criteria Proposer
+Expanding ground-truth criteria for rigorous evaluation.
+The annotated basic criteria may not capture all valid groupings in complex datasets such as COCO-4c.
+We engaged human annotators to expand additional criteria for the six benchmarks in § 3, obtaining {10, 4, 11, 7, 17, 11} criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}, respectively.
+We call the original per-image annotated criteria (Figure 2) the Basic ground truth and the enriched set the Hard ground truth.
+Full annotation details appear in App.
+C.2.
+Which Proposer performs best?
+Figure 4 compares the TPR performance for all Proposer variants.
+The Caption-based Proposer recovers the most comprehensive set of criteria, ranking closest to the human-annotated reference set on every benchmark and on both Basic and Hard splits; in the latter it outperforms the runner-up Tag-based Proposer by $+32.2\%$ TPR.
+Captions offer richer, nuanced cues that help the LLM discover grouping criteria more comprehensively, whereas tags provide little context and therefore struggle on complex datasets (e.g., COCO-4c).
+The Image-based Proposer lags behind, as reasoning over small image grids omits fine visual cues and obscures overarching themes.
+Impact of image quantity on criterion discovery.
+Figure 5 shows the TPR performance of the Caption-based Proposer as the number of input images for criteria discovery varies.
+Interestingly, for object-centric datasets (Card-2c, Clevr-4c), only a handful of samples – even a single image – recover most criteria, thanks to their uniform structure, i.e., seeing one playing card is enough to suggest criteria like Suit and Rank in Card-2c.
+In contrast, performance drops sharply on the more diverse and realistic datasets (COCO-4c, Food-4c, Action-3c) when fewer images are provided, because inferring their richer and intricate thematic criteria requires wider visual coverage.
+Since $\mathcal{X}$-Cluster assumes no dataset priors, we use all images by default.
+Figure 5: Impact of Image Quantity on Criteria Discovery. We evaluate the TPR performance of the Caption-based Proposer at different image scales against the Hard ground-truth criteria set.
+Figure 6: Comparison of Semantic Groupers.
+We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across six benchmarks.
+CLIP zero-shot classification (Refer to caption) serves as a pseudo upper bound, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline.
+Best performers are marked with Refer to caption.
+See App.
+G.2 for expanded numerical results and App.
+H for clustering visualizations.
+Table 2: Comparison with TCMC methods.
+For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria.
+We provide CLIP L/14 zero-shot performance as the pseudo upper-bound reference (UB).
+Note: -marked methods used the ground-truth criteria and the number of clusters ($K_{l}$) as prior input.
+MMaP and MSub do not build semantic clusters.
 See expanded results in App.
-E.1.
-4.1 Study of the Criteria Proposer
-Expanding ground-truth criteria for comprehensive evaluation: The newly introduced benchmarks provide up to four ground-truth criteria to evaluate the predicted criteria, but for image collections with high scene complexity (e.g.,, COCO-4c), these are still insufficient to cover all valid grouping options.
-To address this, we expanded the ground-truth criteria for each of the six benchmarks in Sec.
-2 by using human annotators to identify additional criteria.
-This resulted in expanded sets that have: {10, 4, 11, 7, 17, 11} distinct criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}.
-We refer to the per-image annotated criteria set (see Fig.
-2) as Basic ground truth and the expanded set as Hard.
-See App.
-B.2 for the exact criteria and annotation details.
-Which Criteria Proposer is best?
-Our comparative results, presented in Fig.
-4(a-b), show that the Caption-based Criteria Proposer consistently outperforms its counterparts by a significant margin across all six benchmarks.
-As shown in Fig.
-4(a), the caption-based proposer achieves an average of 100% TPR on the Basic criteria set and 75.1% TPR on the Hard set across the benchmarks.
-Its superior performance is particularly evident under the Hard criteria, where it surpasses the second-best tag-based proposer by +32.2% TPR.
-Meanwhile, the criteria predicted by the caption-based proposer also show greater diversity in Fig.
-4(b), coming closest to the diversity of human-annotated criteria.
-Although the tag-based proposer performs well on object-centric benchmarks like Clevr-4c or Food-4c, it significantly lags behind the caption-based proposer on more challenging benchmarks such as COCO-4c and Action-3c, which feature more complex, realistic images.
-This discrepancy arises because image tagging is a more object-centric way of representing image semantics, potentially missing the rich context depicted in images.
-Conversely, while the image-based proposer outperforms the tag-based proposer on COCO-4c and Action-3c, it is less effective overall because it is limited to reasoning over a small subset of images and loses visual details when combining images into a grid.
-Studying the influence of image quantity on Criteria Discovery: Fig.
-4(c) shows the TPR performance of the caption-based proposer across varying image scales (100% to only 1 image) used for criteria proposing, tested on the Hard criteria sets of six benchmarks.
-Interestingly, satisfactory performance is achieved with just a few images in object-centric benchmarks like Card-2c and Clevr-4c.
-In fact, even a single image often provides sufficient information for reasonable criteria discovery in these object-centric datasets, where a brief glimpse often represents the entirety adequately.
-For example, seeing one playing card allows the proposer to easily suggest criteria like “Rank” and “Suit” in the Card-2c dataset.
-However, this does not hold for more complex datasets like COCO-4c, Food-4c, and Action-3c, which contain diverse and realistic scenarios.
-Here, a reduction in image scale leads to a clear drop in TPR performance, as these datasets require a larger set of images to capture their intricate and varied thematic criteria.
-Since TeDeSC operates without prior knowledge of the dataset, we default to using the entire dataset to ensure comprehensive criteria discovery.
-Figure 5: Comparison of the Semantic Groupers.
-We show CAcc, SAcc, their Harmonic Mean (HM) evaluated on the basic criteria.
-The Caption-based Proposer is used by all groupers.
-See expanded results in App.
-E.2.
-4.2 Study of the Semantic Grouper
-Which Semantic Grouper is best?
-To determine which grouper best uncovers the substructures indicated by the discovered criteria and aligns most closely with the ground truth, we use the Harmonic Mean (HM) of the CAcc and SAcc for each criterion.
-This allows us to identify the best overall performer for each criterion.
-In the comparison, to further illustrate the relative performance position of our framework, we first establish a pseudo upper-bound reference by leveraging the zero-shot classification capabilities of CLIP ViT-L14 (Radford et al., 2021) with ground-truth labels for each criterion, where the grouping criteria, cluster names, and number of clusters are all known.
-Additionally, we apply KMeans (Ahmed et al., 2020) with the ground-truth number of clusters, using robust visual features from CLIP ViT-L14, DINOv1-B/16 (Li et al., 2022), and DINOv2-G/14 (Oquab et al., 2023) to establish baseline positions for clustering performance.
-Our comparative analysis, presented in Fig.
-5, shows that the Caption-based Semantic Grouper performs best, ranking first in 12 out of 19 tested criteria based on the HM, across six benchmarks.
-Specifically, in terms of CAcc, the caption-based grouper achieves an average of 59.9% across all tested criteria, closely aligning with the pseudo-upper-bound average of 58.1% and significantly outperforming feature representation-driven baselines, with the highest being DINOv2 at 40.6% CAcc.
-This demonstrates the effectiveness of our text-driven approach, as the cluster names inferred by the caption-based provide robust guidance for grouping.
-For SAcc, while the caption-based grouper falls short of the upper-bound reference average of 74.2%, where exact ground-truth class names are used, it still achieves a notable average of 60.5% across all tested criteria, surpassing its counterparts.
-To some extent, this gap in semantic consistency is expected due to the open nature of the semantic space—i.e., names like “Joyful”, “Happy”, and “Cheerful” are often used to describe the same Mood, yet they do not exhibit full semantic equivalence.
-The BLIP-2 image-based grouper also shows commendable performance, ranking first in 5 of the 19 tested criteria.
-This suggests that the criterion-customized prompts generated by LLM help BLIP-2 accurately label relevant visual content.
-However, its per-image labeling approach may lead to noisy clusters due to a lack of overall alignment between them.
-The tag-based grouper lags behind in all benchmarks likely due to a mismatch between the generated tags and the concepts in the dataset.
-Table 2: Comparison with criterion-conditioned clustering methods.
-For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria.
-We provide CLIP ViT-L/14 zero-shot performance as the pseudo upper-bound reference (UB).
-See expanded per-criterion results in App.
-E.3.
+G.3.
 COCO-4c Food-4c Clevr-4c Action-3c Card-2c Fruit-2c Avg
@@
 UB 40.1 60.6 64.1 80.2 56.7 72.5 79.8 82.3 41.4 66.9 69.4 88.3 50.2 64.4
+MMaP  (Yao et al., 2024) 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 48.2 -
+MSub  (Yao et al., 2025) 36.0 - 47.3 - 72.2 - 64.3 - 39.6 - 54.4 - 52.3 -
 […]
-MMaP 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 41.3 -
-TeDeSC (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 53.0 53.4
+SSD-LLM  (Luo et al., 2024b) 41.6 52.1 47.5 55.5 54.8 37.6 78.1 52.9 67.3 76.3 62.0 46.8 58.6 53.6
+$\mathcal{X}$-Cluster (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 61.8 62.3
+5.2 Study of the Semantic Grouper
+Which Semantic Grouper performs best?
+In Figure 6, we score each Grouper with CAcc (structural agreement) and SAcc (semantic agreement) evaluated on the ground-truth per-criterion annotation, ranking models by their harmonic mean Harmonic Mean (HM) of CAcc and SAcc.
+To gauge headroom we include a pseudo upper bound: CLIP ViT-L/14 in a zero-shot regime where the true criterion, cluster names and count are all supplied.
+We also report $k$-means baselines on features from CLIP-L/14, DINOv1-B/16 (Li et al., 2023b), and DINOv2-G/14 (Oquab et al., 2024), evaluated with CAcc and given the oracle $k$.
+Figure 6 shows the Caption-based Grouper performs best.
+It ranks first on 12/19 criteria and averages $59.9\%$ CAcc, virtually matching the upper bound’s $58.1\%$.
+Its mean SAcc is $60.5\%$, again best among learned methods but below the upper-bound $74.2\%$, which benefits from exact ground-truth names.
+This gap is expected given lexical variability, e.g., “Joyful,” “Happy,” and “Cheerful” all describe the Mood criterion but lack full semantic equivalence.
+The BLIP-2 Image-based Grouper ranks second.
+Its criterion-specific questions help, yet per-image answers add noise.
+The Tag-based Grouper trails across all benchmarks, reflecting the weak alignment between generated tags and dataset concepts.
+Effect of multi-granularity refinement.
+We ablate the Caption-based Grouper with three cluster naming schemes to study the effect of our multi-granularity design: i) Initial Names, where the initially LLM-assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and our iii) Multi-Granularity Refinement strategy.
+As Figure 7 indicates, either refinement method markedly boosts accuracy over raw initial names, confirming the need for consistent granularity.
+Furthermore, our multi-granularity strategy outperforms the flat one by offering several detail levels, providing greater flexibility in aligning with user-preferred grouping granularity.
+Figure 7: Study of multi-granularity refinement.
+5.3 Comparison with TCMC Methods
 […]
-Figure 6: Ablation study of multi-granularity refinement. For each benchmark, we show average CAcc and SAcc across all criteria. See expanded per-criterion numerical results in App. E.4.
-Necessity of multi-granularity cluster refinement: To validate the effectiveness of the multi-granularity cluster refinement design, we design control experiments with our Caption-based Grouper, using three different methods for constructing cluster names to organize images: i) Initial Names: using the initially assigned names as the final output, ii) Flat Refinement: prompting the LLM to refine the initial names into a flat list with a unified granularity, and our iii) Multi-granularity Refinement.
-We then compare the performance of these approaches in Fig.
-6.
-We observe that both refinement methods significantly improve clustering accuracy compared to using noisy initial names, highlighting the importance of having granularity-consistent cluster names for accurately revealing substructures.
-Additionally, our proposed multi-granularity method surpasses flat refinement, as it enables the grouper to generate clustering results at different granularities, providing greater flexibility in aligning with the label granularity-the user-preferred level of detail.
-Additional studies: In App.
-F, we provide further studies on: i) Fine-grained Scenario: We examine how to integrate our framework with more advanced cross-modal Chain-of-Thought prompting strategies to better handle fine-grained criteria.
-ii) Sensitivity Analysis: We analyze the system sensitivity to various MLLMs and LLMs.
-Additionally, Sec.
-G offers a Qualitative Analysis by visualizing clustering results for different criteria, and Sec.
-H includes a Failure Case Analysis.
+Additional ablations and studies of $\bm{\mathcal{X}}$-Cluster.
+In the technical appendix, we provide extensive ablations and follow-up experiments: i) Cluster visualizations and qualitative analysis (App.
+H); ii) Failure case analysis (App.
+I); iii) Handling of hallucinated criteria (App.
+K); iv) Mitigating model biases (App.
+L); v) Computational cost analysis (App.
+M); vi) Sensitivity analysis of different MLLMs and LLMs (App.
+N); vii) Further investigation on multi-granularity clustering (App.
+J); and (viii) Fine-grained criterion handling (App.
+O).
+Figure 8: Bias Discovery in T2I-Generated Images.
+Bias intensity, dominant clusters, and example images are shown for few occupations.
+Full results for all studied occupations are provided in App.
+P.1.
 6 Applications
-Discovering and mitigating dataset bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024).
-Figure 7: Results of dataset bias discovery and mitigation. Worst group and average accuracies(%) are reported.
-As a case study, we applied the proposed TeDeSC framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
-Additional details are provided in App.
-I.1.
-Findings: As expected, our method identified the grouping criteria Hair color and Gender.
-Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters.
-As shown in Fig.
-7(a), we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%).
-Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”.
-To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels.
-As shown in Fig.
-7(b), our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
-Figure 8: Bias discovery results in T2I models. Bias evaluation results are shown for the associated occupations.
-Discovering novel bias in Text-to-Image Diffusion Models: Stereotypical biases related to gender or race (Naik & Nushi, 2023) in images generated by widely accessible text-to-image (T2I) models like Stable Diffusion XL (SDXL) (Podell et al., 2024) and DALL·E3 (Betker et al., 2023) have sparked intense discussions (Nicoletti & Bass, 2023) and significant research (Bianchi et al., 2023).
-However, beyond these well-known biases extensively studied in social science, we wonder: Are there other biases present in T2I-generated images?
-To explore this question, we selected nine occupations (e.g.,, Nurse or CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) as subjects for our case study on bias.
-For each occupation, we generated 100 images using the prompt “A portrait photo of a $<$OCCUPATION$>$” with both DALL-E3 and SDXL, resulting in a total of 1.8k images (see Fig.
-20 and Fig.
-21 for examples).
-We applied TeDeSC to these images and automatically discovered 10 grouping criteria (possible bias dimensions) along with their predicted distributions for each occupation image set.
-To inspect the biases, we calculated the normalized entropy of each distribution to quantify the bias intensity, as proposed by D’Incà et al.
-(2024), and identified the dominant cluster—the one containing the most images—as the possible bias direction.
-Furthermore, we conducted a user study with 54 anonymous participants to assess how well our bias discovery results align with human judgment.
-We found that our system’s bias intensity scores closely matched human ratings—with an Absolute Mean Error (AME) of 0.1396 on a scale from 0 to 1—and that its predicted bias directions agreed with human evaluations 72.3% of the time.
-We highlight some key findings below and provide complete results and evaluations in App.
-I.2.
-Findings: In Fig.
-8, we present some of our notable findings.
-Without any predefined biases, our method successfully identifies well-known social biases associated with certain occupations.
-As shown in Fig.
-8(a–c), the SDXL-generated images exhibit near-total social bias for occupations like Nurse, Firefighter, and Basketball Player, displaying gender and racial imbalances that exceed official statistics such as those from U.S.
-Bureau of Labor Statistics (2021).
-In contrast, DALL·E3 shows markedly improved performance in mitigating social biases—likely due to its bias mitigation “guardrails” (OpenAI, 2022).
-More importantly, as shown in Fig.
-8(d–f), our method revealed previously unnoticed novel bias dimensions where both models exhibit moderate biases.
-For example, we found that SDXL strongly associates CEOs with “Grey” hair color, while DALL·E3 favors “Dark” hair.
-Interestingly, and contrary to trends observed in well-known social biases, DALL·E3 shows stronger biases than SDXL in Hair style and Grooming for Nurse and Teacher.
-For instance, 91% of the Teacher images generated by DALL·E3 depict individuals who are clean-shaven (Fig.
-8(f)).
-The gap in DALL·E3’s performance between addressing well-known and novel biases suggests that even industrial T2I models with guardrail systems may overlook some novel biases, underscoring the importance of thoroughly examining biases beyond the commonly recognized ones.
-Figure 9: Analysis of social media photo popularity on SPID dataset.
-We show the viral and major popular clusters along with the popularity distribution of data points within these clusters across five criteria (in Grey).
-Analyzing Social media image popularity: Given a set of social media photos with significantly varying popularity for each photo, it is natural to ask: What visual element makes a photo popular?
-As a case study, we applied TeDeSC to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), each accompanied by a popularity score based on the number of views.
-After our method grouped the photos into semantic clusters based on 10 discovered criteria, we calculated two scores for each cluster: i) $\bar{P}$: the average popularity of photos within the cluster, and ii) $\bar{P}_{w}$: the average popularity $\bar{P}$ weighted by the cluster’s proportion relative to the entire dataset.
-Here, we consider the cluster with the highest $\bar{P}$ as viral, meaning its photos are highly popular but may be few in number, and the cluster with the highest $\bar{P}_{w}$ as major popular, where the majority of popular photos are concentrated.
-We discuss the key findings below and provide complete results in App.
-I.3.
-Findings: As shown in our analysis in Fig.
-9, by combining the interpretable groupings from our method with popularity scores, the visual elements that drive virality (clusters with the highest $\bar{P}$) and the common traits shared among widely popular images (clusters with the highest $\bar{P}_{w}$) across different dimensions clearly emerged.
-Interestingly, we further observe that viral visual elements often exhibit sharp semantic contrast with those of the majority of popular images, such as Musical activities vs.
-Rest and relaxation, and High-intensity expressions vs.
-Neutral emotion.
-This suggests that viral content is more attention-grabbing, likely due to novel, unexpected, or intense visuals that stand out, especially in an era where the average attention span is less than 9 seconds (McSpadden, 2015; Farid, 2024).
-During our analysis, as shown in Fig.
-9, we unexpectedly found that some of the most popular images in certain clusters contained not safe for work (NSFW) content, which had not been previously identified in the SPID dataset.
-This suggests that some images gain popularity due to negative or provocative visuals.
-It also highlights the need for more thorough inspection of visual datasets to avoid harm, and our proposed framework can be a useful tool for this purpose.
+In this section, we apply the best caption-based $\mathcal{X}$-Cluster on three applications to showcase its ability to surface novel, human-interpretable insights.
+Discovering Novel Bias in Text-to-Image (T2I) Diffusion Models: Do T2I models encode biases beyond the well-studied gender and race stereotypes?
+(Naik and Nushi, 2023) To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) and generated 100 images per occupation via the prompt “A portrait photo of a <OCCUPATION>” using DALL-E3 (Betker et al., 2023) and SDXL (Podell et al., 2024), resulting in 1.8k images.
+Applying $\mathcal{X}$-Cluster, we automatically identified 10 grouping criteria (bias dimensions) and their distributions for each occupation.
+To quantify bias, we measured the normalized entropy of each dimension’s distribution (D’Incà et al., 2024) as bias intensity and marked the dominant (largest) cluster as the potential bias direction.
+We conducted a user study with 54 participants to validate our results: $\mathcal{X}$-Cluster’s predicted bias intensity closely matched ratings from the interviewees with an Absolute Mean Error of 0.1396 on a 0–1 scale, and bias directions agree 72.3% of the time with the interviewees.
+See App.
+P.1 for user study details.
+Findings: As shown in Figure 8, our method identifies both well-known and previously unnoticed novel biases without relying on predefined categories.
+For SDXL, roles such as Nurse, Firefighter and Basketball Player show pronounced gender and racial skews that exceed labor-force statistics (U.S.
+Bureau of Labor Statistics, 2021), whereas DALL-E3’s guardrails (OpenAI, 2022b) mitigate many of these effects.
+More notably, Figure 8(d–f) highlights previously unnoticed bias dimensions.
+For example, SDXL strongly associates CEOs with “Grey” hair, while DALL·E3 favors “Dark” hair.
+Additionally, DALL·E3 shows stronger biases in Hair style and Grooming for occupations like Nurse (Figure 8(e)) and Teacher (Figure 8(f)).
+These findings suggest that while industrial T2I models with guardrails may have mitigated well-known biases, they often overlook emerging or less-discussed ones – highlighting the need for broader audits, where $\mathcal{X}$-Cluster can serve as a valuable tool.
+See further findings in App.
+P.1.
+Analyzing Social Media Image Popularity: What visual traits drive popularity?
+To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), where popularity is measured by image view count.
+$\mathcal{X}$-Cluster discovered ten grouping criteria and organizes photos into semantic clusters under each.
+Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters for three criteria.
+Findings: As shown in Figure 9, combining $\mathcal{X}$-Cluster’s grouping with popularity scores provides a direct interpretation of the visual elements that drive trends versus those that define widely uploaded images.
+Interestingly, we find that trending elements often contrast with mainstream ones, such as “Musical activities” vs.
+“Rest and relaxation” or “High-intensity expressions” vs.
+“Neutral emotion”.
+These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era (McSpadden, 2015; Farid, 2024), underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior.
+For full findings and additional analysis, see App.
+P.2.
+Confirming and Mitigating Dataset Bias: Lastly, in App.
+P.3, we used $\mathcal{X}$-Cluster to confirm the gender bias in 162k CelebA (Liu et al., 2015) training images and successfully trained a debiased model using the predicted groupings with DRO (Sagawa et al., 2020).
+Tab.
+35 shows our debiased model achieves robustness comparable to B2T (Kim et al., 2024), which uses annotations for bias detection.
+Figure 9: Social media image popularity analysis.
+We show the popularity score distributions for Top Trending (have highest average popularity score) and Top Mainstream (contain most images) clusters, discovered by $\mathcal{X}$-Cluster across three criteria.
 7 Conclusion
-In this work, we introduce the task of semantic multiple clustering and propose TeDeSC, a system that automatically discovers grouping criteria in natural language from large image collections and uncovers interpretable data substructures based on these criteria.
-We rigorously evaluate various design choices of TeDeSC on four existing and two newly proposed benchmarks, and demonstrate its ability to reveal valuable insights that might otherwise go unnoticed in various real-world applications.
+We introduced the OpenSMC task and proposed $\mathcal{X}$-Cluster for discovering grouping criteria and semantic substructures in an unstructured image collections, expressed in natural language.
+We evaluated $\mathcal{X}$-Cluster on six benchmarks and three applications, showing its ability to automatically discover valuable insights stored in visual data.
 Acknowledgments
@@
 This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN.
+We thank OpenAI for supporting our experiments through the Researcher Access Program (0000004498).
 We extend our gratitude to Feng Xue for his valuable suggestions on plot creation.
@@
 would like to thank Margherita Potrich for her unwavering support.
+Limitations
+As discussed in § 5.1 and § 5.2, $\mathcal{X}$-Cluster still has significant room for improvement.
+It is thus not guaranteed to uncover all valid grouping criteria or perfectly align semantic substructures with user preferences.
+Moreover, although designed as an automatic, assistive tool for large-scale visual data analysis, $\mathcal{X}$-Cluster requires human oversight; users bear the ultimate responsibility for correctly interpreting the generated results.
+Since $\mathcal{X}$-Cluster relies on multimodal large language models (MLLMs) and large language models (LLMs), it inherits their limitations, including potential errors, biases, or hallucinations.
+Foundation models such as LLaVA-Next (Liu et al., 2024b), CLIP (Radford et al., 2021), and Llama series (Meta, 2024b, a) reflect biases present in their internet-scale training data (Goh et al., 2021; Schramowski et al., 2022; Kirk et al., 2021; Yamada et al., 2023; Wan et al., 2023).
+These biases can be challenging to identify or quantify due to the vast—and sometimes proprietary—nature of the training sets (Radford et al., 2021; OpenAI, 2022a, 2023).
+Furthermore, both MLLMs and LLMs may suffer from hallucinations.
+For LLMs (Ji et al., 2023; Yao et al., 2023a), this means generating information that appears accurate but is not.
+For MLLMs (or VQA models) (Ge et al., 2024), it involves creating image descriptions that seem reasonable but do not accurately reflect the visual content.
+To alleviate these limitations, we provide detailed analyses of failure cases (App.
+I) for our proposed $\mathcal{X}$-Cluster framework, strategies for handling hallucinated or invalid criteria (App.
+K), and methods to mitigate biases (App.
+L) in the technical appendix.
+Lastly, due to the zero-shot nature of $\mathcal{X}$-Cluster, its limitations may be mitigated by incorporating improved foundation models (OpenAI, 2025), employing bias alleviation techniques (Kong et al., 2024), or applying hallucination reduction strategies (Sahoo et al., 2024; Ge et al., 2024).
 Ethics Statement
 We do not anticipate any immediate negative societal impacts from our work.
-However, we encourage future researchers building on this work to remain vigilant, as we have, about the potential for TeDeSC, which integrates LLMs and MLLMs–particularly their human-like reasoning abilities– to be used both for good and for harm.
-The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs, as well as our exploration of social media image popularity, is to reveal and address these biases and the presence of sensitive or not-safe-for-work (NSFW) content that objectively exist in the datasets and models.
+However, we encourage future researchers building on this work to remain vigilant, as we have, about the potential for $\mathcal{X}$-Cluster, which integrates LLMs and MLLMs–particularly their human-like reasoning abilities– to be used both for good and for harm.
+The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs is to reveal and address these biases that objectively exist in the datasets and models.
 We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content.
-Specifically, in Sec.
-5, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016).
-Furthermore, our framework reveals previously undisclosed sensitive and sexual content in the SPID dataset (Ortis et al., 2019).
-We responsibly present these findings in Sec.
-5, applying significant blurring to disturbing content, with the intention of raising community awareness about the need to further scrutinize NSFW content in existing benchmarks.
+Specifically, in § 6, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016).
 However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups.
@@
 Our project, including data annotation and the user study involving human subjects, was approved by the Ethical Review Board of our university.
-Lastly, we emphasize that our proposed framework, TeDeSC, relies on open-source LLMs and MLLMs, allowing full deployment on local machines.
+Lastly, we emphasize that our proposed framework, $\mathcal{X}$-Cluster, relies on open-source LLMs and MLLMs, allowing full deployment on local machines.
 We refrain from using APIs from industrial LLMs or MLLMs, both to ensure reproducibility and to protect data privacy.
-Reproducibility Statement
-We will release all essential resources required to reproduce the experimental results presented in this project, including source code, exact prompts, benchmarks with their data splits, and generated images, upon publication.
-Our proposed framework, TeDeSC, is built on open-source, publicly accessible models to ensure reproducibility.
-In Sec.
-3, we provide a detailed description of how our framework is constructed.
-Additionally, App.
-C contains further implementation details, including exact prompts, to help practitioners easily reproduce our method.
-Details regarding the implementation of the compared methods are also provided in App.
-D.
-Moreover, we present extended numerical experimental results in App.
-E, alongside comprehensive findings for the application study in App.
-I.
-We believe that the thorough descriptions of our methodology, the extensive presentation of experimental results, and the open-source nature of our framework ensure that this work is highly reproducible, enabling future researchers and practitioners to readily apply our method to various domains.
 Appendix
@@
 […]
-This appendix provides detailed supplementary information supporting the implementations, experiments, and findings presented in the paper.
-First, in App.
-A, we offer an extensive discussion of Related Work, covering tasks and methods pertinent to our study.
-In App.
-B, we describe the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, along with the process for creating hard ground-truth criteria for evaluating different proposers.
+Overview
+This supplementary material provides additional details supporting the implementations, experiments, findings, and discussions in the main paper.
+In App. A, we first provide a reproducibility statement to ensure the transparency and replicability of our work.
 App.
-C outlines the prompts and implementation details for our proposed framework, including both Criteria Proposers and Semantic Groupers.
+B expands on Related Work, covering relevant tasks and methods.
 App.
-D provides a deep dive into the implementation specifics of the compared methods used in our experiments.
-In App.
-E, we present additional quantitative results that supplement the findings reported in the main paper, covering evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods.
+C describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation.
 App.
-F extends our analysis of the proposed framework, including system sensitivity and studies on fine-grained image collections.
+D details the evaluation metrics used in this work.
 App.
-G presents further qualitative results for the predicted clusters, while App.
-H offers a more detailed analysis of the failure cases encountered by our method.
+E presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers.
 App.
-I provides additional findings, along with implementation details and user study results for the three applications explored in this work.
+F provides implementation specifics of the compared methods.
+Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in App.
+G.
+For qualitative analysis, App.
+H presents further visualizations of predicted clusters, while App.
+I examines failure cases.
+App.
+J investigates the impact of multi-granularity clustering.
+App.
+K explores the effect of invalid (hallucinated) criteria on system performance, and App.
+L studies the influence of foundation model hallucinations and biases.
+App.
+M analyzes the computational cost and runtime of $\mathcal{X}$-Cluster.
+App.
+N and App.
+O extend our analysis with studies on system sensitivity and fine-grained image collections.
+App.
+P provides additional findings, implementation details, and user study results for the three explored applications.
+App.
+Q discusses how LLMs enhance image clustering.
 Finally, App.
-J presents a deeper discussion of related settings pertinent to our research.
-We will release all essential resources for reproducing this work, including code, prompts, benchmarks, and annotations, upon publication.
-Appendix A Related Work
-Image Clustering.
-Image clustering (Xu & Wunsch, 2005) discovers hidden grouping structures within large, unstructured, and unlabeled image collections, serving as a tool for various data-driven applications (Wazarkar & Keshavamurthy, 2018).
-To achieve this, deep clustering (DC) methods such as DEC (Xie et al., 2016) and SCAN (Van Gansbeke et al., 2020) focus on simultaneously learning feature representations and cluster assignments using deep neural networks via self-supervised techniques (Caron et al., 2020; Zhong et al., 2021; Ren et al., 2024).
-Furthermore, large-scale pre-trained feature representations like DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and CLIP (Radford et al., 2021) have also been shown to be effective at clustering image collections in a zero-shot fashion with the help of KMeans (Vaze et al., 2022; Liu et al., 2024f; Han et al., 2023).
-However, it is well-known that “clusters are in the eye of the beholder” (Estivill-Castro, 2002); there often exist multiple ways to partition the same image collection into clusters, and what constitutes a cluster depends on the user’s needs.
-This insight has led to the study of Multiple Clustering (MC) (Ren et al., 2022; Yao et al., 2023; Yu et al., 2024), which aims to simultaneously learn feature representations and cluster assignments from different perspectives to find various ways of grouping the same data, enabling alternative interpretations from different viewpoints.
-Although existing MC methods have achieved impressive results (Yao et al., 2023) on some benchmarks, they share similar limitations with deep clustering approaches (Xie et al., 2016; Van Gansbeke et al., 2020).
-Their results require intensive manual post-analysis, and they also hold strong assumptions: users must specify (i) the number of potential clusterings and (ii) the number of clusters within each clustering.
-However, when dealing with millions of unstructured images, it is infeasible for users—who are trying to understand the data—to know this information a priori.
-[…]
-In stark contrast to prior work, we introduce and study the task of semantic multiple clustering (SMC).
-Instead of requiring users to specify the grouping criteria, SMC seeks to actively and automatically discover criteria expressed in natural language from large visual data and uncover the corresponding semantic substructures, without access to any of the aforementioned human priors.
-As demonstrated in Sec.
-5, the flexibility provided by our proposed SMC framework adds significant value to various data-driven applications, unveiling novel insights about the data that might not have been noticed before.
+R outlines potential future research directions for our proposed OpenSMC task.
+The Table of Contents on the next page outlines the main topics in this supplementary material, with hyperlinks for direct navigation to each section.
+Appendix A Reproducibility Statement
+Upon publication, we will open-source all essential resources for reproducing this work.
+Specifically, we will provide the full code implementation of $\mathcal{X}$-Cluster, along with the exact prompts used in each module.
+Additionally, we will release our two newly proposed benchmarks, COCO-4c and Food-4c, including annotations for each grouping criterion.
+Lastly, we will provide the code for our evaluation protocol, experiments, and application studies.
+Appendix B Additional Related Work
 Topic Discovery.
-The setting of semantic multiple clustering (SMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund & Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g.,, documents).
+The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund and Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents).
 Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data.
 In contrast, our work focuses on discovering thematic criteria from large visual content.
-However, indeed, the core challenges of SMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data.
-Nevertheless, SMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them.
-Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024b).
+However, indeed, the core challenges of OpenSMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data.
+Nevertheless, OpenSMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them.
+Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024a).
 Multimodal Large Language Model.
 Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data.
-The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024a; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024b) alongside a pre-trained vision encoder (Radford et al., 2021).
+The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024b; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024a) alongside a pre-trained vision encoder (Radford et al., 2021).
 A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs.
-Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2023), and the LLaVA series (Liu et al., 2024c; b; Li et al., 2024).
-In our proposed TeDeSC framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria.
-Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2024) to establish a baseline image-based proposer for the SMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2023) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
+Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023a), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2024), and the LLaVA series (Liu et al., 2024c, b; Li et al., 2025).
+In our proposed $\mathcal{X}$-Cluster framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria.
+Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2025) to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2024) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
 […]
 Text-Driven Image Retrieval.
-Given a query text (e.g.,, “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query.
+Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query.
 In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques.
 In this context, the query can be considered as a sort of “cluster name”.
-However, this differs significantly from the proposed task of semantic multiple clustering (SMC), because SMC requires both discovering the textual criteria and the corresponding textual clusters.
-Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish SMC.
-Appendix B Benchmark Details
-B.1 Benchmark Construction of COCO-4c and Food-4c
+However, this differs significantly from the proposed task of open-ended semantic multiple clustering (OpenSMC), because OpenSMC requires both discovering the textual criteria and the corresponding textual clusters.
+Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish OpenSMC.
+Appendix C Benchmark Details
+C.1 Construction of COCO-4c and Food-4c
 Table 3: Summary of number of classes for the basic criteria annotation across the six benchmarks.
-Dataset Basic Criterion Number of Classes
-COCO-4c Activity 64
+Dataset Number of Images Basic Criterion Number of Classes
+COCO-4c 5,000 Activity 64
 Location 19
@@
 Time of Day 6
-Food-4c Food Type 101
+Food-4c 25,250 Food Type 101
 Cuisine 15
@@
 Diet 4
-Action-3c Action 40
+Action-3c 1,000 Action 40
 Location 10
 Mood 4
-Clevr-4c Color 10
+Clevr-4c 10,000 Color 10
 Texture 10
@@
 Count 10
-Card-2c Rank 14
+Card-2c 8,029 Rank 14
 Suit 5
-Fruit-2c Species 34
+Fruit-2c 103 Species 34
 Color 15
@@
 Diet “omnivore”, “vegan”, “vegetarian”, “gluten free”
-B.2 Details on Hard Grouping Criteria Annotation
+C.2 Details on Hard Grouping Criteria Annotation
 Table 6: Annotated criteria for the six benchmarks.
-The basic criteria are annotated on per-image level for each benchmark, while the hard criteria (those not in the basic criteria) are further exhaustively annotated by human annotators for further evaluating the performance of the rule proposer in SMC task.
+The basic criteria are annotated on per-image level for each benchmark, while the hard criteria (those not in the basic criteria) are further exhaustively annotated by human annotators for further evaluating the performance of the rule proposer in OpenSMC task.
 COCO-4c Food-4c Action-3c
@@
 This is particularly true for more complex and realistic datasets like COCO-4c, Food-4c, and Action-3c.
-As a result, the performance differences between different criteria proposers on these basic criteria, as shown in Fig.
-4, tend to be close to each other, limiting our understanding of each proposer’s ability to generate comprehensive grouping criteria.
+As a result, the performance differences between different criteria proposers on these basic criteria, as shown in Figure 4, tend to be close to each other, limiting our understanding of each proposer’s ability to generate comprehensive grouping criteria.
 To address this limitation, we employed human annotators to further identify and propose grouping criteria across the six benchmarks, resulting in a more extensive ground-truth set for each benchmark.
@@
 The proposed criteria from each annotator were then combined to form a comprehensive set of grouping criteria for the dataset.
-(2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g.,, Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
+(2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g., Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
 By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in Tab.
@@
 These expanded ground-truth sets enable us to more effectively evaluate the capabilities of various criteria discovery methods, providing a clearer understanding of different criteria proposers.
-Appendix C Further Implementation Details
+Appendix D Further Details of Evaluation Protocol
+Further Discussion on Clustering Accuracy (CAcc).
+Clustering Accuracy (CAcc) (Han et al., 2021) is evaluated by applying the Hungarian algorithm (Kuhn, 1955) to determine the optimal assignment between the predicted cluster indices and ground-truth labels.
+As extensively discussed in the GCD (Vaze et al., 2022) literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation.
+On the other hand, if the number of predicted clusters is lower than the number of ground-truth classes, the extra classes are assigned to a null set, and all instances with those ground-truth labels are similarly considered incorrect.
+Thus, CAcc is maximized only when the number of predicted clusters matches the number of ground-truth clusters.
+[…]
+Appendix E Further Implementation and Prompt Details
 In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App.
-C.1 and the Semantic Grouper in App.
-C.2.
-C.1 Prompts and Implementation Details of Criteria Proposer
+E.1 and the Semantic Grouper in App.
+E.2.
+E.1 Further Details of Criteria Proposer
 Image-based Proposer:
@@
 In Tab.
-7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2024).
+7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2025).
 Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images.
@@
 In Tab.
-8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024b).
+8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024a).
 For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image.
@@
 The criteria proposed by the LLM are then added to the criteria pool.
-Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g.,, 128k for Llama-3.1-8B) for the datasets used in our experiments.
+Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g., 128k for Llama-3.1-8B) for the datasets used in our experiments.
 However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
@@
 We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab.
-9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024b) in Tab.
+9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024a) in Tab.
 10.
@@
 To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset.
-This strategy allows us to remain within the context length limits of modern LLMs (e.g.,, 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
+This strategy allows us to remain within the context length limits of modern LLMs (e.g., 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
 Criteria Pool Refinement:
@@
 Output Instruction Please respond with the cleaned and optimized list of clustering criteria, formatted as bullet points (using “*”). Your response:
+Since the Proposers operate on subsets, the criteria accumulated in $\mathcal{R}$ may include redundant or noisy entries.
+Therefore, we further include a criteria poop refinement process.
+Specifically, we input all initially proposed criteria from $\mathcal{R}$ into a LLM, prompting it to consolidate similar criteria and discard noisy ones.
+This process refines and updates the criteria pool $\mathcal{R}$ for next stage.
 In Tab.
-11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024b).
+11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024a).
 Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab.
@@
 This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
-C.2 Prompts and Implementation Details of Semantic Grouper
+E.2 Further Details of Semantic Grouper
 Image-based Grouper:
@@
 19, and use this to query the LLM for final cluster assignment at each granularity level.
-Appendix D Further Implementation Details of the Compared Methods
+Appendix F Further Details of the Compared Methods
 […]
 […]
-Implementation details of MMaP (Yao et al., 2024): We closely followed the training configuration outlined in the original MMaP paper.
+[…]
+Implementation details of MMaP and MSub (Yao et al., 2024, 2025): We closely followed the training configuration outlined in the original MMaP and MSub paper.
 Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset.
@@
 After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
-Appendix E Supplementary Results of the Quantitative Experiments
+Appendix G Further Quantitative Experimental Results
 In this section, we present additional numerical experiment results to supplement the figures in the main paper.
-In Sec.
-E.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework.
-In Sec.
-E.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks.
+In § G.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework.
+In § G.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks.
 Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods.
-Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in Sec.
-E.4.
-E.1 Supplementary Results for Criteria Proposer Evaluation
-We provide detailed numerical results corresponding to Fig. 4(a) in Tab. 20 and Fig. 4(c) in Tab. 21 for the six tested benchmarks.
-Although captions generated by the MLLM may exhibit some information loss (e.g.,, ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g.,, introducing objects not present in the images) Liu et al.
+Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § G.4.
+G.1 Further Results for Criteria Proposer Study
+We provide detailed numerical results corresponding to Figure 4 in Tab. 20 and Figure 5 in Tab. 21 for the six tested benchmarks.
+Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g., introducing objects not present in the images) Liu et al.
 (2024a), these issues generally occur at the object or fine-grained attribute level.
-However, when reasoning about grouping criteria for SMC task, the focus is on identifying general thematic elements shared across the image set.
+However, when reasoning about grouping criteria for OpenSMC task, the focus is on identifying general thematic elements shared across the image set.
 As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
-Table 20: Comparison of True Positive Rate (TPR) (%) for criteria proposers across the six SMC benchmarks.
+Table 20: Comparison of True Positive Rate (TPR) (%) for criteria proposers across the six OpenSMC benchmarks.
 TPR performance is reported for both Basic and Hard ground-truth criteria.
@@
 Table 21: Study of the impact of data scale on criteria discovery.
-The Caption-based Proposer is used for criteria discovery, and TPR performance (%) is reported on the Hard ground-truth criteria sets across the six SMC benchmarks for different data scales.
+The Caption-based Proposer is used for criteria discovery, and TPR performance (%) is reported on the Hard ground-truth criteria sets across the six OpenSMC benchmarks for different data scales.
 The best performance is highlighted in bold.
@@
 1 img 23.5 36.4 27.3 57.1 75.0 50.0 44.9
-E.2 Supplementary Results for Semantic Grouper Evaluation
-In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Fig.
-5.
+G.2 Further Results for Semantic Grouper Study
+In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Figure 6.
 Specifically, we provide detailed results for the evaluation of the six tested datasets as follows:
-  • COCO-4c (Fig. 5(a)) in Tab. 22
-  • Card-2c (Fig. 5(b)) in Tab. 23
-  • Action-3c (Fig. 5(c)) in Tab. 24
-  • Food-4c (Fig. 5(d)) in Tab. 25
-  • Fruit-2c (Fig. 5(e)) in Tab. 26
-  • Clevr-4c (Fig. 5(f)) in Tab. 27
+In addition, we present the statistics of the predicted clusters at each granularity level in Tab. 28.
 Table 22: Comparison of Semantic Groupers on COCO-4c.
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(a).
+These results are plotted in Figure 6(a).
 Methods Activity Location Mood Time of Day
@@
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(b).
+These results are plotted in Figure 6(b).
 Methods Suit Rank
@@
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(c).
+These results are plotted in Figure 6(c).
 Methods Action Location Mood
@@
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(d).
+These results are plotted in Figure 6(d).
 Methods Food Type Cuisine Course Diet
@@
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(e).
+These results are plotted in Figure 6(e).
 Methods Species Color
@@
 We report Clustering Accuracy (CAcc), Semantic Accuracy (SAcc), and their Harmonic Mean (HM) in percentages (%).
-These results are plotted in Fig.
-5(f).
+These results are plotted in Figure 6(f).
 Methods Color Texture Count Shape
@@
 Caption-based 70.3 63.4 66.7 65.3 42.1 51.2 65.7 73.3 69.3 58.4 38.5 46.4
-E.3 Supplementary Results for Comparison Criteria-Conditioned Clustering Methods
+Table 28: Summary of cluster counts across six benchmarks for the comparison of semantic groupers.
+The results yield by the main Caption-based Grouper is reported.
+Specifically, we report: i) GT: the number of ground-truth clusters; ii) Pred-Init: predicted clusters from initial names; iii) Pred-Coarse: predicted coarse-grained clusters after multi-granularity refinement; iv) Pred-Middle: predicted middle-grained clusters after multi-granularity refinement; and v) Pred-Fine: predicted fine-grained clusters after multi-granularity refinement.
+Dataset Criteria GT Pred-Init Pred-Corase Pred-Middle Pred-Fine
+COCO-4c Activity 64 203 12 23 52
+Location 19 145 7 14 28
+Mood 20 122 15 25 30
+Time of Day 6 96 2 8 31
+Food-4c Food Type 101 301 7 37 127
+Cuisine 15 141 9 18 53
+Course 5 97 4 12 78
+Diet 4 139 5 8 64
+Action-3c Action 40 71 8 15 51
+Location 10 82 5 10 67
+Mood 4 95 6 18 55
+Clevr-4c Color 10 25 6 12 17
+Texture 10 23 2 5 12
+Shape 10 22 5 11 14
+Count 10 11 2 4 11
+Card-2c Rank 14 147 4 7 16
+Suit 5 56 4 7 30
+Fruit-2c Species 34 54 8 25 38
+Color 15 66 5 15 39
+G.3 Further Comparative Results with TCMC Methods
 We provide expanded results in Tab. 29 for each criterion and benchmark, detailing the comparison of criteria-conditioned clustering methods presented in Tab. 2 in the main paper.
@@
 […]
-CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
-COCO-4c Activity 62.6 73.5 51.3 53.2 33.8 - 44.1 48.9
-Location 34.3 51.5 58.5 54.0 35.3 - 55.2 55.6
-Mood 22.4 43.3 23.2 40.4 20.9 - 38.1 32.6
-Time of Day 40.6 74.1 62.8 65.2 45.7 - 67.6 56.7
-$Avg.$ 40.1 60.6 48.9 53.2 33.9 - 51.2 48.4
-Food-4c Food Type 90.6 94.6 36.0 52.6 48.9 - 34.6 54.2
-Cuisine 54.9 81.4 46.8 42.4 31.7 - 47.0 65.9
-Course 63.5 84.7 70.5 89.5 48.6 - 69.1 85.7
-Diet 47.6 59.9 48.5 62.1 45.9 - 41.5 54.0
-$Avg.$ 64.1 80.2 50.5 61.7 43.8 - 48.1 64.9
-Clevr-4c Color 77.7 94.0 51.2 43.2 75.3 - 70.3 63.4
-Texture 34.1 41.9 64.9 26.4 56.5 - 65.3 42.1
-Count 43.7 81.5 46.9 39.0 53.9 - 65.7 73.3
-Shape 71.1 72.7 70.0 38.7 65.5 - 58.4 38.5
-$Avg.$ 56.7 72.5 58.3 36.8 62.8 - 64.9 54.3
-Action-3c Action 97.1 99.2 86.4 58.7 51.3 - 82.8 76.3
-Location 66.7 67.1 82.0 52.9 59.4 - 69.8 55.2
-Mood 75.5 80.7 60.8 57.4 71.0 - 52.3 50.2
-$Avg.$ 79.8 82.3 76.4 56.3 60.6 - 68.3 60.6
-Card-2c Suit 47.9 69.5 54.9 65.6 41.3 - 54.5 73.6
-Rank 35.0 64.2 94.6 96.8 32.6 - 92.1 95.1
-$Avg.$ 41.4 66.9 74.8 81.2 36.9 - 73.3 84.3
-Fruit-2c Species 84.0 93.1 69.3 66.9 58.8 - 76.9 70.7
-Color 54.8 83.5 57.2 43.3 43.3 - 53.3 51.5
-$Avg.$ 69.4 88.3 63.3 55.1 51.0 - 65.1 61.1
-E.4 Supplementary Results for Studying the Necessity of Multi-granularity Cluster Generation.
-We present expanded results in Tab. 29 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Fig. 6 in the main paper.
-Table 29: Ablation study of multi-granularity refinement on the six SMC benchmarks.
+CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc CAcc SAcc
+COCO-4c Activity 62.6 73.5 51.3 53.2 44.0 52.1 33.8 - 35.9 - 44.1 48.9
+Location 34.3 51.5 58.5 54.0 51.2 52.9 35.3 - 37.4 - 55.2 55.6
+Mood 22.4 43.3 23.2 40.4 15.9 39.3 20.9 - 23.0 - 38.1 32.6
+Time of Day 40.6 74.1 62.8 65.2 55.5 64.1 45.7 - 47.8 - 67.6 56.7
+$Avg.$ 40.1 60.6 48.9 53.2 41.6 52.1 33.9 - 36.0 - 51.2 48.4
+Food-4c Food Type 90.6 94.6 36.0 52.6 33.1 46.5 48.9 - 52.4 - 34.6 54.2
+Cuisine 54.9 81.4 46.8 42.4 43.9 36.3 31.7 - 35.2 - 47.0 65.9
+Course 63.5 84.7 70.5 89.5 67.6 83.4 48.6 - 52.1 - 69.1 85.7
+Diet 47.6 59.9 48.5 62.1 45.6 56.0 45.9 - 49.4 - 41.5 54.0
+$Avg.$ 64.1 80.2 50.5 61.7 47.5 55.5 43.8 - 47.3 - 48.1 64.9
+Clevr-4c Color 77.7 94.0 51.2 43.2 47.8 44.0 75.3 - 84.7 - 70.3 63.4
+Texture 34.1 41.9 64.9 26.4 61.5 27.2 56.5 - 65.9 - 65.3 42.1
+Count 43.7 81.5 46.9 39.0 43.5 39.8 53.9 - 63.3 - 65.7 73.3
+Shape 71.1 72.7 70.0 38.7 66.6 39.5 65.5 - 74.9 - 58.4 38.5
+$Avg.$ 56.7 72.5 58.3 36.8 54.8 37.6 62.8 - 72.2 - 64.9 54.3
+Action-3c Action 97.1 99.2 86.4 58.7 88.1 55.3 51.3 - 55.0 - 82.8 76.3
+Location 66.7 67.1 82.0 52.9 83.7 49.5 59.4 - 63.1 - 69.8 55.2
+Mood 75.5 80.7 60.8 57.4 62.5 54.0 71.0 - 74.7 - 52.3 50.2
+$Avg.$ 79.8 82.3 76.4 56.3 78.1 52.9 60.6 - 64.3 - 68.3 60.6
+Card-2c Suit 47.9 69.5 54.9 65.6 47.5 60.7 41.3 - 44.0 - 54.5 73.6
+Rank 35.0 64.2 94.6 96.8 87.2 91.9 32.6 - 35.3 - 92.1 95.1
+$Avg.$ 41.4 66.9 74.8 81.2 67.3 76.3 36.9 - 39.6 - 73.3 84.3
+Fruit-2c Species 84.0 93.1 69.3 66.9 68.1 58.6 58.8 - 62.2 - 76.9 70.7
+Color 54.8 83.5 57.2 43.3 56.0 35.0 43.3 - 46.7 - 53.3 51.5
+$Avg.$ 69.4 88.3 63.3 55.1 62.0 46.8 51.0 - 54.4 - 65.1 61.1
+G.4 Further Results for Multi-granularity Clustering Study.
+We present expanded results in Tab. 30 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Figure 7 in the main paper.
+Table 30: Ablation study of multi-granularity refinement on the six OpenSMC benchmarks.
 We compare three ways of constructing cluster names: Initial Names (IN), Flat Refinement (FR), Multi-granularity Refinement (MR).
@@
 Average (Avg.) CAcc and SAcc across different criteria on each dataset is also provided.
-These expanded results correspond to the plotting shown in Fig.
-6.
+These expanded results correspond to the plotting shown in Figure 7.
 Benchmark Criterion IN FR MR
@@
 $Avg.$ 50.5 62.7 54.4 54.7 65.1 61.1
-Appendix F Further Studies of the Proposed Framework
-In this section, we provide additional studies on our proposed framework, using the main configuration (Caption-based Proposer and Caption-based Grouper) for the analysis.
-In Sec.
-F.1, we conduct control experiments to examine the system sensitivity of our framework to different multimodal large language models (MLLMs) and large language models (LLMs).
-In Sec.
-F.2, we demonstrate how incorporating advanced prompting strategies can further enhance the framework’s performance on fine-grained criteria.
-F.1 Further System Sensitivity Analysis of Various MLLMs and LLMs
-Figure 10: Sensitivity analysis of different MLLMs and LLMs on the six SMC benchmarks.
+Appendix H Qualitative Analysis
+In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
+Specifically, we present example clustering results across different criteria for COCO-4c in Figure 10, Food-4c in Figure 11, Action-3c in Figure 12, Clevr-4c in Figure 13, and Card-2c in Figure 14.
+Additionally, we showcase example clustering results at different predicted granularity levels for COCO-4c in Figure 15.
+Figure 10: Example predicted clusters of COCO-4c.
+Figure 11: Example predicted clusters of Food-4c.
+Figure 12: Example predicted clusters of Action-3c.
+Figure 13: Example predicted clusters of Clevr-4c.
+Figure 14: Example predicted clusters of Card-2c.
+Figure 15: Example predicted clusters of COCO-4c at different granularities.
+Appendix I Failure Case Analysis
+In Figure 16, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
+As observed, our method frequently mis-assigns “Surfing” to the “Kayaking” cluster under the Activity criterion.
+Upon examining the intermediate criterion captions generated by the MLLM, we found that this error is largely due to the MLLM incorrectly describing a “Surfboard” as a “Kayak”.
+This highlights the importance of the MLLM’s ability to accurately describe images, as it is critical for the performance of our system.
+Potential improvements could include majority voting or model ensembling using different MLLM models.
+Another issue arises in crowded scenes.
+When multiple people are present in an image, the model consistently assigns the Mood label “Communal” to the images.
+We speculate that this occurs because, in the presence of multiple people, the model struggles to accurately determine the mood of one key individual.
+Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels.
+For example, as shown in Figure 16, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine.
+The “Edamame” shown in Figure 16 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine.
+This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
+Figure 16: Failure case analysis. We show wrongly predicted images with their ground-truth label for four clusters.
+Appendix J Further Study on Multi-granularity Clustering
+In this section, we provide a detailed study on how different levels of multi-granularity output from our $\mathcal{X}$-Cluster framework impact grouping results.
+Specifically, for the Action-3c dataset, we employed human annotators to label two additional granularity levels for the criteria Action and Location.
+For the Action criterion, we consider the original annotation as fine-grained (L3) and tasked annotators to name the action in the image using more abstract and general coarse-grained (L1) and middle-grained (L2) labels.
+For the Location criterion, we consider the original annotation as middle-grained (L2) and tasked annotators to provide both more abstract coarse-grained (L1) labels and more specific fine-grained (L3) labels.
+This process resulted in expanded ground-truth annotations at three distinct semantic granularity levels for both the Action and Location criteria of the Action-3c dataset.
+Figure 17: Further study on the influence of multi-granularity clustering output.
+We evaluate the CAcc and SAcc of the multi-granularity grouping results at each predicted clustering granularity level against each ground-truth annotation granularity level for the Action and Location criteria of the Action-3c dataset.
+The Harmonic Mean of CAcc and SAcc is reported for each granularity pair.
+L1, L2, and L3 represent the coarse-grained, middle-grained, and fine-grained levels, respectively, for both predictions and annotations.
+Next, we quantitatively evaluated the multi-granularity grouping results at each predicted clustering granularity level against each ground-truth annotation granularity level by measuring clustering accuracy (CAcc) and semantic accuracy (SAcc).
+The main caption-based $\mathcal{X}$-Cluster framework was used for this experiment.
+In Figure 17, we report the Harmonic Mean of CAcc and SAcc for the Action and Location criteria of Action-3c, across each predicted clustering granularity level evaluated against each ground-truth annotation level.
+As clearly shown, the highest grouping performance consistently appears along the diagonal.
+This indicates that the best grouping performance is achieved when the predicted granularity matches the annotation granularity.
+These experimental results not only highlight the importance of the multi-granularity output of our framework but also validate the effectiveness of our multi-granularity design in aligning with user-preferred granularities that is reflected by the annotations in these experiments.
+Appendix K Study on Handling Invalid Criteria
+At the criteria refinement step, invalid grouping criteria (False Positives) may be proposed due to hallucinations from large language models (LLMs).
+While we did not observe hallucinated criteria being introduced during our experiments across six datasets and three application studies, it is important to further investigate the potential impact of such invalid criteria on the proposed $\mathcal{X}$-Cluster system.
+To this end, we design and conduct a control experiment using the Fruit-2c dataset (Muresan and Oltean, 2018), where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool.
+These invalid criteria were then used in the subsequent grouping process to evaluate their effect on our system.
+We apply the main Caption-based Grouper to group fruit images based on these “hallucinated” criteria.
+The grouping results for the two invalid criteria are presented in Tab.
+31.
+As observed, when processing invalid “hallucinated” criteria, nearly all images are assigned to a cluster named “Not visible” by our framework.
+This occurs because, in the absence of relevant visual content in the images, the MLLM-generated captions do not include descriptors corresponding to the invalid criteria.
+Consequently, the LLM creates a “Not visible” cluster and assigns the images to it.
+Since the system provides interpretable outputs, users can easily identify and disregard such invalid groupings.
+This control experiment highlights the robustness of our system against hallucination in practical scenarios.
+Table 31: Study of the Influence of Invalid Grouping Criteria (False Positives) on the Fruit-2c Dataset.
+We report the distribution of predicted groupings under the two “hallucinated” invalid grouping criteria.
+The main Caption-based Semantic Grouper is used for this experiment.
+†: For simplicity, all other minority clusters are grouped as “Others”.
+Predicted Clusters Action (%) Clothing Style (%)
+Not visible 98.3 96.7
+Others† 1.7 3.3
+Appendix L Study on Model Hallucination and Bias
+Model hallucination.
+LLM hallucination (Wang et al., 2024c) typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like "Who was the 70th president of the United States?" might lead to a fabricated response.
+However, in our system, the use of LLMs is fully grounded in the visual descriptions (tags or captions) of the images.
+Consequently, the LLM output is strongly constrained to analyzing these visual descriptions, significantly reducing the likelihood of hallucination.
+That said, LLM hallucination can still have mild effects on clustering results.
+For example, as discussed in the failure case analysis in Appendix I, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 16).
+MLLMs also play a crucial role in our system, as they are responsible for translating images into text for subsequent processing steps.
+MLLM hallucination (Wang et al., 2024c) typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image.
+However, since our proposed system operates at the dataset level rather than on a per-image basis, it is largely insensitive to such hallucinations, especially at the fine-grained visual detail level.
+Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker (Ge et al., 2024), which we leave as a direction for future work.
+Model Bias.
+Foundation models such as LLMs and MLLMs are known to inherit biases from their training data (Bommasani et al., 2021).
+In our system, we addressed potential biases using Hard Positive Prompting techniques: i) MLLM Bias Mitigation: The MLLM is further prompted to generate criterion-specific captions that focus solely on describing the criterion-related content in each image.
+This approach constrains the MLLM from generating irrelevant content influenced by inherent biases; ii) LLM Bias Mitigation: Similarly, when prompting the LLM to assign image captions to clusters, we condition it to concentrate exclusively on the Criterion depicted in each image (see Tab.
+17).
+To validate the effectiveness of these bias mitigation techniques, we conducted a fair clustering experiment.
+Specifically, following Kwon et al.
+(2024), we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET (Gustafson et al., 2023) dataset, which contains images from 52 occupations.
+For each occupation, we selected 10 images of men and 10 images of women, totaling 80 images, ensuring a ground-truth gender proportion disparity of 0% for each occupation.
+Using our main $\mathcal{X}$-Cluster system, we grouped these images based on the criterion Occupation using three bias mitigation strategies: i) No mitigation: using general descriptions from the MLLM for LLM grouping; ii) Our default hard positive prompting strategy: using criterion-specific captions from the MLLM for LLM grouping; and iii) Our default strategy with additional negative prompt: adding a simple negative prompt, “Do not consider gender,” to both the MLLM captioning and LLM grouping prompts.
+In this experiment, non-biased result is defined as achieving equal gender proportions within each cluster.
+Tab.
+32 presents the average gender ratios of the clustering results for each method across the four occupations.
+As observed, without bias mitigation, $\mathcal{X}$-Cluster exhibits noticeable gender bias in the studied occupations, with a gender disparity of 19.4%.
+However, our default bias mitigation techniques effectively reduce this disparity to 4.9%, achieving performance comparable to the addition of a manual negative prompt.
+This experiment demonstrates the effectiveness of our bias mitigation strategy and highlights the potential for further reducing model bias in our framework using more advanced techniques.
+Table 32: Average gender ratio and disparity across the four studied occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET dataset.
+Images sampled from each occupation have an equal proportion of genders.
+Results from different bias mitigation strategies are reported.
+Bias Mitigation Strategy Male (%) Female (%) Gender Disparity (%)
+Ground-truth 50.0 50.0 0.0
+No mitigation 40.3 59.7 19.4
+Ours (default) 47.6 52.5 4.9
+Ours w. Negative prompt 48.4 51.6 3.2
+Appendix M Computational Cost Analysis
+The proposed $\mathcal{X}$-Cluster framework is training-free, requiring only inference processes.
+Specifically, our main framework (Caption-based) requires up to 31 GB of GPU memory to operate.
+All experiments reported in the paper were conducted on 4 Nvidia A100 40GB GPUs.
+In Tab.
+33, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations.
+For these experiments, we used LLaVA-NeXT-7B (Liu et al., 2024b) as the MLLM and Llama-3.1-8B (Meta, 2024a) as the LLM.
+As shown in Tab.
+33, organizing 5,000 images based on all four discovered criteria can be completed by $\mathcal{X}$-Cluster in 29.1 hours on a single A100 GPU or 16.7 hours on a single H100 GPU.
+More importantly, most steps in our framework, such as per-image captioning and per-caption cluster assignment, are parallelizable across multiple GPUs, significantly accelerating the process.
+Therefore, when parallelizing the framework on 4 A100 or H100 GPUs, we achieve approximately a 4$\times$ speedup, reducing computational time to 7.6 hours on 4 A100 GPUs and 4.3 hours on 4 H100 GPUs.
+Table 33: Computational cost analysis on the COCO-4c benchmark (5,000 images with four criteria).
+We report the average and total time costs on various machines.
+The time costs were calculated for organizing all 5,000 images according to all the 4 criteria.
+Our main caption-based $\mathcal{X}$-Cluster framework is used in this experiment.
+Method Hardware Average time cost (sec/img) $\downarrow$ Total time cost (hrs) $\downarrow$
+$\mathcal{X}$-Cluster 1 Nvidia A100-40GB 20.9 29.1
+4 Nvidia A100-40GB 5.5 7.6
+1 Nvidia H100-80GB 12.0 16.7
+4 Nvidia H100-80GB 3.1 4.3
+Appendix N System Sensitivity Analysis of Various MLLMs and LLMs
+Figure 18: Sensitivity analysis of different MLLMs and LLMs on the six OpenSMC benchmarks.
 Top (a): We fix the LLM to Llama-3.1-8B and study the impact of different MLLMs.
@@
 The average clustering accuracy(%) across different criteria is reported on the left, while the average semantic accuracy(%) is reported on the right.
-In Fig.
-10, we perform a system-level sensitivity analysis using our default system configuration (caption-based proposer and caption-based grouper) to examine the impact of different MLLMs and LLMs on the system performance.
+In Figure 18, we perform a system-level sensitivity analysis using our default system configuration (caption-based proposer and caption-based grouper) to examine the impact of different MLLMs and LLMs on the system performance.
 Since all variants successfully propose the basic criteria in each benchmark, we report the average clustering accuracy (CAcc) and semantic accuracy (SAcc) across various criteria for comparative analysis.
-Specifically, in Fig.
-10(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024b) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b).
-Next, in Fig.
-10(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024a), and Llama-3.1-8B.
-Findings in Fig.
-10(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language.
+Specifically, in Figure 18(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024a) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b).
+Next, in Figure 18(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024b), and Llama-3.1-8B.
+Findings in Figure 18(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language.
 On the other hand, this scalability demonstrates that our system can enhance performance with more robust MLLMs, thanks to its training-free design, which ensures compatibility with any MLLM.
 Despite this, we use LLaVA-NeXT-7B as our default MLLM due to its reproducibility, being open-source and unaffected by API changes, and its capacity for local deployment, which upholds privacy by not exposing sensitive image data to external entities.
-As for the LLMs, as depicted in Fig.
-10(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM.
+As for the LLMs, as depicted in Figure 18(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM.
 Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM.
-This consistency suggests that the reasoning task for SMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
-F.2 Further Study on Fine-grained Image Collections
+This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
+Appendix O Study on Fine-grained Image Collections
 Image collections may include fine-grained grouping criteria, such as Bird species in bird photography.
-Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He & Peng, 2017).
+Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He and Peng, 2017).
 This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process.
@@
 As shown in Tab.
-30, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself.
+34, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself.
 This demonstrates the flexibility of our system, allowing future adaptations to specific application needs, such as fine-grained image collections.
-Appendix G Further Qualitative Results
-In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
-Specifically, we present example clustering results across different criteria for COCO-4c in Fig.
-11, Food-4c in Fig.
-12, Action-3c in Fig.
-13, Clevr-4c in Fig.
-14, and Card-2c in Fig.
-15.
-Additionally, we showcase example clustering results at different predicted granularity levels for COCO-4c in Fig.
-16.
-Figure 11: Example predicted clusters of COCO-4c.
-Figure 12: Example predicted clusters of Food-4c.
-Figure 13: Example predicted clusters of Action-3c.
-Figure 14: Example predicted clusters of Clevr-4c.
-Figure 15: Example predicted clusters of Card-2c.
-Figure 16: Example predicted clusters of COCO-4c at different granularities.
-Appendix H Further Failure Case Analysis
-Figure 17: Failure case analysis. We show wrongly predicted images with their ground-truth label for four clusters.
-In Fig.
-17, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
-As observed, our method frequently mis-assigns “Surfing” to the “Kayaking” cluster under the Activity criterion.
-Upon examining the intermediate criterion captions generated by the MLLM, we found that this error is largely due to the MLLM incorrectly describing a “Surfboard” as a “Kayak”.
-This highlights the importance of the MLLM’s ability to accurately describe images, as it is critical for the performance of our system.
-Potential improvements could include majority voting or model ensembling using different MLLM models.
-Another issue arises in crowded scenes.
-When multiple people are present in an image, the model consistently assigns the Mood label “Communal” to the images.
-We speculate that this occurs because, in the presence of multiple people, the model struggles to accurately determine the mood of one key individual.
-Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels.
-For example, as shown in Fig.
-17, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine.
-The “Edamame” shown in Fig.
-17 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine.
-This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
-Appendix I Further Details on Application Study
-In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in Sec.
-5 of the main paper.
-Specifically, Sec.
-I.1 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020).
-Sec.
-I.2 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria.
-Finally, Sec.
-I.3 provides additional insights from the analysis of social media image popularity.
-I.1 Further Details on Discovering and Mitigating Dataset Bias
-In this section, we provide additional evaluation results and implementation details for the application study presented in Sec. 5 of the main paper.
+Appendix P Further Details of the Application Study
+In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in § 6 of the main paper.
+Specifically, § P.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020).
+§ P.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria.
+Finally, § P.2 provides additional insights from the analysis of social media image popularity.
+P.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
+Figure 19: Bias quantification results and human evaluation for each occupation and criterion across the two studied T2I models, DALL·E3 and SDXL. The bias intensity score is reported.
+Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator).
+We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study.
+This resulted in a total of 1,800 images.
+For each occupation, we provide some examples of images generated by DALL·E3 in Figure 22, while provide some examples of images generated by SDXL in Figure 23.
+We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
+Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set.
+For this study, we utilized the mid-granularity output of our system.
+To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction.
+We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by D’Incà et al.
+(2024):
+$\mathcal{H}_{bias}^{l}=1+\frac{\sum_{c^{l}\in\mathcal{C}^{l}}\log(p(c^{l}|\mathcal{C}^{l},\mathcal{D}_{\text{Occupation}}))}{\log(|\mathcal{C}^{l}|)}$ (1)
+where $\mathcal{D}_{\text{Occupation}}$ represents the generated images for each occupation, $\mathcal{C}^{l}$ denotes the clusters discovered under the $l$-th criterion, and $p(c^{l}|\mathcal{C}^{l})$ is the probability of each cluster under the current distribution.
+The resulting score is bounded between $\mathcal{H}_{bias}^{l}\in[0,1]$, where 0 indicates no bias towards a specific cluster (concept) under the evaluated criterion, and 1 indicates that the images are completely biased towards a particular cluster (concept) (e.g., “Grey” hair color) within the current bias dimension (e.g., Hair color).
+We used the score defined in 1 to quantify the biases for each occupation across the 10 discovered grouping criteria.
+We report the bias intensity score for each occupation and each model across the 10 discovered grouping criteria in Figure 19.
+Human Evaluation Study Details: To assess the alignment between our method’s predictions and human judgments on bias detection, we conducted a user study to gather human evaluation results for the generated images.
+As shown in the questionnaire example in Figure 24, participants were presented with images generated by DALL-E3 and SDXL for each occupation and were asked to identify the bias direction (dominant class) for each of the 10 discovered criteria and rate the bias intensity on a scale from 0 to 10.
+We collected responses from 54 anonymous participants, resulting in 6 human evaluations for each occupation and each criterion.
+The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396.
+Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach.
+Detailed user study results are provided in § P.1.
+We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers.
+In Figure 19, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
+Complete Results and Additional Findings: In Figure 19, we present the detailed bias detection results for each model, occupation, and criterion, alongside human evaluation scores for reference.
+A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g., Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g., Hair color, Mood, Attire, and Accessories).
+We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022b), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs.
+However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied.
+This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
+P.2 Further Details on Analyzing Social Media Image Popularity
+Figure 20: Complete analysis of social media photo popularity on the SPID dataset.
+We display the Top Trending and Top Mainstream clusters, along with the popularity distribution of data points within these clusters across all ten discovered criteria (in Grey).
+With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems.
+Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts.
+However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored.
+In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria).
+By analyzing these interpretable results alongside image popularity metrics (e.g., number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images.
+These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
+To expand on the discussion in § 6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 20.
+Notably, we consistently observed a sharp semantic contrast between the visual elements in top trending images and those in the mainstream images across all ten criteria.
+For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
+This recurring observation reinforces the idea that viral (or trending) content tends to capture more attention, likely because it features novel, surprising, or striking visual elements.
+Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer and Gore, 2014).
+On the other hand, widely uploaded yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality.
+We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
+P.3 Confirming and Mitigating Dataset Bias
+Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024).
+As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
+Figure 21: Results of dataset bias discovery and mitigation. Worst group and average accuracies(%) are reported.
+Table 35: Debiasing Results and Comparison on CelebA.
+We use the groups discovered by $\mathcal{X}$-Cluster to train DRO and compare it with state-of-the-art debiasing methods.
+Additionally, we present DRO results using the ground-truth distribution (DRO+GT) for reference.
+Method Worst Avg.
+JTT 81.5 88.1
+CNC 88.8 89.9
+DRO+B2T 90.4 93.2
+DRO+$\mathcal{X}$-Cluster 90.9 93.1
+DRO+GT 89.7 93.6
+Findings: As expected, our method identified the grouping criteria Hair color and Gender.
+Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters.
+As shown in Figure 21, we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%).
+Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”.
+To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels.
+As shown in Tab.
+35, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
 Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset (Liu et al., 2015) to assess the classification accuracy of them.
@@
 A value of $H(A)$ or $H(Y)$ equal to 1 indicates a uniform distribution (i.e., no class imbalance).
-We then used the ground-truth distribution from the dataset’s labels and our predicted distribution to estimate the spurious correlation intensity using the score from Eq.
-1.
+We then used the ground-truth distribution from the dataset’s labels and our predicted distribution to estimate the spurious correlation intensity using the score from 2.
 For gender and hair color, our method’s predictions yielded a score of $I_{Pred}=0.10$, which is nearly identical to the ground-truth score of $I_{GT}=0.11$.
@@
 For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set.
-The final evaluation and comparison results are provided in Fig.
-7.
-I.2 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
-Figure 18: Bias quantification results and human evaluation for each occupation and criterion across the two studied T2I models, DALL·E3 and SDXL. The bias intensity score is reported.
-Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator).
-We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study.
-This resulted in a total of 1,800 images.
-For each occupation, we provide some examples of images generated by DALL·E3 in Fig.
-20, while provide some examples of images generated by SDXL in Fig.
-21.
-We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
-Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set.
-For this study, we utilized the mid-granularity output of our system.
-To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction.
-We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by D’Incà et al.
-(2024):
-$\mathcal{H}_{bias}^{l}=1+\frac{\sum_{c^{l}\in\mathcal{C}^{l}}\log(p(c^{l}|\mathcal{C}^{l},\mathcal{D}_{\text{Occupation}}))}{\log(|\mathcal{C}^{l}|)}$ (2)
-where $\mathcal{D}_{\text{Occupation}}$ represents the generated images for each occupation, $\mathcal{C}^{l}$ denotes the clusters discovered under the $l$-th criterion, and $p(c^{l}|\mathcal{C}^{l})$ is the probability of each cluster under the current distribution.
-The resulting score is bounded between $\mathcal{H}_{bias}^{l}\in[0,1]$, where 0 indicates no bias towards a specific cluster (concept) under the evaluated criterion, and 1 indicates that the images are completely biased towards a particular cluster (concept) (e.g.,, “Grey” hair color) within the current bias dimension (e.g., Hair color).
-We used the score defined in Eq.
-2 to quantify the biases for each occupation across the 10 discovered grouping criteria.
-We report the bias intensity score for each occupation and each model across the 10 discovered grouping criteria in Fig.
-18.
-Human Evaluation Study Details: To assess the alignment between our method’s predictions and human judgments on bias detection, we conducted a user study to gather human evaluation results for the generated images.
-As shown in the questionnaire example in Fig.
-22, participants were presented with images generated by DALL-E3 and SDXL for each occupation and were asked to identify the bias direction (dominant class) for each of the 10 discovered criteria and rate the bias intensity on a scale from 0 to 10.
-We collected responses from 54 anonymous participants, resulting in 6 human evaluations for each occupation and each criterion.
-The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396.
-Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach.
-Detailed user study results are provided in Sec.
-I.2.
-We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers.
-In Fig.
-18, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
-Complete Results and Additional Findings: In Fig.
-18, we present the detailed bias detection results for each model, occupation, and criterion, alongside human evaluation scores for reference.
-A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g.,, Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g.,, Hair color, Mood, Attire, and Accessories).
-We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs.
-However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied.
-This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
-I.3 Further Details on Analyzing Social Media Image Popularity
-Figure 19: Complete analysis of social media photo popularity on the SPID dataset.
-We display the viral and major popular clusters, along with the popularity distribution of data points within these clusters across all ten discovered criteria (in Grey).
-With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems.
-Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts.
-However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored.
-In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria).
-By analyzing these interpretable results alongside image popularity metrics (e.g.,, number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images.
-These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
-To expand on the discussion in Sec.
-5 of the main paper, we present the complete findings across all ten discovered criteria in Fig.
-19.
-Notably, we consistently observed a sharp semantic contrast between the visual elements in viral images and those in the majority of popular images across all ten criteria.
-For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
-This recurring observation reinforces the idea that viral content tends to capture more attention, likely because it features novel, surprising, or striking visual elements.
-Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer & Gore, 2014).
-On the other hand, widely popular yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality.
-We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
+The final evaluation and comparison results are provided in Tab.
+35.
+Appendix Q Why LLMs Improve Image Clustering?
+The most compelling aspect of this work lies in our $\mathcal{X}$-Cluster framework’s ability to transform large volumes of unstructured images into natural language and leverage the advanced text understanding and summarization capabilities of LLMs to tackle the challenging Open-ended Semantic Multiple Clustering (OpenSMC) task.
+This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain (Eklund and Forsman, 2022).
+Our core motivation is: “If LLMs can discover topics from documents and organize them, then by converting images into text, we can similarly use LLMs to organize unstructured images.”
+Traditional clustering methods (Estivill-Castro, 2002; Caron et al., 2018; Van Gansbeke et al., 2020; Li et al., 2024; Yu et al., 2024) often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable.
+These limitations hinder their applicability to diverse datasets in open-world scenarios, as they demand significant human priors and retraining for each new dataset.
+In contrast, LLMs (OpenAI, 2022a, 2023; Touvron et al., 2023; Meta, 2024b, a) excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content).
+Operating in a zero-shot (Kojima et al., 2022), interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training.
+By integrating LLMs with MLLMs (Liu et al., 2024b) into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset.
+This design allows our framework to overcome the rigid assumptions of traditional clustering methods, making it automatic, generalizable, and training-free.
+Our approach provides a novel perspective, demonstrating how clustering tasks can evolve beyond traditional paradigms.
+Challenges of employing LLMs to facilitate the SMC task.
+The main challenge of employing LLMs for the SMC task lies in accurately translating visual content from images into natural language that LLMs can effectively reason with.
+This is evident from the sensitivity analysis results in App.
+N: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 18 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 18 (b)).
+In other words, the quality of image captions generated by MLLMs is critical for the effective use of LLMs in the SMC task.
+Specifically, in the first stage of $\mathcal{X}$-Cluster (criteria proposal), captions need to be as comprehensive as possible to provide rich information for LLMs to discover grouping criteria.
+In the second stage (semantic grouping), criterion-specific captions should precisely capture relevant visual content to provide accurate information for assigning images to clusters.
+To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling (Liu et al., 2024c), or stronger models like GPT-4V (OpenAI, 2023) can improve comprehensiveness.
+For better precision, advanced prompting methods like CoT (Wei et al., 2022) or FineR (Liu et al., 2024e) can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker (Ge et al., 2024) can reduce noise caused by hallucinations.
+However, these techniques increase computational costs and framework complexity.
+In this work, we choose to keep $\mathcal{X}$-Cluster simple yet effective, and we outline these potential improvements for future practitioners.
+Appendix R Future Work
+Closed-Loop Optimization.
+In this work, we designed our prompts following the Iterative Prompt Engineering methodology (DeepLearning.AI, 2024) introduced by Isa Fulford and Andrew Ng.
+In App.
+E, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice.
+These explanations cover elements such as system prompts, input formatting, task and sub-task instructions, and output instructions.
+Our focus in this work is on creating a highly generalizable framework, $\mathcal{X}$-Cluster, and we do not perform any closed-loop, dataset-specific prompt optimizations.
+However, in future work or application scenarios where a labeled training/validation dataset is available, practitioners could build upon our design objectives.
+By leveraging our proposed evaluation metrics (see § 5) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
+$\bm{\mathcal{X}}$-Cluster on Other Data Types.
+The core idea of our proposed framework, $\mathcal{X}$-Cluster, is to use text as a proxy (or medium) for reasoning over large volumes of unstructured data, generating human-interpretable insights at scale.
+As such, $\mathcal{X}$-Cluster can be directly applied to textual data (e.g., documents).
+Moreover, since natural language is a highly versatile and widely-used medium of representation, $\mathcal{X}$-Cluster can be extended to other data types by converting these data into text (by replacing the captioning module with suitable tools) in future work, such as:
+  • Audio Data: Speech-to-Text models like Whisper (Radford et al., 2023) can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
+  • Tabular Data: Table-to-Text models, such as TabT5 (Andrejczuk et al., 2022), can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster.
+For tables containing figures, modern MLLMs like LLaVA-Next, which support both OCR and image-to-text capabilities, can handle these elements to create a unified textual representation for $\mathcal{X}$-Cluster.
+  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT (Wang et al., 2024a), can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
+  • Point Cloud Data: 3D captioning models, like Cap3D (Luo et al., 2024a), can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
+We believe the versatile nature of $\mathcal{X}$-Cluster has the potential to open up a broad range of applications across diverse data modalities, fostering new directions in future research.
 Figure 22: Samples of DALL·E3 generated images.
@@
 Figure 24: Example of the questionnaire for human evaluation study.
-Appendix J In-depth Discussion of Related Settings
-Distinction from Multiple Clustering: Multiple clustering involves finding diverse clusterings from the same dataset in either a semi-supervised (Bae & Bailey, 2006; Ren et al., 2022) or unsupervised (Caruana et al., 2006; Mautz et al., 2020; Yao et al., 2023) manner.
-However, multiple clustering methods require users to specify the number of clusterings and clusters (or rely on ground-truth annotations for evaluation) (Yu et al., 2024).
-In other words, $L$ (the number of clusterings) and $K_{l}$ (the number of clusters within each clustering) are assumed as prior knowledge in multiple clustering settings.
-This strong assumption creates a chicken-or-egg dilemma and is often impractical in real-world applications: Users want to employ multiple clustering methods to understand their data, but how can they predefine the number of valid clusterings and clusters without already having a deep understanding of the entire dataset?
-Although some strategies exist to determine the number of clusters, most of these methods (Monti et al., 2003; Zhang et al., 2017; Liu et al., 2024e) only work for single clustering.
-In contrast, SMC requires the model to automatically discover both the number of clusterings and the clusters within them.
-Additionally, while multiple clustering methods can reveal diverse sample division patterns and underlying data structures, they do not provide interpretations of the results—specifically, what rules the output clustering follows and the semantic meaning of the clusters.
-As a result, users often need to manually investigate the clustering outcomes.
-In stark contrast, SMC methods provide interpretability by describing both the clustering rules and the semantic meanings of the clusters in natural language.
-This not only offers a more comprehensive understanding of the data but also allows users to combine different clusterings to gain deeper insights into the data distribution.
-[…]
-Distinction from Multi-label Zero-shot Classification: SMC differs from multi-label zero-shot classification in that the latter (Lee et al., 2018; Huang et al., 2020; Ali & Khan, 2023; Gupta et al., 2023) requires predefined sets of classes under different rules, with the goal being to assign each image to multiple classes from these sets.
-In contrast, SMC requires discovering the class names (or cluster semantics).
-In fact, multi-label zero-shot classification can be viewed as a specific instance of SMC when the user explicitly and precisely defines all the clusters and their semantic meanings under different clustering rules.

### v5

@@
 Abstract
-Organizing unstructured image collections into semantic clusters is a long-standing challenge.
-Traditional deep clustering techniques address this by producing a single data partition, whereas multiple clustering methods uncover diverse alternative partitions, but only when users predefine the clustering criteria.
-Yet expecting users to specify such criteria a priori for large, unfamiliar datasets is unrealistic.
-In this work, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), which aims to automatically discover clustering criteria from large, unstructured image collections, revealing interpretable substructures without human input.
+In this work, we introduce and study the novel task of Open-ended Semantic Multiple Clustering (OpenSMC).
+Given a large, unstructured image collection, the goal is to automatically discover several, diverse semantic clustering criteria (e.g., Activity or Location) from the images, and subsequently organize them according to the discovered criteria, without requiring any human input.
 Our framework, $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, treats text as a reasoning proxy: it concurrently scans the entire image collection, proposes candidate criteria in natural language, and groups images into meaningful clusters per criterion.
-To evaluate progress, we release COCO-4c and Food-4c benchmarks, each annotated with four grouping criteria.
-Experiments show that $\mathcal{X}$-Cluster effectively reveals meaningful partitions and enables downstream applications such as bias discovery and social media image popularity analysis.
-We will open-source code and data to encourage reproducibility and further research.
-Organizing Unstructured Image Collections using Natural Language
-Mingxuan Liu1    Zhun Zhong4    Jun Li6    Gianni Franchi3    Subhankar Roy5    Elisa Ricci1,2 1 University of Trento   2 Fondazione Bruno Kessler  3 ENSTA Paris, Institut Polytechnique de Paris 4 Hefei University of Technology   5 University of Bergamo   6 Technical University of Munich
-Figure 1: Top: Open-ended Semantic Multiple Clustering (OpenSMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable clusters, under multiple shared themes or criteria, without requiring any prior knowledge.
-Bottom: Our system enables various applications like discovering novel biases in text-to-image (T2I) generative models.
-See our project webpage for code and data: here.
+This radically differs from previous works, which either assume predefined clustering criteria or fixed cluster counts.
+To evaluate $\mathcal{X}$-Cluster, we create two new benchmarks, COCO-4C and Food-4C, each annotated with four distinct grouping criteria and corresponding cluster labels.
+Experiments show that $\mathcal{X}$-Cluster can effectively reveal meaningful partitions on several datasets.
+Finally, we use $\mathcal{X}$-Cluster to achieve various real-world applications, including uncovering hidden biases in text-to-image (T2I) generative models and analyzing image virality on social media.
+Code and datasets will be open-sourced for future research.
+Figure 1: Organizing unstructured image collections with $\bm{\mathcal{X}}$-Cluster.
+Mid: We propose $\mathcal{X}$-Cluster, which takes any unstructured image collection as input, automatically discovers multiple criteria (e.g., Activity or Location) that meaningfully group the data, and outputs images organized into semantic clusters under each criterion, without any prior knowledge.
+We demonstrate $\mathcal{X}$-Cluster as a versatile tool for real-world analysis: i) Left - uncovers surprising novel biases (e.g., “Hair color” or “Hair style”) when applied to text-to-image (T2I) model outputs, and ii) Right - reveals visual factors (e.g., “Dramatic” color or “Intensive” emotion) that drive social media posts virality.
+footnotetext: Corresponding author.
 1 Introduction
-When grouping a large, unstructured collection of unlabeled images, several questions arise: i) Which criterion should guide the grouping?
-ii) Do multiple, equally valid criteria exist?
-iii) Given a criterion, how can we uncover the corresponding semantic clusters?
-iv) What should the clusters’ granularity be?
-Because the answers depend on both the collection’s visual content and the objectives of a specific application, each question is inherently open-ended.
-Having the ability to automatically resolve them would unlock hidden insights in the data and enable valuable applications, such as social media recommendation systems (Cheng et al., 2024) and dataset auditing (Bianchi et al., 2023).
-footnotetext: † Corresponding author.
-Previous studies have attempted to answer these questions.
-Deep Clustering (DC) approaches (Caron et al., 2018; Van Gansbeke et al., 2020) iteratively partition an unlabeled dataset but produce only a single partition.
-Multiple Clustering (MC) methods (Yu et al., 2024), instead, uncover several alternative clustering solutions, offering complementary perspectives.
-Yet both DC and MC ultimately reflect the inductive biases of the underlying networks, which can deviate from the partitions a user actually needs.
-Worse, the discovered criteria and clusters, denoted merely by index labels, lack human-interpretable semantics.
-To give users explicit control over clustering, Text Criterion conditioned Multiple Clustering (TCMC) (Yao et al., 2024, 2025) approaches let the users specify the desired criterion in natural language e.g., “group playing cards by Suit.” TCMC methods (Kwon et al., 2024; Luo et al., 2024b; Yao et al., 2024, 2025) first employ large language models (LLMs) to generate candidate cluster semantics conditioned on the user-specified criterion, then use vision–language models to assign images to semantically related clusters.
-Although TCMC methods yield user-guided and interpretable partitions, they presume the user already knows valid grouping criteria for the entire collection.
-As datasets grow and scenes become more complex (e.g., millions of social-media images (Hu et al., 2014)) knowing such prior knowledge becomes impractical.
-This leads to the question: Can clustering criteria and semantic substructures (groups) be discovered directly from visual data, without any human input?
-We formalize this challenge as the task of Open-ended Semantic Multiple Clustering (OpenSMC).
-As depicted in Figure 1(top), the objective in OpenSMC is to discover, in natural language, both the clustering criteria and their associated semantic clusters without relying on any human priors.
-The task is challenging because i) it requires concurrent reasoning over the entire image collection to surface valid criteria, posing a challenge to current vision systems, and ii) it is open-ended as neither user-defined knowledge about the clustering criteria nor clustering granularity is available.
+When organizing a large collection of unlabelled images, a natural question arises: how should we group them?
+One could imagine many possible criteria, i.e., based on Activity, Location, or even Color.
+Yet, it is often unclear which criterion, if any, best describes the dataset, or whether multiple valid grouping principles coexist.
+As a result, the seemingly simple task of clustering images becomes challenging, as it is influenced by both the visual appearance of the data and their underlying semantics.
+However, tackling this open-ended unsupervised task of automatically uncovering diverse and interpretable substructures within large image collections is pivotal for many applications, such as social media recommendation [cheng2024retrieval] and dataset auditing [bianchi2023easily].
+Existing clustering approaches still heavily rely on an iterative, human-in-the-loop interpretation and refinement process.
+Typically, we begin by setting a few hyperparameters (e.g., the number of grouping criteria or clusters) for Deep Clustering (DC) methods [caron2018deep, van2020scan] to get a single partition, or for Multiple Clustering (MC) methods [yu2024multiple] to produce several partitions showing different views of the data.
+We then inspect sample images from each cluster, hoping they correspond to meaningful categories (e.g., “Surfing” or “Skateboarding”) and, ideally, that all clusters follow a coherent criterion (e.g., Activity).
+When such patterns fail to emerge, we tweak the hyperparameters and try again until the clusters finally make sense to us.
+This labor-intensive trial-and-error loop exists because i) the resulting clusters are not directly interpretable, being represented only as index assignments.
+ii) both DC and MC methods converge to solutions shaped by model inductive biases and hyperparameter settings, rather than the data’s intrinsic semantics.
+To enhance controllability and interpretability, recent studies have introduced Text-Conditioned Multiple Clustering (TCMC) [yao2024multi, yao2025customized].
+TCMC approaches employ Multimodal Large Language Models (MLLMs) [openai2024gpt4o, liu2024llava] to generate semantic clusters based on user-defined criteria and assign images accordingly, producing human-understandable cluster labels.
+However, these approaches assume that users already know meaningful ways to organize the dataset.
+As datasets grow in size and complexity, defining such criteria becomes increasingly unrealistic.
+Moreover, by relying on static preset criteria, this paradigm may overlook previously unknown grouping dimensions that organically emerge from ever-evolving data.
+In this paper, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), the first task that aims to automatically generate open-ended and interpretable groupings of large, unstructured image collections without any human priors.
+Specifically, the goal of OpenSMC is to discover clustering criteria directly from the data and uncover their corresponding semantic clusters to organize images accordingly.
+This task is particularly challenging because i) it requires concurrent reasoning over all images to identify valid clustering criteria, and ii) it assumes no access to user knowledge about either the clustering criteria or the number of clusters.
 Tab.
-1 contrasts OpenSMC with existing clustering paradigms.
-In addition, due to the lack of practical benchmarks with diverse criteria for evaluating the OpenSMC, we introduce two realistic and larger benchmarks–COCO-4c and Food-4c–with annotations for up to four criteria.
+1 summarizes the key differences between OpenSMC and other paradigms.
 Table 1: Comparison of different clustering paradigms. Unlike DC, MC, and TCMC settings, the proposed OpenSMC task does not assume any prior knowledge and offers interpretable results.
@@
 Open-ended
-To tackle OpenSMC, we leverage text as a compact proxy for reasoning because current vision models cannot yet encode and reliably reason over large image collections.
-Consequently, we propose $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a two-stage framework that combines multimodal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2025, 2023a) with text-only LLMs (Meta, 2024a).
-$\mathcal{X}$-Cluster comprises two modules: a Criteria Proposer, which jointly reasons over all images via their textual representation (embeddings, tags, or captions) to surface valid grouping criteria, and a Semantic Grouper, which uncovers semantic clusters and assigns images to them under each discovered criterion.
-As shown in Figure 1(top), $\mathcal{X}$-Cluster can help users explore unfamiliar image collections at scale, by automatically grouping the images into semantic clusters (e.g., “Surfing” and “Skateboarding”) under diverse, human-interpretable criteria (e.g., Activity).
-In this work, we investigate three designs for each module and demonstrate the flexibility and effectiveness of our framework.
-Finally, we showcase both the practicality and versatility of $\mathcal{X}$-Cluster across several applications, including uncovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images.
-For instance, as shown in Figure 1(bottom), the method uncovers well-known biases (e.g., Gender) alongside less explored ones (e.g., Hair color) in DALL·E3 -generated images (Betker et al., 2023).
-These results confirm that $\mathcal{X}$-Cluster is an automated, practical tool for extracting interpretable insights from large-scale, unstructured visual data, opening new avenues for research and industry.
+To address OpenSMC, we make the following contributions.
+First, we introduce $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a novel training-free two-stage framework powered by MLLMs [liu2024llava, li2024llava, li2023blip] and LLMs [meta2024llama31].
+$\mathcal{X}$-Cluster consists of two consecutive modules: the Criteria Proposer and the Semantic Grouper.
+The Criteria Proposer employs a LLM to holistically reason over the entire image collection through textual representations to discover potential clustering criteria.
+For each discovered criterion, the Semantic Grouper then organizes images into distinct semantic substructures based on their criterion-related visual content.
+As shown in Figure 1(mid), our $\mathcal{X}$-Cluster automatically discovers clustering criteria (e.g., Activity, Location) and uncovers their corresponding semantic clusters (e.g., “Surfing”, “Skateboarding” under Activity), all expressed in human-interpretable natural language.
+As our second contribution, we introduce two realistic and large-scale benchmarks, COCO-4c and Food-4c, each annotated with ground-truth data for up to four clustering criteria.
+Using them, we comprehensively evaluate the effectiveness of our method in both criteria discovery and semantic grouping.
+As our third contribution, we demonstrate the versatility of $\mathcal{X}$-Cluster by applying it across diverse applications.
+When applied to occupation portrait images generated by text-to-image (T2I) generative models (Figure 1(left)), it uncovers novel occupational biases, such as DALL·E3 [betker2023improving] associating CEOs with “dark” and “short” hair, beyond well-known biases (e.g., Gender).
+When applied to social media image posts, $\mathcal{X}$-Cluster finds that images featuring “dramatic” colors, “intense” emotions, or ‘urban eclectic” clothing styles tend to attract greater popularity online.
+These findings show that $\mathcal{X}$-Cluster is a practical tool for understanding large-scale unstructured visual data, enabling the discovery of novel, unexpected patterns.
 2 Related Work
 Image Clustering.
-Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self-supervision (Caron et al., 2018, 2020; Ren et al., 2024).
-Multiple clustering extends this idea, seeking multiple non-redundant partitions with data augmentations, diversity losses, or subspace methods (Qi and Davidson, 2009; Hu and Pei, 2018; Yao et al., 2023b; Metaxas et al., 2023; Yu et al., 2024).
-Despite steady progress, both paradigms share key limitations: i) their results reflect inductive biases of the network and training algorithm, hampering generalization beyond object-centric data and often misaligning with user intent; and ii) clusters are returned as numeric indices rather than human-readable names.
+Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self‑supervision [caron2018deep, caron2020unsupervised, ren2024deep].
+Multiple clustering extends this idea, seeking multiple non‑redundant partitions with data augmentations, diversity losses, or subspace methods [qi2009principled, hu2018subspace, yao2023augdmc, metaxas2023divclust, yu2024multiple].
+Despite steady progress, both paradigms share key limitations: i) their results are shaped by model inductive biases and training algorithms, limiting generalization beyond object-centric data and often misaligning with user intent or data semantics; and ii) clusters are produced as numeric indices rather than human-readable names.
 In contrast, $\mathcal{X}$-Cluster derives both meaningful criteria and cluster names directly from unlabeled data.
 […]
-While $\mathcal{X}$-Cluster also uses text as its reasoning medium, it departs from TCMC systems in two key aspects: i) it automatically discovers the grouping criteria rather than relying on a user-supplied one; ii) it infers both the number and the names of clusters, without any user input.
-Together, these features make $\mathcal{X}$-Cluster fully automatic and scalable, and well suited to uncover novel, often overlooked insights in large image collections.
 Topic Discovery.
-OpenSMC is also related to Topic Discovery (Blei et al., 2003) in NLP, which uncovers latent themes in large text corpora.
-Our work likewise seeks to reveal common thematic concepts, but from visual collections, and does so by using LLMs as the reasoning core.
-3 OpenSMC
-Task Definition: Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering (OpenSMC) is to build a system $\mathcal{H}$ that automatically i) discovers a set of $L$ grouping criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them.
-Formally, we define an OpenSMC system:
+OpenSMC is also related to Topic Discovery [blei2003latent, wang2009multi, eklund2022topic] in NLP, which identifies latent themes [viswanathan2024large, zhang2023clusterllm, zhong2024explaining] or events [tipirneni2024context, nakshatri2023using] from text corpora.
+Our work similarly aims to uncover common themes from large, unstructured data but operates on images, which is more challenging since i) visual semantics are implicit, unlike text where meaning is explicit, and ii) no current vision model can reliably reason over large image sets.
+3 Open-ended Semantic Multiple Clustering
+Task Definition.
+Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering is to build a system, $\mathcal{H}$, that automatically i) discovers a set of $L$ grouping criteria $\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them.
+Formally, we define an OpenSMC system as:
 $\mathcal{H}:\mathcal{D}\mapsto\left\{\mathcal{O}_{l}=\Bigl\{\mathcal{C}_{k}^{l}=\bigl(s_{k}^{l},\mathcal{D}_{k}^{l}\bigr)\Bigr\}_{k=1}^{K_{l}}\,\middle|\,R_{l}\right\}_{l=1}^{L},$
+where each cluster $\mathcal{C}_{k}^{l}$ is characterized by a semantic name $s_{k}^{l}$ and a subset of images $\mathcal{D}_{k}^{l}\subset\mathcal{D}$ that share the same semantics.
+A criterion $R_{l}$ refers to a theme for grouping images, such that all the clusters under $R_{l}$ should align with the theme.
+As shown in Figure 1(top), if $R_{l}=\texttt{Activity}$, each cluster under this criterion should collect images $\mathcal{D}_{k}^{l}$ that depict an activity, such as $s_{k}^{l}=\text{``Surfing''}$.
+If $R_{l}=\texttt{Location}$, the same dataset should be organized into clusters like “Restaurant”, “Sports facility”, and so on.
+An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language.
+In contrast to TCMC setting, where criteria $\mathcal{R}$ and the corresponding cluster counts $K$ are preset by human operators.
+Figure 2: OpenSMC benchmarks. We introduce two new challenging benchmarks: COCO-4c and Food-4c. We show all annotated criteria and the corresponding labels for the example images.
+Figure 3: $\bm{\mathcal{X}}$-Cluster consists of a Criteria Proposer and a Semantic Grouper.
+(left) Given a set of images, the Proposer discovers and outputs a pool of grouping criteria in natural language.
+(right) The Grouper subsequently extracts criterion-specific descriptions from images relevant to each criterion, discovers the underlying semantic clusters, and groups each image at three semantic granularity levels.
+Results shows an example, as how an unstructured image collection can be grouped into clusters of different semantic granularity corresponding to criterion “Location”.
+See Supp.
+F for implementation and prompt details.
+Benchmark.
+Evaluating OpenSMC methods requires benchmarks that can be partitioned under multiple criteria.
+Currently, only a few benchmarks [yu2024multiple] support the evaluation of OpenSMC methods: Fruit-2c [muresan2018fruit], Card-2c [kaggle2022card], Action-3c [kwon2023image], and Clevr-4c [vaze2024no].
+As shown in Figure 2, these benchmarks are limited by their object-centric nature with simple backgrounds (e.g.
+Fruit-2c), an insufficient number of criteria (e.g.
+up to three in Action-3c), and a lack of photorealism due to synthetic generation (e.g.
+Clevr-4c).
+Given that the data encountered in real-world applications is more complex, we annotate and propose two new benchmarks for OpenSMC: Food-4c and COCO-4c.
+Food-4c is sourced from Food-101 [bossard2014food], which includes 101 Food type (original annotations), along with new annotations for 15 Cuisine types, 5 Courses types, and 4 Diet preferences, totaling four clustering criteria.
+Additionally, we introduced COCO-4c using images from COCO-val [lin2014microsoft], where we annotated four criteria with varying number of clusters: 64 Activity, 19 Location, 20 Mood, and 6 Time of day.
+Examples of these newly constructed benchmarks are shown in Figure 2.
+Further details, such as cluster names and the annotation pipeline, are provided in Supp.
+D.
+4 Method
+The goal of an OpenSMC system is to first discover meaningful grouping criteria (or themes) from an unstructured image collection by finding commonalities among the images, and then group them into semantic clusters as per the discovered criteria.
+This is particularly a challenging task because it requires reasoning over the visual content of all images simultaneously.
+To address OpenSMC, we diverge from representation learning-based MC approaches [yao2024multi, yao2025customized], as no existing model can yet encode large image sets and reason over them reliably.
+Instead, we convert the visual content of all images into text and use text descriptions as a proxy to discover the grouping criteria and the semantic substructures.
+System Overview: As illustrated in Figure 3, our proposed $\mathcal{X}$-Cluster is a two-stage framework that is composed of two modules: Criteria Proposer and Semantic Grouper.
+The Criteria Proposer processes the entire image set $\mathcal{D}$ to discover diverse common themes among the images and proposes grouping criteria $\mathcal{R}$ in natural language (e.g.
+Location).
+Once the criteria are proposed, the Semantic Grouper uncovers the substructure $\mathcal{O}_{l}$ of $\mathcal{D}$ by discovering distinct semantic clusters and assigning images to their respective clusters (e.g.
+“Tennis Court”), adhering to each criterion $R_{l}\in\mathcal{R}$.
+As the OpenSMC task operates without user priors to guide semantic granularity, we also design our Semantic Grouper to automatically discovers $\mathcal{O}_{l}$ across multiple granularity levels, from coarse (e.g., “Outdoor”) to mid (e.g., “Recreation”) to fine (e.g., “Tennis Court”), and organizes the images accordingly.
+In this work, we explore three design choices for both the Proposer and the Grouper.
+Due to space constraints, only the main variant of $\mathcal{X}$-Cluster is illustrated in Figure 3, while the illustrations of the alternative variants and additional implementation details, including the exact prompts, are provided in Supp.
+F.
+Next, we describe each variant in detail.
+4.1 Criteria Proposer
+As shown in Figure 3(left), the Proposer takes as input a set of input images and generates distinct grouping criteria (or Criteria Pool) in natural language.
+Its core design principle is the ability to concurrently reason across a large set of images.
+Next, we explore three systematic approaches.
+dlgCaption-based Proposer (main):
+To enable reasoning over a large image set for criterion discovery, we first leverage a MLLM [liu2024llava] to generate a comprehensive caption $e_{n}$ for each image, converting its visual content into text representations $e_{n}=\text{MLLM}(\mathbf{x}_{n})$.
+The resulting caption set $\{e_{n}\}_{n=1}^{N}$ serves as a rich and holistic semantic proxy for the image collection $\mathcal{D}$.
+Using these compact textual proxies, we then prompt a LLM [meta2024llama31, openai2024gpt4o] to jointly analyze the aggregated visual content and propose multiple valid grouping criteria, denoted as $\widetilde{\mathcal{R}}=\text{LLM}(\{e_{n}\}_{n=1}^{N})$.
+As an example, the LLM could use different cues such as “Tennis”, “grassy field”, “rock wall” in the captions (see Figure 3) and its reasoning capability to discover the criterion “Location”, since they are usually associated to a physical location, albeit locations may not explicitly appear in any caption.
+dlgTag-based Proposer (alternative):
+Instead of using captions as textual proxies for reasoning, we further explore an approach that relies on image tags.
+Specifically, using the WordNet [miller1995wordnet] vocabulary as the candidate tag set, we employ an open-vocabulary tagger (e.g., CLIP [radford2021learning]) to assign ten tags $t_{i,n}$ to each image as $\{t_{i,n}\}_{i=1}^{10}=\text{Tagger}(\mathbf{x}_{n},\text{WordNet})$.
+These tags act as concise semantic descriptors that summarize the key elements present in each image.
+We then aggregate all assigned tags and prompt a LLM to analyze them jointly and propose grouping criteria as $\widetilde{\mathcal{R}}=\text{LLM}(\{\{t_{i,n}\}_{i=1}^{10}\}_{n=1}^{N})$.
+dlgImage-based Proposer (alternative):
+Lastly, we explore an approach that reasons directly over images rather than their textual proxies.
+Since no existing model can reliably encode large image sets at once, we adopt a simple workaround: divide $\mathcal{D}$ into 64-image batches, stitch each batch into an $8\times 8$ grid as a single composite image, and feed the resulting image grids to a MLLM.
+The model is prompted to propose grouping criteria from each grid, and we aggregate and deduplicate these subset proposals to obtain the final criteria set $\widetilde{\mathcal{R}}$.
 […]
-An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language.
-By contrast, in the TCMC setting humans predefine the criterion $\mathcal{R}$ and the desired cluster count $K$.
-Figure 2: OpenSMC benchmarks. We introduce two new challenging benchmarks: COCO-4c and Food-4c. We show all annotated criteria and the corresponding labels for the example images.
-Refer to caption(Prompts in App. E.1)(Prompts in App. E.2)
-Figure 3: $\bm{\mathcal{X}}$-Cluster overview.
-(Left) The Criteria Proposer reason over the entire image set via their textual proxies and deposits diverse, natural language grouping criteria in a Criteria Pool.
-(Middle) The Semantic Grouper draws each criterion from the pool, discovers semantic clusters at three levels of granularity, and assigns images to their proper clusters.
-(Right) Aggregating these assignments reveals the collection’s semantic substructure across multiple criteria and granularities.
-We evaluate three design variants for each module and adopt the best-performing configuration (marked Refer to caption) as the default $\mathcal{X}$-Cluster.
-See App.
-E for implementation details.
-Benchmark: Only a few benchmarks exist that are suitable for evaluating OpenSMC.
-As shown in Figure 2, current benchmarks: Fruit-2c (Muresan and Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), Clevr-4c (Vaze et al., 2024) are limited because they are object-centric (Fruit-2c), lacking photorealism (Clevr-4c), and too few criteria (e.g., at most three in Action-3c).
-To alleviate this issue, we introduce two new benchmarks: Food-4c and COCO-4c.
-Examples of images are presented in Figure 2.
-We extended Food-101 dataset (Bossard et al., 2014) to create Food-4c which in total has four grouping criteria.
-Similarly, we extend COCO-val (Lin et al., 2014) to create COCO-4c that has four grouping criteria.
-We believe that the newly introduced more realistic benchmarks will play a critical role in evaluating OpenSMC systems.
-Further benchmark statistics and annotation details are provided in App.
-C.
-4 Method
-To recap, the objective of an OpenSMC system is twofold: i) discover meaningful grouping criteria, or themes, in an unstructured image collection, and ii) group the images into semantic clusters for each discovered criterion.
-This task is particularly challenging because it requires simultaneous reasoning over a large number of images, which the current vision systems are not equipped to handle.
-System overview.
-We propose $\bm{\mathcal{X}}$-Cluster, a two-stage framework that is designed to perform joint reasoning over an entire image collection, as required by the OpenSMC task.
-At a high level, $\mathcal{X}$-Cluster addresses this task by first translating the visual content of the image collection into a lower-dimensional representation, i.e., natural language descriptions.
-We then exploit the large context window of the LLMs/MLLMs to discover the grouping criteria, followed by their semantic clusters, using the derived textual descriptions.
-Concretely, as shown in Figure 3, $\mathcal{X}$-Cluster comprise: (1) a Criteria Proposer that is in charge of discovering the grouping criteria (e.g., Activity, Mood); and (2) a Semantic Grouper that finds semantic clusters by conditioning the predictions on each discovered criterion (e.g., for a given image “Climbing gym” is predicted when conditioned on the criterion Activity, whereas “Adventurous” is predicted when conditioned on Mood).
-Next, we elaborate on each component of $\mathcal{X}$-Cluster.
-4.1 Criteria Proposer
-We explore three Proposer designs that convert visual information into text, enabling the LLM to infer grouping criteria.
-Each design differs in how we convert the visual content into text and in how we use the text to derive the grouping criteria.
-Image-based Proposer: We prompt a MLLM (Li et al., 2025) to directly infer grouping criteria $\mathcal{R}$ from images, bypassing the translation of visual content into text.
-In detail, we present a mini-batch of images, tiled into a composite of $8{\times}8$ image grid, to the MLLM.
-This allows the MLLM to reason concurrently over a group of images.
-The image-based approach balances broader coverage of criteria with the preservation of fine-grained details, a trade-off introduced by the tiling operation.
-Tag-based Proposer.
-This variant adopts a tag-driven strategy.
-We use an open-vocabulary tagger (Radford et al., 2021) to assign up to ten WordNet (Miller, 1995) tags to each image, yielding concise textual descriptors.
-A LLM (Meta, 2024a) then summarizes all the aggregated tags in $\mathcal{D}$ to propose the grouping criteria $\mathcal{R}$.
-Caption-based Proposer (Main).
-While image tags capture certain visual semantics, they primarily focus on object-related content.
-To incorporate richer visual cues, such as scene context and interactions, we use a MLLM (Liu et al., 2024b) to caption every image in $\mathcal{D}$.
-We then split the captions into the largest chunks that fit the LLM’s context window (e.g.128k or 1M tokens) (Meta, 2024a; OpenAI, 2024) and prompt the LLM on each chunk to propose grouping criteria.
-The deduplicated criteria from all chunks form the pool $\mathcal{R}$.
 4.2 Semantic Grouper
-We explore three Grouper designs that partition the dataset $\mathcal{D}$ into semantic clusters by conditioning the partition on each criterion $R_{l}\in\mathcal{R}$.
-The end result is a partition (or substructure) $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\}_{k=1}^{K_{l}}$, where an image is not only assigned to a cluster, but also has an associated semantic name (e.g., “Climbing gym” instead of a cluster index), making it human interpretable.
-Each design differs in how we condition the predictor to incorporate the grouping criterion.
-Since data can be clustered at varying granularity, we design the Grouper to cluster images at three granular levels: coarse, middle, and fine.
-As an example, under the Cuisine criterion, the Grouper can cluster $\mathcal{D}$ at a coarse continental level (“European”, “Asian”), a middle regional level (“Mediterranean”, “Southeast Asian”), or a fine national level (“Italian”, “Thai”).
-Image-based Grouper.
-We task an LLM to generate a question $q_{l}$ conditioned on $R_{l}$.
-For e.g., for $R_{l}=\leavevmode\nobreak$Mood, the $q_{l}=$ “What mood does this image convey?
-Return an abstract, common, and specific category.” We then query a visual question answering (VQA) model (Li et al., 2023a) model to output a coarse-, mid-, and fine-grained label for every image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$.
-Subsequently, aggregating these labels over $\mathcal{D}$ at each level yields multi-granularity substructures.
-Tag-based Grouper: We first prompt a LLM to generate a list of middle-grained tags (e.g., “Recreational facility”) conditioned on a given criterion $R_{l}$.
-Following Liu et al.
-(2024d), we then query the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag, obtaining coarse- and fine-grained tags.
-Given the vocabulary of tags, we use CLIP (Radford et al., 2021) to tag an image with the most likely label, one at each granular level.
-Caption-based Grouper (Main): We condition a MLLM with a given criterion $R_{l}$ to generate a set of captions for each image as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$.
-Next, we use a LLM in a three-step process to group images to clusters at multiple granularity levels: i) Initial Naming: The LLM assigns a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels as $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names each granularity level as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$.
+Each discovered criterion $R_{l}\in\mathcal{R}$ serves as a thematic indicator for a distinct semantic substructure $\mathcal{O}_{l}$ within the image set $\mathcal{D}$.
+To uncover these substructures, as shown in Figure 3(right), the Grouper takes $\mathcal{D}$ and each criterion $R_{l}$ as inputs, discovers cluster names $\{s_{k}^{l}\}_{k=1}^{K_{l}}$, and groups images $\mathcal{D}_{k}^{l}$ to their corresponding clusters.
+As a result, the interpretable substructure $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl(s_{k}^{l},\mathcal{D}_{k}^{l}\bigr)\}_{k=1}^{K_{l}}$ emerges for each $R_{l}$.
+The core design of the Grouper focuses on aligning semantic substructure discovery with the given partitioning criterion.
+Like the Proposer, we explore three distinct approaches for the Grouper.
+Furthermore, as clusters under a given criterion can be formed at varying semantic granularities based on user preferences, we have designed our Grouper to clusters $\mathcal{D}$ at three levels: coarse, middle, and fine-grained.
+This allows $\mathcal{X}$-Cluster to provide insights at different granularities.
+For example, under the Cuisine criterion, $\mathcal{X}$-Cluster can organize images at a coarse continental level (e.g., “European” or “Asian”), a middle regional level (e.g., “Mediterranean” or “Southeast Asian”), or a fine national level (e.g., “Italian” or “Thai”).
+See Supp.
+F.2 for design details.
+dlgCaption-based Grouper (main):
+Given a target criterion $R_{l}$, we prompt the MLLM to generate criterion-specific captions that focus exclusively on the visual content relevant to $R_{l}$ for each image, as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$.
+Next, we design a Multi-granularity Group Assignment (MGA) module that uses the LLM to group images into clusters across multiple semantic granularity levels through a three-step process: i) Initial Naming: The LLM assigns a provisional class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels: $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names at different granularity levels as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l})$.
+By aggregating these cluster assignments across $\mathcal{D}$ at different levels, we derive multi-granularity semantic substructures.
+As we will show in § 5.3, the Caption-based Grouper outperforms other alternatives, making it our main method.
+dlgTag-based Grouper (alternative):
+Given a target criterion $R_{l}$, we prompt the LLM to generate a set of common categories (e.g., “Commercial Space”) related to the criterion $\mathcal{S}_{\text{mid}}^{l}=\text{LLM}(R_{l})$ as the mid-grained tags.
+Following liu2024shine, we further query the LLM to infer potential super- and sub-categories (e.g., “Indoor” and “Restaurant”) for each mid-grained tag, thereby obtaining the corresponding coarse- and fine-grained tag sets, $\mathcal{S}_{\text{coarse}}^{l}$ and $\mathcal{S}_{\text{fine}}^{l}$.
+Finally, we employ an open-vocabulary image tagger [radford2021learning] to assign the most relevant tag at each granularity level to each image, $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{Tagger}(\mathbf{x}_{n},\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l})$, yielding multi-granularity substructures after aggregation.
+dlgImage-based Grouper (alternative):
+Given a target criterion $R_{l}$, we first prompt a LLM to generate a question $q_{l}$ tailored to $R_{l}$, For e.g., for the criterion Mood the generated question is: “What mood is conveyed by this image?
+Answer with an abstract, common, and specific category name, respectively”.
+We then use $q_{l}$ to guide a visual question answering (VQA) model [li2023blip] in directly inferring semantic cluster names and assignments for each image at different granularity levels as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$.
 5 Experiments
+5.1 Experimental Protocol
 […]
-Evaluation metric for criteria discovery.
-We assess criterion discovery performance of the proposers using the True Positive Rate (TPR) (Csurka et al., 2024), defined as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, i.e., the fraction of ground-truth criteria $\mathcal{Y}$ recovered by the predicted set $\mathcal{R}$.
-Because the space of plausible criteria is inherently open-ended and subjective, False Positives are hard to define, so we use TPR as our primary metric: higher scores indicate more complete coverage of the reference annotated criteria.
-Evaluation metrics for substructure uncovering.
-We measure how well each criterion-specific substructure predicted by the Grouper matches the ground truth along two complementary axes: i) Semantic consistency: We quantify the semantic agreement between predicted cluster names and ground-truth labels by computing their cosine similarity with Sentence-BERT (Reimers and Gurevych, 2019) for each image; the dataset-level mean is reported as Semantic Accuracy (SAcc) (Conti et al., 2023; Liu et al., 2024e).
-ii) Structural consistency: We quantify the structural agreement between predicted and ground-truth clusters with clustering accuracy (CAcc) (Vaze et al., 2022), which computes Hungarian-matched accuracy between the two partitions (Kuhn, 1955).
-Because ground-truth granularity is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for evaluation.
-Unlike TCMC methods, which use the true cluster count during both training and evaluation, our protocol offers a fair and practical benchmark for OpenSMC.
-Figure 4: Comprehensiveness Comparison of Criteria Proposers: TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart.
+Evaluation Metric for Criteria Discovery: We use True Positive Rate (TPR) [csurka2024could] to evaluate the criteria discovery performance of different proposers.
+Specifically, we compute TPR as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, measuring to what extent the predicted set covers the ground-truth criteria $\mathcal{Y}$.
+It is important to note that the number of grouping criteria is subjective and can be as extensive as one’s preferences allow (open-ended), making False Positives hard to define.
+Thus, we use TPR as the primary metric.
+A higher TPR means better coverage of predicted criteria compared to the ground truth.
+Evaluation Metrics for Substructure Uncovering: To assess each criterion-specific substructure uncovered by the Grouper, we evaluate its alignment with the ground-truth substructure along two dimensions: i) Semantic Consistency: For each image, we compute the semantic similarity between its assigned cluster name and the ground-truth label under the current criterion using Sentence-BERT.
+The average similarity across the dataset, reported as Semantic Accuracy (SAcc) [liu2024democratizing], measures how well the predicted substructure aligns semantically with the ground truth.
+ii) Structural Consistency: We use clustering accuracy (CAcc) [han2021autonovel, vaze2022generalized] to measure the degree of structural match between the predicted and ground-truth substructures (clusters) using Hungarian matching algorithm [kuhn1955hungarian].
+Since the granularity of ground-truth annotations is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for assessment.
+Unlike TCMC methods [kwon2023image, yao2025customized] that rely on ground-truth cluster counts for perfect matching, our strategy provides a fair and practical evaluation for open-ended OpenSMC systems.
+Figure 4: Comprehensiveness Comparison of Criteria Proposers.
+TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart.
 Each block represents one ground-truth criterion, with Colored blocks indicating successfully discovered criteria and Gray blocks representing undiscovered criteria.
 5.2 Study of the Criteria Proposer
-Expanding ground-truth criteria for rigorous evaluation.
-The annotated basic criteria may not capture all valid groupings in complex datasets such as COCO-4c.
-We engaged human annotators to expand additional criteria for the six benchmarks in § 3, obtaining {10, 4, 11, 7, 17, 11} criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}, respectively.
-We call the original per-image annotated criteria (Figure 2) the Basic ground truth and the enriched set the Hard ground truth.
-Full annotation details appear in App.
-C.2.
-Which Proposer performs best?
-Figure 4 compares the TPR performance for all Proposer variants.
-The Caption-based Proposer recovers the most comprehensive set of criteria, ranking closest to the human-annotated reference set on every benchmark and on both Basic and Hard splits; in the latter it outperforms the runner-up Tag-based Proposer by $+32.2\%$ TPR.
-Captions offer richer, nuanced cues that help the LLM discover grouping criteria more comprehensively, whereas tags provide little context and therefore struggle on complex datasets (e.g., COCO-4c).
-The Image-based Proposer lags behind, as reasoning over small image grids omits fine visual cues and obscures overarching themes.
-Impact of image quantity on criterion discovery.
-Figure 5 shows the TPR performance of the Caption-based Proposer as the number of input images for criteria discovery varies.
-Interestingly, for object-centric datasets (Card-2c, Clevr-4c), only a handful of samples – even a single image – recover most criteria, thanks to their uniform structure, i.e., seeing one playing card is enough to suggest criteria like Suit and Rank in Card-2c.
-In contrast, performance drops sharply on the more diverse and realistic datasets (COCO-4c, Food-4c, Action-3c) when fewer images are provided, because inferring their richer and intricate thematic criteria requires wider visual coverage.
-Since $\mathcal{X}$-Cluster assumes no dataset priors, we use all images by default.
+We also evaluate the performance of our design for the Proposer module.
+To properly assess his effectiveness, we realize that for complex datasets like COCO-4c, four ground-truth criteria may not cover all valid grouping options.
+Therefore, we expanded the ground-truth criteria for each of the six benchmarks in Sec.
+3 using human annotators, resulting in {10, 4, 11, 7, 17, 11} distinct criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}.
+We refer to the original per-image annotated criteria set (see Figure 2) as Basic ground truth and the expanded set as Hard during evaluation.
+See Supp.
+D.2 for annotations.
+Which Criteria Proposer Performs the Best?
+In Figure 4, we compare different approaches for the Proposers in terms of the comprehensiveness of the discovered criteria using TPR.
+From Figure 4, we observe that our caption-based Proposer discovers the most comprehensive criteria, making it the closest to the human-annotated set among all methods.
+It consistently outperforms other variants in both the Basic and Hard sets across all six benchmarks.
+Its superior performance is particularly evident under the Hard criteria set, where it surpasses the second-best Tag-based Proposer by +32.2% TPR.
+Intuitively, the Caption-based Proposer works better because captions capture more diverse and nuanced aspects of the image set, which further guides the LLM to comprehensively discover different grouping criteria.
+Contrarily, the Tag-based Proposer is less effective in complex benchmarks (e.g.
+COCO-4c and Action-3c) since tags provide less contextual and descriptive information.
+Similarly, the Image-based Proposer is subpar in terms of performance since it is limited to reasoning over a small subset of images and loses visual details when combining images into a grid.
 Figure 5: Impact of Image Quantity on Criteria Discovery. We evaluate the TPR performance of the Caption-based Proposer at different image scales against the Hard ground-truth criteria set.
 Figure 6: Comparison of Semantic Groupers.
-We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across six benchmarks.
-CLIP zero-shot classification (Refer to caption) serves as a pseudo upper bound, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline.
-Best performers are marked with Refer to caption.
-See App.
-G.2 for expanded numerical results and App.
-H for clustering visualizations.
+We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across four benchmarks.
+CLIP zero-shot classification (Refer to caption) serves as an oracle, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline.
+The dlgbest performer for each criterion, determined by HM, is highlighted in green.
+Overall, our Caption-based Grouper performs best, ranking first in 10 of 15 evaluated criteria.
+See Supp.
+H.2 for clustering visualizations.
+Impact of Image Quantity on Criteria Discovery: Figure 5 shows the TPR performance of the Caption-based Proposer across different image scales.
+Interestingly, in object-centric benchmarks like Card-2c and Clevr-4c, satisfactory performance is achieved with just a few images.
+In fact, even a single image often suffices for reasonable criteria discovery, as object-centric datasets tend to have uniform structures, i.e., seeing one playing card is enough to suggest criteria like Suit.
+However, this does not hold for more complex datasets like COCO-4c, Food-4c, and Action-3c, which feature diverse and realistic scenarios.
+Here, reducing the number of images leads to a clear drop in TPR performance, as capturing intricate and varied thematic criteria requires a larger image set.
+Since $\mathcal{X}$-Cluster operates without prior knowledge of the dataset, we use the entire dataset by default to ensure comprehensive discovery.
+5.3 Study of the Semantic Grouper
 Table 2: Comparison with TCMC methods.
@@
 MMaP and MSub do not build semantic clusters.
-See expanded results in App.
-G.3.
+See expanded results in Supp.
+H.3.
 COCO-4c Food-4c Clevr-4c Action-3c Card-2c Fruit-2c Avg
@@
 UB 40.1 60.6 64.1 80.2 56.7 72.5 79.8 82.3 41.4 66.9 69.4 88.3 50.2 64.4
-MMaP  (Yao et al., 2024) 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 48.2 -
-MSub  (Yao et al., 2025) 36.0 - 47.3 - 72.2 - 64.3 - 39.6 - 54.4 - 52.3 -
+MMaP [yao2024multi] 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 48.2 -
+MSub [yao2025customized] 36.0 - 47.3 - 72.2 - 64.3 - 39.6 - 54.4 - 52.3 -
 […]
-SSD-LLM  (Luo et al., 2024b) 41.6 52.1 47.5 55.5 54.8 37.6 78.1 52.9 67.3 76.3 62.0 46.8 58.6 53.6
+SSD-LLM [luo2024llm] 41.6 52.1 47.5 55.5 54.8 37.6 78.1 52.9 67.3 76.3 62.0 46.8 58.6 53.6
 $\mathcal{X}$-Cluster (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 61.8 62.3
-5.2 Study of the Semantic Grouper
-Which Semantic Grouper performs best?
-In Figure 6, we score each Grouper with CAcc (structural agreement) and SAcc (semantic agreement) evaluated on the ground-truth per-criterion annotation, ranking models by their harmonic mean Harmonic Mean (HM) of CAcc and SAcc.
-To gauge headroom we include a pseudo upper bound: CLIP ViT-L/14 in a zero-shot regime where the true criterion, cluster names and count are all supplied.
-We also report $k$-means baselines on features from CLIP-L/14, DINOv1-B/16 (Li et al., 2023b), and DINOv2-G/14 (Oquab et al., 2024), evaluated with CAcc and given the oracle $k$.
-Figure 6 shows the Caption-based Grouper performs best.
-It ranks first on 12/19 criteria and averages $59.9\%$ CAcc, virtually matching the upper bound’s $58.1\%$.
-Its mean SAcc is $60.5\%$, again best among learned methods but below the upper-bound $74.2\%$, which benefits from exact ground-truth names.
-This gap is expected given lexical variability, e.g., “Joyful,” “Happy,” and “Cheerful” all describe the Mood criterion but lack full semantic equivalence.
+Which Semantic Grouper Performs the Best?
+In Figure 6, we evaluate different design choices for the Grouper using CAcc and SAcc for each criterion, determining the best performer based on Harmonic Mean (HM).
+To contextualize performance, we establish an oracle using CLIP ViT-L/14 in a zero-shot classification setup, where grouping criteria, cluster names, and the number of clusters are all known.
+We also use KMeans with ground-truth cluster numbers and visual features from CLIP-L/14, DINOv1-B/16 [li2022emergent], and DINOv2-G/14 [oquab2023dinov2] as CAcc baselines.
+From Figure 6, we observe that the proposed Caption-based Grouper performs best, ranking first in 10 out of 15 tested criteria based on the HM across four benchmarks.
+It achieves an average CAcc of 59.9%, closely matching the oracle performance of 58.1%, highlighting the effectiveness of our text-driven approach.
+For SAcc, the Caption-based Grouper achieves an average of 60.5%, surpassing its counterparts, but falling short of the oracle 74.2% which benefits from exact ground-truth class names.
+This gap is expected due to the open nature of the semantic space, i.e., terms like “Joyful”, “Happy”, and “Cheerful” often describe the same Mood but lack full semantic equivalence.
 The BLIP-2 Image-based Grouper ranks second.
-Its criterion-specific questions help, yet per-image answers add noise.
-The Tag-based Grouper trails across all benchmarks, reflecting the weak alignment between generated tags and dataset concepts.
-Effect of multi-granularity refinement.
-We ablate the Caption-based Grouper with three cluster naming schemes to study the effect of our multi-granularity design: i) Initial Names, where the initially LLM-assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and our iii) Multi-Granularity Refinement strategy.
-As Figure 7 indicates, either refinement method markedly boosts accuracy over raw initial names, confirming the need for consistent granularity.
-Furthermore, our multi-granularity strategy outperforms the flat one by offering several detail levels, providing greater flexibility in aligning with user-preferred grouping granularity.
-Figure 7: Study of multi-granularity refinement.
+Its criterion-specific questions improve labeling accuracy, but per-image predictions can introduce noise in clustering.
+Figure 7: Ablation study of multi-granularity refinement.
+Necessity of Multi-Granularity Cluster Refinement: To evaluate the effectiveness of multi-granularity cluster refinement design, we conduct controlled experiments using our Caption-based Grouper with three cluster naming strategies: i) Initial Names, where the initially assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and iii) Multi-Granularity Refinement, our proposed approach.
+As shown in Figure 7, both refinement methods significantly improve clustering accuracy compared to using noisy initial names, highlighting the importance of granularity-consistent cluster names for revealing substructures.
+Moreover, our multi-granularity refinement outperforms flat refinement by enabling clustering at different levels of detail, providing greater flexibility in aligning with user-preferred grouping granularity.
 5.4 Comparison with TCMC Methods
 […]
-Additional ablations and studies of $\bm{\mathcal{X}}$-Cluster.
-In the technical appendix, we provide extensive ablations and follow-up experiments: i) Cluster visualizations and qualitative analysis (App.
-H); ii) Failure case analysis (App.
-I); iii) Handling of hallucinated criteria (App.
-K); iv) Mitigating model biases (App.
-L); v) Computational cost analysis (App.
-M); vi) Sensitivity analysis of different MLLMs and LLMs (App.
-N); vii) Further investigation on multi-granularity clustering (App.
-J); and (viii) Fine-grained criterion handling (App.
-O).
-Figure 8: Bias Discovery in T2I-Generated Images.
-Bias intensity, dominant clusters, and example images are shown for few occupations.
-Full results for all studied occupations are provided in App.
-P.1.
+Further Analysis of $\bm{\mathcal{X}}$-Cluster is provided in the supplementary material: i) Supp.
+I presents qualitative results; ii) Supp.
+J examines failure cases; iii) Supp.
+L explores how $\mathcal{X}$-Cluster handles invalid (hallucinated) criteria; iv) Supp.
+M investigates model biases; v) Supp.
+N analyzes computational costs; vi) Supp.
+O studies system sensitivity to different MLLMs and LLMs; vii) Supp.
+K further investigates the impact of multi-granularity clustering; and viii) Supp.
+P explores improvements for handling fine-grained criteria.
+Figure 8: Bias Discovery in T2I-Generated Images. Bias intensity, dominant clusters, and example images are shown for few occupations.
 6 Applications
-In this section, we apply the best caption-based $\mathcal{X}$-Cluster on three applications to showcase its ability to surface novel, human-interpretable insights.
-Discovering Novel Bias in Text-to-Image (T2I) Diffusion Models: Do T2I models encode biases beyond the well-studied gender and race stereotypes?
-(Naik and Nushi, 2023) To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) and generated 100 images per occupation via the prompt “A portrait photo of a <OCCUPATION>” using DALL-E3 (Betker et al., 2023) and SDXL (Podell et al., 2024), resulting in 1.8k images.
+We apply $\mathcal{X}$-Cluster to three applications, demonstrating its ability to generate novel, human-interpretable criteria for real-world analysis.
+Below, we present results for the first two applications, while additional results for the third application on confirming and mitigating gender bias in 162k CelebA [liu2015deep] images are provided in Supp.
+Q.3.
+6.1 Discovering Biases in T2I Diffusion Models
+Do T2I models exhibit biases beyond the widely studied ones, such as gender and racial stereotypes?
+[naik2023social, nicoletti2023bias] To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies [bianchi2023easily, bolukbasi2016man] and generated 100 images per occupation using the prompt “A portrait photo of a <<OCCUPATION>>” with DALL-E3 [betker2023improving] and SDXL [podell2023sdxl], resulting in 1.8k images.
 Applying $\mathcal{X}$-Cluster, we automatically identified 10 grouping criteria (bias dimensions) and their distributions for each occupation.
-To quantify bias, we measured the normalized entropy of each dimension’s distribution (D’Incà et al., 2024) as bias intensity and marked the dominant (largest) cluster as the potential bias direction.
-We conducted a user study with 54 participants to validate our results: $\mathcal{X}$-Cluster’s predicted bias intensity closely matched ratings from the interviewees with an Absolute Mean Error of 0.1396 on a 0–1 scale, and bias directions agree 72.3% of the time with the interviewees.
-See App.
-P.1 for user study details.
-Findings: As shown in Figure 8, our method identifies both well-known and previously unnoticed novel biases without relying on predefined categories.
-For SDXL, roles such as Nurse, Firefighter and Basketball Player show pronounced gender and racial skews that exceed labor-force statistics (U.S.
-Bureau of Labor Statistics, 2021), whereas DALL-E3’s guardrails (OpenAI, 2022b) mitigate many of these effects.
-More notably, Figure 8(d–f) highlights previously unnoticed bias dimensions.
+To quantify bias, we measured the normalized entropy of each distribution [d2024openbias] as bias intensity and identified the dominant cluster (the largest group) as the potential bias direction.
+We conducted a user study with 54 participants to validate our findings.
+$\mathcal{X}$-Cluster’s predicted bias intensity closely matched human ratings with an Absolute Mean Error of 0.1396 (0–1 scale) and aligned with human-identified bias directions 72.3% of the time.
+User study details are provided in Supp.
+Q.1.
+Findings: As shown in Figure 8, our method identifies both well-known and novel biases in occupational images without relying on predefined categories.
+For instance, Figure 8(a–c) reveals strong gender and racial imbalances in SDXL-generated images for roles like Nurse, Firefighter, and Basketball Player, exceeding official statistics [blsEmployedPersons].
+In contrast, DALL·E3 exhibits improved bias mitigation, likely due to its built-in “guardrails” [dalle_guards].
+More notably, Figure 8(d–f) highlights previously unrecognized bias dimensions.
 For example, SDXL strongly associates CEOs with “Grey” hair, while DALL·E3 favors “Dark” hair.
 Additionally, DALL·E3 shows stronger biases in Hair style and Grooming for occupations like Nurse (Figure 8(e)) and Teacher (Figure 8(f)).
-These findings suggest that while industrial T2I models with guardrails may have mitigated well-known biases, they often overlook emerging or less-discussed ones – highlighting the need for broader audits, where $\mathcal{X}$-Cluster can serve as a valuable tool.
-See further findings in App.
-P.1.
-Analyzing Social Media Image Popularity: What visual traits drive popularity?
-To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), where popularity is measured by image view count.
-$\mathcal{X}$-Cluster discovered ten grouping criteria and organizes photos into semantic clusters under each.
-Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters for three criteria.
+These findings suggest that while industrial T2I models with guardrails may address well-known biases, they may still overlook emerging or less-discussed ones, underscoring the need for broader bias analysis.
+For additional findings and experimental details, see Supp.
+Q.1.
+6.2 Analyzing Social Media Image Popularity
+What makes a photo popular?
+To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset [ortis2019prediction], where popularity is measured by image view count.
+$\mathcal{X}$-Cluster discovered 10 grouping criteria and organizes photos into semantic clusters under each.
+Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters across three criteria.
 Findings: As shown in Figure 9, combining $\mathcal{X}$-Cluster’s grouping with popularity scores provides a direct interpretation of the visual elements that drive trends versus those that define widely uploaded images.
@@
 “Neutral emotion”.
-These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era (McSpadden, 2015; Farid, 2024), underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior.
-For full findings and additional analysis, see App.
-P.2.
-Confirming and Mitigating Dataset Bias: Lastly, in App.
-P.3, we used $\mathcal{X}$-Cluster to confirm the gender bias in 162k CelebA (Liu et al., 2015) training images and successfully trained a debiased model using the predicted groupings with DRO (Sagawa et al., 2020).
-Tab.
-35 shows our debiased model achieves robustness comparable to B2T (Kim et al., 2024), which uses annotations for bias detection.
+These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era [mcspadden2015you, farid_2024], underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior.
+For full findings and additional analysis, see Supp.
+Q.2.
 Figure 9: Social media image popularity analysis.
@@
 7 Conclusion
-We introduced the OpenSMC task and proposed $\mathcal{X}$-Cluster for discovering grouping criteria and semantic substructures in an unstructured image collections, expressed in natural language.
-We evaluated $\mathcal{X}$-Cluster on six benchmarks and three applications, showing its ability to automatically discover valuable insights stored in visual data.
+We introduce the OpenSMC task and propose $\mathcal{X}$-Cluster, a system that discovers interpretable grouping criteria and substructures in image collections, effectively extracting valuable insights across six datasets and three applications.
 Acknowledgments
-This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGenerationEU.
-E.R.
-acknowledges partial support from the PRECRISIS project, funded by the EU Internal Security Fund (ISFP-2022-TFI-AG-PROTECT-02-101100539), the EU project SPRING (No.
-871245), and the PRIN project LEGO-AI (Prot.
-2020TA3K9N).
-This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN.
-We thank OpenAI for supporting our experiments through the Researcher Access Program (0000004498).
-We extend our gratitude to Feng Xue for his valuable suggestions on plot creation.
+This work was supported by the EU Horizon projects ELIAS (No.
+101120237) and ELLIOT (No.
+101214398).
+We greatly acknowledge CINECA and the ISCRA initiative for providing high-performance computing resources.
+We also extend our gratitude to Feng Xue for his valuable suggestions on plot creation.
 M.L.
 would like to thank Margherita Potrich for her unwavering support.
-Limitations
-As discussed in § 5.1 and § 5.2, $\mathcal{X}$-Cluster still has significant room for improvement.
-It is thus not guaranteed to uncover all valid grouping criteria or perfectly align semantic substructures with user preferences.
-Moreover, although designed as an automatic, assistive tool for large-scale visual data analysis, $\mathcal{X}$-Cluster requires human oversight; users bear the ultimate responsibility for correctly interpreting the generated results.
-Since $\mathcal{X}$-Cluster relies on multimodal large language models (MLLMs) and large language models (LLMs), it inherits their limitations, including potential errors, biases, or hallucinations.
-Foundation models such as LLaVA-Next (Liu et al., 2024b), CLIP (Radford et al., 2021), and Llama series (Meta, 2024b, a) reflect biases present in their internet-scale training data (Goh et al., 2021; Schramowski et al., 2022; Kirk et al., 2021; Yamada et al., 2023; Wan et al., 2023).
-These biases can be challenging to identify or quantify due to the vast—and sometimes proprietary—nature of the training sets (Radford et al., 2021; OpenAI, 2022a, 2023).
-Furthermore, both MLLMs and LLMs may suffer from hallucinations.
-For LLMs (Ji et al., 2023; Yao et al., 2023a), this means generating information that appears accurate but is not.
-For MLLMs (or VQA models) (Ge et al., 2024), it involves creating image descriptions that seem reasonable but do not accurately reflect the visual content.
-To alleviate these limitations, we provide detailed analyses of failure cases (App.
-I) for our proposed $\mathcal{X}$-Cluster framework, strategies for handling hallucinated or invalid criteria (App.
-K), and methods to mitigate biases (App.
-L) in the technical appendix.
-Lastly, due to the zero-shot nature of $\mathcal{X}$-Cluster, its limitations may be mitigated by incorporating improved foundation models (OpenAI, 2025), employing bias alleviation techniques (Kong et al., 2024), or applying hallucination reduction strategies (Sahoo et al., 2024; Ge et al., 2024).
-Ethics Statement
+[…]
+Supplementary Material
+Overview
+This supplementary material provides additional details supporting the implementations, experiments, findings, and discussions in the main paper.
+In Supp.
+A, we provide a reproducibility statement to ensure the transparency and replicability of our work.
+Additionally, in Supp.
+B, we present an ethics statement to discuss and address the potential ethical implications and concerns associated with the proposed task and methodology, highlighting possible societal impacts and mitigation strategies.
+Supp.
+C expands on Related Work, covering relevant tasks and methods.
+Supp.
+D describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation.
+Supp.
+E details the evaluation metrics used in this work.
+Supp.
+F presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers.
+Supp.
+G provides implementation specifics of the compared methods.
+Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in Supp.
+H.
+For qualitative analysis, Supp.
+I presents further visualizations of predicted clusters, while Supp.
+J examines failure cases.
+Supp.
+K investigates the impact of multi-granularity clustering.
+Supp.
+L explores the effect of invalid (hallucinated) criteria on system performance, and Supp.
+M studies the influence of foundation model hallucinations and biases.
+Supp.
+N analyzes the computational cost and runtime of $\mathcal{X}$-Cluster.
+Supp.
+O and Supp.
+P extend our analysis with studies on system sensitivity and fine-grained image collections.
+Supp.
+Q provides additional findings, implementation details, and user study results for the three explored applications.
+Supp.
+R discusses how LLMs enhance image clustering.
+Finally, Supp.
+S outlines potential future research directions for our proposed OpenSMC task.
+The Table of Contents on the next page outlines the main topics in this supplementary material, with hyperlinks for direct navigation to each section.
+Appendix A Reproducibility Statement
+Upon publication, we will open-source all essential resources for reproducing this work.
+Specifically, we will provide the full code implementation of $\mathcal{X}$-Cluster, along with the exact prompts used in each module.
+Additionally, we will release our two newly proposed benchmarks, COCO-4c and Food-4c, including annotations for each grouping criterion.
+Lastly, we will provide the code for our evaluation protocol, experiments, and application studies.
+Appendix B Ethics Statement
 We do not anticipate any immediate negative societal impacts from our work.
@@
 We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content.
-Specifically, in § 6, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016).
+Specifically, in Sec.
+6, we use well-established benchmarks, such as CelebA [liu2015deep], for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies [bianchi2023easily, bolukbasi2016man].
 However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups.
@@
 Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
-For the creation of our two new benchmarks, COCO-4c and Food-4c, we sourced images exclusively from the COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) datasets, strictly adhering to their licensing agreements.
+For the creation of our two new benchmarks, COCO-4c and Food-4c, we sourced images exclusively from the COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] datasets, strictly adhering to their licensing agreements.
 Additionally, we utilized voluntary human annotators for proposing valid grouping criteria and creating annotations along these criteria, rather than employing annotators from crowdsourcing platforms.
@@
 We refrain from using APIs from industrial LLMs or MLLMs, both to ensure reproducibility and to protect data privacy.
-Appendix
-[…]
-[…]
-Overview
-This supplementary material provides additional details supporting the implementations, experiments, findings, and discussions in the main paper.
-In App. A, we first provide a reproducibility statement to ensure the transparency and replicability of our work.
-App.
-B expands on Related Work, covering relevant tasks and methods.
-App.
-C describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation.
-App.
-D details the evaluation metrics used in this work.
-App.
-E presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers.
-App.
-F provides implementation specifics of the compared methods.
-Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in App.
-G.
-For qualitative analysis, App.
-H presents further visualizations of predicted clusters, while App.
-I examines failure cases.
-App.
-J investigates the impact of multi-granularity clustering.
-App.
-K explores the effect of invalid (hallucinated) criteria on system performance, and App.
-L studies the influence of foundation model hallucinations and biases.
-App.
-M analyzes the computational cost and runtime of $\mathcal{X}$-Cluster.
-App.
-N and App.
-O extend our analysis with studies on system sensitivity and fine-grained image collections.
-App.
-P provides additional findings, implementation details, and user study results for the three explored applications.
-App.
-Q discusses how LLMs enhance image clustering.
-Finally, App.
-R outlines potential future research directions for our proposed OpenSMC task.
-The Table of Contents on the next page outlines the main topics in this supplementary material, with hyperlinks for direct navigation to each section.
-Appendix A Reproducibility Statement
-Upon publication, we will open-source all essential resources for reproducing this work.
-Specifically, we will provide the full code implementation of $\mathcal{X}$-Cluster, along with the exact prompts used in each module.
-Additionally, we will release our two newly proposed benchmarks, COCO-4c and Food-4c, including annotations for each grouping criterion.
-Lastly, we will provide the code for our evaluation protocol, experiments, and application studies.
-Appendix B Additional Related Work
+Appendix C Additional Related Work
 Topic Discovery.
-The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund and Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents).
+The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery [blei2003latent, wang2009multi, eklund2022topic] in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents).
 Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data.
@@
 Nevertheless, OpenSMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them.
-Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024a).
+Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models [meta2024llama31].
 Multimodal Large Language Model.
 Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data.
-The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024b; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024a) alongside a pre-trained vision encoder (Radford et al., 2021).
+The typical paradigm [liu2024visual] involves using a pre-trained large language model (LLM) [meta2024llama3, vicuna2023, jiang2023mistral, meta2024llama31] alongside a pre-trained vision encoder [radford2021learning].
 A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs.
-Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023a), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2024), and the LLaVA series (Liu et al., 2024c, b; Li et al., 2025).
+Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 [li2023blip], BLIP-3 [xue2024xgen], Kosmos-2 [peng2023kosmos], and the LLaVA series [liu2024visual, liu2024llava, li2024llava].
 In our proposed $\mathcal{X}$-Cluster framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria.
-Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2025) to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2024) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
+Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave [li2024llava] to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style [shao2023prompting, zhu2023chatgpt] as the image-based grouper to form semantic clusters linked to specific visual content within the images.
 […]
 Text-Driven Image Retrieval.
-Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query.
+Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods [karthik2024visionbylanguage, liu2023zero, wu2023cap4video] aim to find images from an image collection that are relevant to the query.
 In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques.
@@
 Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish OpenSMC.
-Appendix C Benchmark Details
-C.1 Construction of COCO-4c and Food-4c
+Appendix D Benchmark Details
+D.1 Construction of COCO-4c and Food-4c
 Table 1: Summary of number of classes for the basic criteria annotation across the six benchmarks.
@@
 To create high-quality benchmarks for COCO-4c and Food-4c, we designed a four-step annotation pipeline:
-(1) Criteria Identification: We first split COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images into batches of 100.
-Each batch was stitched into a 10$\times$10 grid to form a single image.
+(1) Criteria Identification: We first split COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] images into batches of 100.
+Each batch was stitched into a $\times10$ grid to form a single image.
 These grid images were then distributed to 5 human annotators, who were tasked with identifying grouping criteria.
 For each dataset, we selected the 4 most frequently occurring criteria, as shown in Tab.
-3, to proceed with per-image annotation.
-(2) Label Candidate Generation: To facilitate the annotation process, we used GPT-4V (OpenAI, 2023) to generate an initial list of candidate labels for each criterion.
+1, to proceed with per-image annotation.
+(2) Label Candidate Generation: To facilitate the annotation process, we used GPT-4V [openai2023gpt4] to generate an initial list of candidate labels for each criterion.
 Specifically, for each criterion of COCO-4c and Food-4c, GPT-4V was prompted to assign a label that reflected the criterion for each image.
@@
 The entire annotation process took 25 days to complete.
-[…]
+(4) Label Merging: Image annotation is inherently subjective, with annotators potentially assigning different labels for the same criterion.
+For example, one annotator might label the Mood criterion as “Happy”, while another might label it as “Joyful” or “Delightful”.
+To resolve such discrepancies, we used majority voting to determine the final label for each image.
+Specifically, the most frequently assigned label among the 10 annotators was chosen as the final label for each criterion.
 Following these steps, we constructed COCO-4c and Food-4c.
-Note that we used the official COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images for our benchmarks and did not collect any new images.
+Note that we used the official COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] images for our benchmarks and did not collect any new images.
 We adhered strictly to the licenses of the datasets during their creation.
 The exact number of classes is presented in Tab.
+1.
+Additionally, the annotated class names for each criterion of COCO-4c are provided in Tab.
+2, and for Food-4c in Tab.
 3.
-Additionally, the annotated class names for each criterion of COCO-4c are provided in Tab.
-4, and for Food-4c in Tab.
-5.
 Table 2: Full class names for COCO-4c across the four basic criteria.
@@
 Diet “omnivore”, “vegan”, “vegetarian”, “gluten free”
-C.2 Details on Hard Grouping Criteria Annotation
+D.2 Details on Hard Grouping Criteria Annotation
 Table 4: Annotated criteria for the six benchmarks.
@@
 Fruit Quantity and Arrangement
-In Tab. 6, we present the additional annotated Hard grouping criteria ground truth alongside the Basic criteria for each benchmark.
+In § D.2, we present the additional annotated Hard grouping criteria ground truth alongside the Basic criteria for each benchmark.
 While we have established more rigorous and challenging benchmarks such as COCO-4c and Food-4c, which feature up to four distinct grouping criteria, these annotated criteria sets do not encompass all potential grouping criteria within the image collections.
@@
 The procedure for obtaining the Hard grouping criteria is as follows:
-(1) Criteria Discovery: We divided each dataset into batches of 100 images, displaying each batch in a 10$\times$10 grid.
+(1) Criteria Discovery: We divided each dataset into batches of 100 images, displaying each batch in a $\times10$ grid.
 Five human annotators were assigned to each batch and instructed to identify as many valid grouping criteria as possible.
@@
 (2) Criteria Merging: After collecting the annotated criteria from all five annotators, we aggregated the criteria and manually cleaned the set by merging semantically similar criteria (e.g., Location and Place) and discarding binary grouping criteria, as the inclusion of binary criteria can result in an unmanageable number of grouping criteria for complex datasets.
-By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in Tab.
-6.
+By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in § D.2.
 This resulted in sets containing 8 criteria for Fruit-2c, 4 criteria for card, 11 criteria for Action-3c, 7 criteria for Clevr-4c, 17 criteria for COCO-4c, and 11 criteria for Food-4c.
 These expanded ground-truth sets enable us to more effectively evaluate the capabilities of various criteria discovery methods, providing a clearer understanding of different criteria proposers.
-Appendix D Further Details of Evaluation Protocol
+Appendix E Further Details of Evaluation Protocol
 Further Discussion on Clustering Accuracy (CAcc).
-Clustering Accuracy (CAcc) (Han et al., 2021) is evaluated by applying the Hungarian algorithm (Kuhn, 1955) to determine the optimal assignment between the predicted cluster indices and ground-truth labels.
-As extensively discussed in the GCD (Vaze et al., 2022) literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation.
+Clustering Accuracy (CAcc) [han2021autonovel] is evaluated by applying the Hungarian algorithm [kuhn1955hungarian] to determine the optimal assignment between the predicted cluster indices and ground-truth labels.
+As extensively discussed in the GCD [vaze2022generalized] literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation.
 On the other hand, if the number of predicted clusters is lower than the number of ground-truth classes, the extra classes are assigned to a null set, and all instances with those ground-truth labels are similarly considered incorrect.
@@
 […]
-Appendix E Further Implementation and Prompt Details
-In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App.
-E.1 and the Semantic Grouper in App.
-E.2.
-E.1 Further Details of Criteria Proposer
+Appendix F Further Implementation and Prompt Details
+In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in Supp.
+F.1 and the Semantic Grouper in Supp.
+F.2.
+Figure 1: All three variants of the proposed $\bm{\mathcal{X}}$-Cluster framework.
+We explore different design choices for both the Criteria Proposer (left) and the Semantic Grouper (right), and designate the best-performing Caption-based system (marked with Refer to caption) as the main $\mathcal{X}$-Cluster configuration in our experiments.
+Further, we present a system overview illustrating all three variants of our Criteria Proposer and Semantic Grouper, namely the caption based (main), tag based, and image based versions, in Figure 1.
+F.1 Further Details of Criteria Proposer
 Image-based Proposer:
@@
 In Tab.
-7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2025).
+5, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B [li2024llava].
 Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images.
@@
 For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab.
-7.
+5.
 After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set.
@@
 In Tab.
-8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024a).
-For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image.
-These tags are selected from the WordNet (Miller, 1995) vocabulary, which contains 118k English synsets, and represent the semantic content of the images.
+6, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B [meta2024llama31].
+For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 [radford2021learning], to assign 10 related natural language tags to each image.
+These tags are selected from the WordNet [miller1995wordnet] vocabulary, which contains 118k English synsets, and represent the semantic content of the images.
 We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging.
@@
 Your response:
-We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab.
-9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024a) in Tab.
-10.
+We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B [liu2024llava] in Tab.
+7, and the prompt for the LLM Llama-3.1-8B [meta2024llama31] in Tab.
+8.
 Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language.
@@
 Output Instruction Please respond with the cleaned and optimized list of clustering criteria, formatted as bullet points (using “*”). Your response:
-Since the Proposers operate on subsets, the criteria accumulated in $\mathcal{R}$ may include redundant or noisy entries.
-Therefore, we further include a criteria poop refinement process.
-Specifically, we input all initially proposed criteria from $\mathcal{R}$ into a LLM, prompting it to consolidate similar criteria and discard noisy ones.
-This process refines and updates the criteria pool $\mathcal{R}$ for next stage.
 In Tab.
-11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024a).
+9, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B [meta2024llama31].
 Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab.
@@
 This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
-E.2 Further Details of Semantic Grouper
+F.2 Further Details of Semantic Grouper
 Image-based Grouper:
@@
 Task Instruction Therefore, I need your assistance in designing a prompt for the Visual Question Answering (VQA) model to help it identify the "{CRITERION}" category in a given image at three different granularity.
-Please help me design and generate this prompt using the following template: "Question: [Generated VQA Prompt Question] Answer (reply with an abstract, a common, and a specific category name, respectively):".
+Please help me design and generate this prompt using the following template: ”Question: [Generated VQA Prompt Question] Answer (reply with an abstract, a common, and a specific category name, respectively):”.
 The generated prompt should be simple and straightforward.
@@
 The prompt for generating criterion-specific tags is shown in Tab.
-13, while the prompts for generating coarse-grained and fine-grained tags are shown in Tab.
+11, while the prompts for generating coarse-grained and fine-grained tags are shown in Tab.
 12 and Tab.
-15, respectively.
+13, respectively.
 In the tag-based grouper, we begin by embedding the target criterion into the prompt from Tab.
@@
 To enhance the diversity and coverage of the tags, we query the LLM 10 times and take the union of the generated tags after deduplication as candidates.
-Following the SHiNe framework (Liu et al., 2024d), for each middle-grained tag, we further embed it into the prompts from Tab.
+Following the SHiNe framework [liu2024shine], for each middle-grained tag, we further embed it into the prompts from Tab.
 12 and Tab.
@@
 After tagging all the images, we group those sharing the same tag into clusters, yielding the clustering result.
-Note that we do not utilize lexical databases such as WordNet (Miller, 1995) or ConceptNet (Speer et al., 2017) for tag generation, as they do not support free-form input and may not capture certain discovered criteria.
+Note that we do not utilize lexical databases such as WordNet [miller1995wordnet] or ConceptNet [speer2017conceptnet] for tag generation, as they do not support free-form input and may not capture certain discovered criteria.
 Caption-based Grouper:
@@
 Task Instruction Understand the provided description carefully and assign a label to this image based on what is the "{CRITERION}" depicted in this image.
-Output Instruction Please respond in the following format within five words: "*Answer*". Do not talk about the description and do not respond long sentences. The answer should be within five words.
+Output Instruction Please respond in the following format within five words: ”*Answer*”. Do not talk about the description and do not respond long sentences. The answer should be within five words.
 Task Reinforcement Again, your job is to understand the description and assign a label to this image based on what is the "{CRITERION}" shown in this image. Your response:
@@
 The categories in the third-level are subcategories of the second-level.
-Step 5 - Output the generated three-level class hierarchy as a JSON object where the keys are the level numbers and the values are a flat list of generated categories at each level, structured like: { “level 1": [“categories”], “level 2": [“categories”], “level 3": [“categories”] }
+Step 5 - Output the generated three-level class hierarchy as a JSON object where the keys are the level numbers and the values are a flat list of generated categories at each level, structured like: { “level 1”: [“categories”], “level 2”: [“categories”], “level 3”: [“categories”] }
 Output Instruction Please only output the JSON object in your response and simply use a flat list to store the generated categories at each level. Your response:
@@
 Your response:
-We first present the MLLM prompt used for LLaVA-NeXT-7B (Liu et al., 2024b) to generate criterion-specific captions in Tab.
-16.
+We first present the MLLM prompt used for LLaVA-NeXT-7B [liu2024llava] to generate criterion-specific captions in Tab.
+14.
 Following this, we present the LLM Llama-3.1-8B prompts used in the caption-based grouper for the Initial Naming step in Tab.
-17, the Multi-granularity Cluster Refinement step in Tab.
-18, and the Final Assignment step in Tab.
-19.
+15, the Multi-granularity Cluster Refinement step in Tab.
+16, and the Final Assignment step in Tab.
+17.
 Specifically, we begin by generating criterion-specific captions for each image using LLaVA-NeXT-7B with the prompt shown in Tab.
-16.
+14.
 For each image, we then embed its criterion-specific caption and the relevant criterion into the LLM prompt shown in Tab.
-17, querying the LLM to assign an initial name based on the target criterion.
+15, querying the LLM to assign an initial name based on the target criterion.
 Once the initial names for all images in the dataset are obtained, we embed these names along with the target criterion into the prompt in Tab.
@@
 Finally, for each image, we embed the target criterion, its criterion-specific caption, and cluster candidates from each granularity level into the prompt shown in Tab.
-19, and use this to query the LLM for final cluster assignment at each granularity level.
-Appendix F Further Details of the Compared Methods
+17, and use this to query the LLM for final cluster assignment at each granularity level.
+Appendix G Further Details of the Compared Methods
 […]
@@
 […]
-Implementation details of MMaP and MSub (Yao et al., 2024, 2025): We closely followed the training configuration outlined in the original MMaP and MSub paper.
+Implementation details of MMaP and MSub [yao2024multi, yao2025customized]: We closely followed the training configuration outlined in the original MMaP and MSub paper.
 Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset.
@@
 After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
-Appendix G Further Quantitative Experimental Results
+Appendix H Further Quantitative Experimental Results
 In this section, we present additional numerical experiment results to supplement the figures in the main paper.
-In § G.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework.
-In § G.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks.
+In § H.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework.
+In § H.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks.
 Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods.
-Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § G.4.
-G.1 Further Results for Criteria Proposer Study
+Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § H.4.
+H.1 Further Results for Criteria Proposer Study
 We provide detailed numerical results corresponding to Figure 4 in Tab. 18 and Figure 5 in Tab. 19 for the six tested benchmarks.
-Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g., introducing objects not present in the images) Liu et al.
-(2024a), these issues generally occur at the object or fine-grained attribute level.
+Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) [he2024incorporating] and hallucinations (e.g., introducing objects not present in the images) [liu2024survey], these issues generally occur at the object or fine-grained attribute level.
 However, when reasoning about grouping criteria for OpenSMC task, the focus is on identifying general thematic elements shared across the image set.
@@
 1 img 23.5 36.4 27.3 57.1 75.0 50.0 44.9
-G.2 Further Results for Semantic Grouper Study
+H.2 Further Results for Semantic Grouper Study
 In this section, we present the expanded numerical results comparing different semantic groupers to supplement the summary in Figure 6.
@@
 Color 15 66 5 15 39
-G.3 Further Comparative Results with TCMC Methods
+H.3 Further Comparative Results with TCMC Methods
 We provide expanded results in Tab. 27 for each criterion and benchmark, detailing the comparison of criteria-conditioned clustering methods presented in Tab. 2 in the main paper.
@@
 $Avg.$ 69.4 88.3 63.3 55.1 62.0 46.8 51.0 - 54.4 - 65.1 61.1
-G.4 Further Results for Multi-granularity Clustering Study.
+H.4 Further Results for Multi-granularity Clustering Study.
 We present expanded results in Tab. 28 for the ablation study on multi-granularity refinement, providing a detailed breakdown of the summary shown in Figure 7 in the main paper.
@@
 $Avg.$ 50.5 62.7 54.4 54.7 65.1 61.1
-Appendix H Qualitative Analysis
+Appendix I Qualitative Analysis
 In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
@@
 Figure 7: Example predicted clusters of COCO-4c at different granularities.
-Appendix I Failure Case Analysis
+Appendix J Failure Case Analysis
 In Figure 8, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper).
@@
 The “Edamame” shown in Figure 8 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine.
-This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
+This oversight of fine-grained details could be improved by employing a more advanced prompting strategy [liu2024democratizing].
 Figure 8: Failure case analysis. We show wrongly predicted images with their ground-truth label for four clusters.
-Appendix J Further Study on Multi-granularity Clustering
+Appendix K Further Study on Multi-granularity Clustering
 In this section, we provide a detailed study on how different levels of multi-granularity output from our $\mathcal{X}$-Cluster framework impact grouping results.
@@
 These experimental results not only highlight the importance of the multi-granularity output of our framework but also validate the effectiveness of our multi-granularity design in aligning with user-preferred granularities that is reflected by the annotations in these experiments.
-Appendix K Study on Handling Invalid Criteria
+Appendix L Study on Handling Invalid Criteria
 At the criteria refinement step, invalid grouping criteria (False Positives) may be proposed due to hallucinations from large language models (LLMs).
 While we did not observe hallucinated criteria being introduced during our experiments across six datasets and three application studies, it is important to further investigate the potential impact of such invalid criteria on the proposed $\mathcal{X}$-Cluster system.
-To this end, we design and conduct a control experiment using the Fruit-2c dataset (Muresan and Oltean, 2018), where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool.
+To this end, we design and conduct a control experiment using the Fruit-2c dataset [muresan2018fruit], where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool.
 These invalid criteria were then used in the subsequent grouping process to evaluate their effect on our system.
@@
 The grouping results for the two invalid criteria are presented in Tab.
-31.
+29.
 As observed, when processing invalid “hallucinated” criteria, nearly all images are assigned to a cluster named “Not visible” by our framework.
@@
 Others† 1.7 3.3
-Appendix L Study on Model Hallucination and Bias
+Appendix M Study on Model Hallucination and Bias
 Model hallucination.
-LLM hallucination (Wang et al., 2024c) typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like "Who was the 70th president of the United States?" might lead to a fabricated response.
+LLM hallucination [wang2024factuality] typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like ”Who was the 70th president of the United States?” might lead to a fabricated response.
 However, in our system, the use of LLMs is fully grounded in the visual descriptions (tags or captions) of the images.
@@
 That said, LLM hallucination can still have mild effects on clustering results.
-For example, as discussed in the failure case analysis in Appendix I, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 16).
+For example, as discussed in the failure case analysis in Appendix J, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 8).
 MLLMs also play a crucial role in our system, as they are responsible for translating images into text for subsequent processing steps.
-MLLM hallucination (Wang et al., 2024c) typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image.
+MLLM hallucination [wang2024factuality] typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image.
 However, since our proposed system operates at the dataset level rather than on a per-image basis, it is largely insensitive to such hallucinations, especially at the fine-grained visual detail level.
-Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker (Ge et al., 2024), which we leave as a direction for future work.
+Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker [ge2024visual], which we leave as a direction for future work.
 Model Bias.
-Foundation models such as LLMs and MLLMs are known to inherit biases from their training data (Bommasani et al., 2021).
+Foundation models such as LLMs and MLLMs are known to inherit biases from their training data [bommasani2021opportunities].
 In our system, we addressed potential biases using Hard Positive Prompting techniques: i) MLLM Bias Mitigation: The MLLM is further prompted to generate criterion-specific captions that focus solely on describing the criterion-related content in each image.
 This approach constrains the MLLM from generating irrelevant content influenced by inherent biases; ii) LLM Bias Mitigation: Similarly, when prompting the LLM to assign image captions to clusters, we condition it to concentrate exclusively on the Criterion depicted in each image (see Tab.
-17).
+15).
 To validate the effectiveness of these bias mitigation techniques, we conducted a fair clustering experiment.
-Specifically, following Kwon et al.
-(2024), we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET (Gustafson et al., 2023) dataset, which contains images from 52 occupations.
+Specifically, following kwon2023image, we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET [gustafson2023facet] dataset, which contains images from 52 occupations.
 For each occupation, we selected 10 images of men and 10 images of women, totaling 80 images, ensuring a ground-truth gender proportion disparity of 0% for each occupation.
@@
 Ours w. Negative prompt 48.4 51.6 3.2
-Appendix M Computational Cost Analysis
+Appendix N Computational Cost Analysis
 The proposed $\mathcal{X}$-Cluster framework is training-free, requiring only inference processes.
@@
 In Tab.
-33, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations.
-For these experiments, we used LLaVA-NeXT-7B (Liu et al., 2024b) as the MLLM and Llama-3.1-8B (Meta, 2024a) as the LLM.
-As shown in Tab.
-33, organizing 5,000 images based on all four discovered criteria can be completed by $\mathcal{X}$-Cluster in 29.1 hours on a single A100 GPU or 16.7 hours on a single H100 GPU.
-More importantly, most steps in our framework, such as per-image captioning and per-caption cluster assignment, are parallelizable across multiple GPUs, significantly accelerating the process.
-Therefore, when parallelizing the framework on 4 A100 or H100 GPUs, we achieve approximately a 4$\times$ speedup, reducing computational time to 7.6 hours on 4 A100 GPUs and 4.3 hours on 4 H100 GPUs.
+31, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations.
+For these experiments, we used LLaVA-NeXT-7B [liu2024llava] as the MLLM and Llama-3.1-8B [meta2024llama31] as the LLM.
+[…]
 Table 31: Computational cost analysis on the COCO-4c benchmark (5,000 images with four criteria).
@@
 4 Nvidia H100-80GB 3.1 4.3
-Appendix N System Sensitivity Analysis of Various MLLMs and LLMs
+Appendix O System Sensitivity Analysis of Various MLLMs and LLMs
 Figure 10: Sensitivity analysis of different MLLMs and LLMs on the six OpenSMC benchmarks.
@@
 Since all variants successfully propose the basic criteria in each benchmark, we report the average clustering accuracy (CAcc) and semantic accuracy (SAcc) across various criteria for comparative analysis.
-Specifically, in Figure 18(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024a) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b).
-Next, in Figure 18(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024b), and Llama-3.1-8B.
+Specifically, in Figure 10(a), we first fix the LLM in our system to Llama-3.1-8B [meta2024llama31] and assess the influence of various MLLMs: GPT-4V [openai2023gpt4], BLIP-3-4B [xue2024xgen], and LLaVA-NeXT-7B [liu2024llava].
+Next, in Figure 10(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo [openai2023gpt4], GPT-4o [openai2024gpt4o], Llama-3-8B [meta2024llama3], and Llama-3.1-8B.
 Findings in Figure 10(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language.
@@
 Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM.
-This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
-Appendix O Study on Fine-grained Image Collections
+This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems [street2024llms], is relatively straightforward.
+Appendix P Study on Fine-grained Image Collections
 Image collections may include fine-grained grouping criteria, such as Bird species in bird photography.
-Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He and Peng, 2017).
+Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations [zhang2014part, vedaldi2014understanding, he2017fine].
 This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process.
@@
 However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
-We demonstrate this by enhancing our Caption-based Grouper with FineR (Liu et al., 2024e), a cross-modal CoT prompt method specifically designed for fine-grained visual recognition.
+We demonstrate this by enhancing our Caption-based Grouper with FineR [liu2024democratizing], a cross-modal CoT prompt method specifically designed for fine-grained visual recognition.
 When the proposer identifies fine-grained criteria, such as Bird species, the framework switches to a FineR-enhanced captioning strategy that provides more detailed attribute descriptions, such as “Wing color: Blue-grey,” to enrich the captions and capture per-attribute visual characteristics to better support the subsequent substructure uncovering process.
@@
 We report clustering accuracy (CAcc) and semantic accuracy (SAcc) as percentages (%).
-The pseudo upper-bound (UB) performance is obtained using CLIP (Radford et al., 2021) ViT-L/14 in a zero-shot transfer setting with the ground-truth class names.
-†: We compare with FineR (Liu et al., 2024e) without its post-class name refinement step to ensure a fair comparison.
+The pseudo upper-bound (UB) performance is obtained using CLIP [radford2021learning] ViT-L/14 in a zero-shot transfer setting with the ground-truth class names.
+†: We compare with FineR [liu2024democratizing] without its post-class name refinement step to ensure a fair comparison.
 CUB200 Car196
@@
 Ours + FineR 45.1 68.9 31.1 47.3
-We evaluate this on two image collections containing fine-grained criteria: CUB200 (Wah et al., 2011) and Stanford Cars196 (Khosla et al., 2011).
+We evaluate this on two image collections containing fine-grained criteria: CUB200 [wah2011caltech] and Stanford Cars196 [khosla2011novel].
 Our framework successfully discovers the fine-grained criteria Bird species for CUB200 and Car model for Cars196.
 As shown in Tab.
-34, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself.
+32, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself.
 This demonstrates the flexibility of our system, allowing future adaptations to specific application needs, such as fine-grained image collections.
-Appendix P Further Details of the Application Study
-In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in § 6 of the main paper.
-Specifically, § P.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020).
-§ P.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria.
-Finally, § P.2 provides additional insights from the analysis of social media image popularity.
-P.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
+Appendix Q Further Details of the Application Study
+In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in Sec.
+6 of the main paper.
+Specifically, § Q.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO [sagawa2019distributionally].
+§ Q.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria.
+Finally, § Q.2 provides additional insights from the analysis of social media image popularity.
+Q.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
 Figure 11: Bias quantification results and human evaluation for each occupation and criterion across the two studied T2I models, DALL·E3 and SDXL. The bias intensity score is reported.
-Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator).
-We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study.
+Image Generation for the Subject Occupation: Following prior studies [bianchi2023easily, bolukbasi2016man], we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator).
+We used two state-of-the-art T2I diffusion model, DALL·E3 [betker2023improving] and Stable Diffusion (SDXL) [podell2023sdxl] to generate 100 images for each occupation for our study.
 This resulted in a total of 1,800 images.
 For each occupation, we provide some examples of images generated by DALL·E3 in Figure 14, while provide some examples of images generated by SDXL in Figure 15.
-We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
+We only used the simple prompt “A portrait photo of a <<OCCUPATION>>” for image generation for all occupations and did not include any potential biases in the prompt.
 Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set.
@@
 To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction.
-We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by D’Incà et al.
-(2024):
+We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by d2024openbias:
 $\mathcal{H}_{bias}^{l}=1+\frac{\sum_{c^{l}\in\mathcal{C}^{l}}\log(p(c^{l}|\mathcal{C}^{l},\mathcal{D}_{\text{Occupation}}))}{\log(|\mathcal{C}^{l}|)}$ (1)
@@
 Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach.
-Detailed user study results are provided in § P.1.
+Detailed user study results are provided in § Q.1.
 We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers.
@@
 A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g., Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g., Hair color, Mood, Attire, and Accessories).
-We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022b), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs.
+We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” [dalle_guards], designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs.
 However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied.
 This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
-P.2 Further Details on Analyzing Social Media Image Popularity
+Q.2 Further Details on Analyzing Social Media Image Popularity
 Figure 12: Complete analysis of social media photo popularity on the SPID dataset.
@@
 With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems.
-Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts.
+Traditional research often approaches image popularity as a regression problem [ortis2019prediction, cheng2024retrieval], utilizing metadata like hashtags, titles, or follower counts.
 However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored.
@@
 These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
-To expand on the discussion in § 6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 20.
+To expand on the discussion in Sec.
+6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 12.
 Notably, we consistently observed a sharp semantic contrast between the visual elements in top trending images and those in the mainstream images across all ten criteria.
@@
 This recurring observation reinforces the idea that viral (or trending) content tends to capture more attention, likely because it features novel, surprising, or striking visual elements.
-Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer and Gore, 2014).
+Humans are inherently attracted to stimuli that deviate from the norm [priester2004and, bruni2012role, palmer2014theory].
 On the other hand, widely uploaded yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality.
 We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
-P.3 Confirming and Mitigating Dataset Bias
-Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024).
-As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
+Q.3 Confirming and Mitigating Dataset Bias
+Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations [geirhos2020shortcut], we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations [sagawa2019distributionally] or post hoc misclassified images [kim2024discovering].
+As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA [liu2015deep] dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
 Figure 13: Results of dataset bias discovery and mitigation. Worst group and average accuracies(%) are reported.
@@
 DRO+B2T 90.4 93.2
-DRO+$\mathcal{X}$-Cluster 90.9 93.1
+[…]
 DRO+GT 89.7 93.6
@@
 Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”.
-To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels.
+To further validate this observation, following B2T [kim2024discovering], we used our predicted distributions to train a debiased model with GroupDRO [sagawa2019distributionally] and compared it with other unsupervised bias mitigation methods, including JTT [liu2021just], CNC [zhang2022correct], B2T, and GroupDRO trained with ground-truth labels.
 As shown in Tab.
-35, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
-Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset (Liu et al., 2015) to assess the classification accuracy of them.
+33, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
+Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset [liu2015deep] to assess the classification accuracy of them.
 Our method achieved an impressive classification accuracy of 99.1% for gender and 87.4% for hair color on the 162,770 training images, demonstrating its effectiveness for uncovering gender and hair color substructures within the training set.
-In addition, we quantified the spurious correlation between hair color and gender using the metric proposed by Yang et al.
-(2023).
+In addition, we quantified the spurious correlation between hair color and gender using the metric proposed by yang2023change.
 Specifically, given the correlated gender attribute distribution $A$ and the target hair color distribution $Y$, we computed the normalized mutual information between $A$ and $Y$ to quantify the spurious correlation as:
@@
 This demonstrates that our method effectively identifies and confirms the bias directly from the training set.
-Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO (Sagawa et al., 2020), we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution.
-We closely followed the training protocol outlined in B2T (Kim et al., 2024) and GroupDRO (Sagawa et al., 2020).
-Specifically, we fine-tuned a ResNet-50 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), using the training split of the CelebA dataset (Liu et al., 2015).
-The training was performed using the SGD optimizer (Ruder, 2016) with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$.
+Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO [sagawa2019distributionally], we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution.
+We closely followed the training protocol outlined in B2T [kim2024discovering] and GroupDRO [sagawa2019distributionally].
+Specifically, we fine-tuned a ResNet-50 [he2016deep] model pre-trained on ImageNet [deng2009imagenet], using the training split of the CelebA dataset [liu2015deep].
+The training was performed using the SGD optimizer [ruder2016overview] with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$.
 We applied a weight decay of 0.1 and set the group adjustment parameter to zero.
@@
 The final evaluation and comparison results are provided in Tab.
-35.
-Appendix Q Why LLMs Improve Image Clustering?
+33.
+Appendix R Why LLMs Improve Image Clustering?
 The most compelling aspect of this work lies in our $\mathcal{X}$-Cluster framework’s ability to transform large volumes of unstructured images into natural language and leverage the advanced text understanding and summarization capabilities of LLMs to tackle the challenging Open-ended Semantic Multiple Clustering (OpenSMC) task.
-This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain (Eklund and Forsman, 2022).
+This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain [eklund2022topic].
 Our core motivation is: “If LLMs can discover topics from documents and organize them, then by converting images into text, we can similarly use LLMs to organize unstructured images.”
-Traditional clustering methods (Estivill-Castro, 2002; Caron et al., 2018; Van Gansbeke et al., 2020; Li et al., 2024; Yu et al., 2024) often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable.
+Traditional clustering methods [estivill2002so, caron2018deep, van2020scan, li2023image, yu2024multiple] often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable.
 These limitations hinder their applicability to diverse datasets in open-world scenarios, as they demand significant human priors and retraining for each new dataset.
-In contrast, LLMs (OpenAI, 2022a, 2023; Touvron et al., 2023; Meta, 2024b, a) excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content).
-Operating in a zero-shot (Kojima et al., 2022), interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training.
-By integrating LLMs with MLLMs (Liu et al., 2024b) into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset.
+In contrast, LLMs [chatgpt, openai2023gpt4, touvron2023llama, meta2024llama3, meta2024llama31] excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content).
+Operating in a zero-shot [kojima2022large], interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training.
+By integrating LLMs with MLLMs [liu2024llava] into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset.
 This design allows our framework to overcome the rigid assumptions of traditional clustering methods, making it automatic, generalizable, and training-free.
@@
 The main challenge of employing LLMs for the SMC task lies in accurately translating visual content from images into natural language that LLMs can effectively reason with.
-This is evident from the sensitivity analysis results in App.
-N: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 18 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 18 (b)).
+This is evident from the sensitivity analysis results in Supp.
+O: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 10 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 10 (b)).
 In other words, the quality of image captions generated by MLLMs is critical for the effective use of LLMs in the SMC task.
@@
 In the second stage (semantic grouping), criterion-specific captions should precisely capture relevant visual content to provide accurate information for assigning images to clusters.
-To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling (Liu et al., 2024c), or stronger models like GPT-4V (OpenAI, 2023) can improve comprehensiveness.
-For better precision, advanced prompting methods like CoT (Wei et al., 2022) or FineR (Liu et al., 2024e) can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker (Ge et al., 2024) can reduce noise caused by hallucinations.
+To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling [liu2024visual], or stronger models like GPT-4V [openai2023gpt4] can improve comprehensiveness.
+For better precision, advanced prompting methods like CoT [wei2022chain] or FineR [liu2024democratizing] can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker [ge2024visual] can reduce noise caused by hallucinations.
 However, these techniques increase computational costs and framework complexity.
 In this work, we choose to keep $\mathcal{X}$-Cluster simple yet effective, and we outline these potential improvements for future practitioners.
-Appendix R Future Work
+Appendix S Future Work
 Closed-Loop Optimization.
-In this work, we designed our prompts following the Iterative Prompt Engineering methodology (DeepLearning.AI, 2024) introduced by Isa Fulford and Andrew Ng.
-In App.
-E, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice.
+In this work, we designed our prompts following the Iterative Prompt Engineering methodology [deeplearningai_2024] introduced by Isa Fulford and Andrew Ng.
+In Supp.
+F, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice.
 These explanations cover elements such as system prompts, input formatting, task and sub-task instructions, and output instructions.
@@
 However, in future work or application scenarios where a labeled training/validation dataset is available, practitioners could build upon our design objectives.
-By leveraging our proposed evaluation metrics (see § 5) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
+By leveraging our proposed evaluation metrics (see LABEL:sec:method_assessment) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
 $\bm{\mathcal{X}}$-Cluster on Other Data Types.
@@
 Moreover, since natural language is a highly versatile and widely-used medium of representation, $\mathcal{X}$-Cluster can be extended to other data types by converting these data into text (by replacing the captioning module with suitable tools) in future work, such as:
-  • Audio Data: Speech-to-Text models like Whisper (Radford et al., 2023) can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
-  • Tabular Data: Table-to-Text models, such as TabT5 (Andrejczuk et al., 2022), can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster.
+  • Audio Data: Speech-to-Text models like Whisper [radford2023robust] can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
+  • Tabular Data: Table-to-Text models, such as TabT5 [andrejczuk2022table], can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster.
 For tables containing figures, modern MLLMs like LLaVA-Next, which support both OCR and image-to-text capabilities, can handle these elements to create a unified textual representation for $\mathcal{X}$-Cluster.
-  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT (Wang et al., 2024a), can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
-  • Point Cloud Data: 3D captioning models, like Cap3D (Luo et al., 2024a), can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
+  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT [wang2024protchatgpt], can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
+  • Point Cloud Data: 3D captioning models, like Cap3D [luo2024scalable], can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
 We believe the versatile nature of $\mathcal{X}$-Cluster has the potential to open up a broad range of applications across diverse data modalities, fostering new directions in future research.
 Figure 14: Samples of DALL·E3 generated images.
-For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to DALL·E3 to generate 100 images.
+For each occupation, the simple prompt “A portrait photo of a <<OCCUPATION>>”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to DALL·E3 to generate 100 images.
 We present a random sample of 30 generated images.
 Figure 15: Samples of SDXL generated images.
-For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to SDXL to generate 100 images.
+For each occupation, the simple prompt “A portrait photo of a <<OCCUPATION>>”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to SDXL to generate 100 images.
 We present a random sample of 30 generated images.